### v1

# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

Abstract
Differential privacy (DP) enables private data analysis but is hard to use in practice. For data controllers who decide what output to release, choosing the amount of noise to add to the output is a non-trivial task because of the difficulty of interpreting the privacy parameter $\epsilon$. For data analysts who submit queries, it is hard to understand the impact of the noise introduced by DP on their tasks.
To address these two challenges: 1) we define a privacy risk indicator that indicates the impact of choosing $\epsilon$ on individuals’ privacy and use that to design an algorithm that chooses $\epsilon$ automatically; 2) we introduce a utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks. We implement the algorithm and the protocol inside a new platform built on top of a data escrow, which allows the controller to control the data flow and achieve trustworthiness while maintaining high performance. We demonstrate our contributions through an IRB-approved user study, extensive experimental evaluations, and comparison with other DP platforms. All in all, our work contributes to making DP easier to use by lowering adoption barriers.
1 Introduction
Analyzing private data is essential for various tasks, and differential privacy (DP) is a promising tool to achieve privacy. In a typical DP deployment, data analyst runs queries on a dataset hosted on a platform that is accessible to a data controller. The dataset contains information about individuals, and releasing this information risks revealing the identity of those individuals. The data controller wants to protect their identity while releasing useful output to the data analyst. DP guarantees that the output of the query is likely to be the same if we change or remove any individual’s record in the dataset. This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: smaller $\epsilon$ leads to more accurate output but weaker privacy guarantees. Despite bespoke deployments by large organizations such as the US Census Bureau[3], Apple[5], and Google[24], it remains difficult to use DP in practice [23][43]. In this paper, we consider two important challenges controllers and analysts face.
Challenge 1. The controller’s perspective. Controllers need to decide the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$. Dwork et al.[23] showed that controllers have no clear consensus on how to choose $\epsilon$ or how to approach this problem in a principled manner.
Our contribution. In this paper, we observe the crux of the problem is that $\epsilon$ does not directly translate into the degree of privacy loss incurred to specific individuals in the dataset. In response, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual. Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a query output affects individuals’ privacy. And because the PRI and $\epsilon$ are connected in a principled way, there is an opportunity to automatically find an $\epsilon$ that produces a profile of privacy risk indicators consistent with the controller’s privacy requirement. We exploit this opportunity by designing an algorithm that automatically finds an $\epsilon$ and releases query output such that all individuals receive adequate protection based on the controller’s privacy requirement.
Challenge 2. The analyst’s perspective. Analysts submit queries and receive noisy output that deviates from the ground truth output, i.e., the output if DP had not been applied. Without access to such ground truth, analysts do not know the impact of DP on the quality of their downstream tasks; this hinders the analysts’ ability to use the DP output in practice.
Our contribution. In this paper, we propose a utility signaling protocol to help analysts interpret the impact of DP on their downstream tasks while still allowing controllers to fully control the data flow. In the protocol we propose, the analyst delegates their downstream task with an associated evaluation function to the platform. With the controller’s permission, the platform invokes the evaluation function on both the DP output and the ground truth output and releases both results, thereby signaling the utility loss on the analyst’s downstream task due to noise introduced by DP.
Putting everything together: Escrow-Based Platform. The utility signaling protocol requires analysts to place more trust in the platform than in a traditional DP deployment. We want to make the platform trustworthy while maintaining high performance, which is at odds with the increased overhead introduced by the algorithm we propose to help controllers. Our solution is to build a DP platform on top of a data escrow, which permits controlling data flow and improving trust with significantly lower overhead compared to other DP platforms.
Summary of related work. Many approaches have been proposed to help controllers choose $\epsilon$ [30] [26] [23] [36] [39] [40] [28] [48] [32] [28], [26] [30] [36] [39] [40], however, none concentrates on communicating the effects of $\epsilon$ on specific individuals represented in the dataset, as we do. In addition, approaches for communicating the privacy guarantees offered by DP are mainly targeted toward end-users[41][25], i.e., the privacy risk when sharing their data, whereas our approach focuses on helping controllers, who can only choose $\epsilon$ if they understand its privacy implications on those end-users.
Exposing error statistics, e.g., error bound, can help analysts interpret the output; such is the approach followed by the US Census [13]. However, such statistics are not sufficient in many cases. Our utility signaling protocol also complements existing approaches [47] to help analysts gain a deeper understanding of the effect of DP on their output.
Lastly, different deployment of DP has received a lot of attention [20] [42] [9] [4] [45] [46] [49] [10] [44] [11]. Our escrow-based implementation hits a new point in the design space between the degree of trust required by participants and the resulting performance.
[…]
The rest of the paper is organized as follows. We present the new privacy risk indicator in Section 2 and the algorithm we designed to automatically find $\epsilon$ and release DP output in Section 3. Section 4 presents the utility signaling protocol and the escrow-based platform. We finish with evaluation (Section 5), related work (Section 6) and conclusions in Section 7.
2 Privacy Risk Indicator (PRI)
In this section, we present the new privacy risk indicator. We first provide background on differential privacy.
Differential privacy (DP) [19][21] is a mathematical definition of privacy that bounds the effect a single individual can have on the output of a query. Formally, a randomized mechanism $\mathcal{M}$ satisfies $\epsilon$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
$\Pr(\mathcal{M}(x)\in O)\leq e^{\epsilon}\Pr(\mathcal{M}(x^{\prime})\in O).$
The parameter $\epsilon$ is a tunable parameter called privacy loss budget. A typical way of achieving differential privacy is by adding noise to the output. The amount of noise is controlled by $\epsilon$; smaller $\epsilon$ leads to larger noise and less accurate output. Therefore, $\epsilon$ is the key parameter for navigating the privacy-accuracy trade-off in differential privacy.
We present two important properties of differential privacy:
Post-processing. [22] DP is immune to post-processing, which means any computation on the output of a DP mechanism will not undermine the DP guarantees. Formally, let $\mathcal{M}(x)$ be an $\epsilon$-DP mechanism. For any arbitrary function $f$, $f(\mathcal{M}(x))$ is $\epsilon$-DP.
Sequential composition. [22] Sequentially combining multiple DP mechanisms on the same input dataset results in quantifiable degradation of privacy. Formally, given $k$ independent mechanisms $\mathcal{M}_{1},\ldots,\mathcal{M}_{k}$ that satisfy $\epsilon_{1},\ldots,\epsilon_{k}$-DP, then any arbitrary function $g(\mathcal{M}_{1},\ldots,\mathcal{M}_{k})$ is $\sum_{i=1}^{i=k}\epsilon_{i}$-DP.
2.1 Overview
A good $\epsilon$ provides adequate privacy protection to all individuals while making the output still useful to analysts. However, it is inherently difficult to reason about the privacy implications of choosing $\epsilon$ for an specific individual in the dataset. Because $\epsilon$ bounds the privacy loss for any individual in expectation, but its concrete effect on a specific individual’s privacy is unknown. This is the source of difficulty for controllers whose ultimate goal is to protect individuals in the dataset.
We observe that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice. To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to others. For instance, a controller can set an $\epsilon$, compute the corresponding query output, and observe the resulting PRI for each individual. If the PRIs are equal, then all individuals will be equally protected. If they look different, that is an indication that certain individuals are at a higher risk of being identified and a suggestion for the controller to try a different $\epsilon$. The direct connection between the new PRI and differential privacy that we explain next is key to both providing a principled understanding of the new indicator and designing strategies to find $\epsilon$, including an algorithm that finds $\epsilon$ automatically, which we will present in Section 3.
2.2 Privacy Risk Indicator: Formal Definition
We represent a dataset as a set of $n$ records $x=(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, drawn from a data universe $\mathcal{X}$. WLOG, we assume each individual (labeled $1$ to $n$) contributes exactly one record to $x$.
We denote an $\epsilon$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$. We consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers. Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and the privacy loss budget $\epsilon$ as input, we define the privacy risk indicator of individual $i\in[n]$ as follows:
Definition 1 (Privacy Risk Indicator).
$PRI_{i}=\lVert r-q(x_{-i})\rVert_{1},$
where $x_{-i}$ is the neighboring dataset of $x$ with individual $i$’s record $x_{i}$ removed. The difference is in the L1 norm.
Computing PRI is $\epsilon$-DP. Since $r$ is the output of an $\epsilon$-differentially private mechanism, computing PRI also satisfies $\epsilon$-differential privacy by DP’s post-processing property.
Connection between PRI and privacy loss. The PRI has a direct connection to the concept of privacy loss as defined in differential privacy. We formally show that when the Laplace Mechanism is used to compute the query output $r$, the connection is simple: larger PRI corresponds to larger privacy loss and vice versa.
First, we introduce the definition of privacy loss in DP[22].
Definition 2 (Privacy Loss).
Given two neighboring dataset $x,x^{\prime}\in\mathcal{X}^{n}$ such that $\lVert x-x^{\prime}\rVert_{1}\leq 1$, for an output $r\sim\mathcal{Y}$, the privacy loss incurred upon observing $r$ is defined as:
$\mathcal{L}_{\mathcal{M}(x,q,\epsilon)\|\mathcal{M}(x^{\prime},q,\epsilon)}^{(r)}=\ln\left(\frac{\Pr\left[\mathcal{M}(x,q,\epsilon)=r\right]}{\Pr\left[\mathcal{M}(x^{\prime},q,\epsilon)=r)\right]}\right).$
Next, we introduce the definition of Laplace Mechanism[22], a classic $\epsilon$-differentially private mechanism that we use to compute the query output $r$.
Definition 3 (Laplace Mechanism).
Given a query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$ and $\epsilon>0$, the Laplace Mechanism is defined as:
$\mathcal{M}_{L}(x,q,\epsilon)=q(x)+(Z_{1},\ldots,Z_{k}),$
where $Z_{i}\sim Lap(\frac{GS_{q}}{\epsilon})$ are i.i.d random variables drawn from the Laplace distribution  [^1] The global sensitivity of q in the $\ell_{1}$ norm is defined as $GS_{q}=\max|q(x)-q(x^{\prime})|$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
Using the Laplace Mechanism, the relative privacy loss of a specific individual $i$ can be expressed as:
$\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}$ $=\ln\left(\prod_{j=1}^{k}\frac{\exp(-\epsilon|r_{j}-q(x)_{j}|/GS_{q})}{\exp(-\epsilon|r_{j}-q(x_{-i})_{j}|/GS_{q})}\right)$
$=\frac{\epsilon(\lVert r-q(x_{-i})\rVert_{1}-\lVert r-q(x)\rVert_{1})}{GS_{q}}.$
Since $\lVert r-q(x_{-i})\rVert_{1}$ is the definition of $PRI_{i}$, there exists a clear connection between $PRI_{i}$ and individual $i$’s privacy loss: larger PRI corresponds to larger privacy loss. Similarly to PRI, the privacy loss for a specific individual is data and query-dependent; essentially, it is the privacy loss individual $i$ will suffer if they participate in the analysis and the corresponding query output $r$ is released, compared to if they do not participate. Note that since we are using the Laplace Mechanism, which is $\epsilon$-DP, the privacy loss will always be bounded by $\epsilon$.
We have shown that the PRI indicates the relative privacy loss of each individual if a given output is released using DP. A data controller can leverage the PRI for many purposes; one natural goal is to exploit the relationship between PRI and DP to choose an $\epsilon$ and the corresponding output to release. In the next section, we explain different approaches the controller can use to achieve this goal and propose an algorithm.
3 Deriving Epsilon based on PRI
In this section, we describe how to leverage the PRI we introduced to address Challenge 1. We explain the two general approaches of using the PRI to find $\epsilon$ in Section 3.1. We describe the key observations we made to determine whether the PRIs offer adequate protection and the detailed algorithm for choosing $\epsilon$ and releasing query output in Section 3.2. In Section 3.3, we consider multiple queries and discuss how to bound the total privacy loss. Finally, we discuss the algorithm complexities and extensions in Section 3.4.
3.1 Equal PRI means Equal Protection
The goal of DP is to protect each individual equally in expectation, such that no individual suffers additional privacy loss than others by participating in the analysis; the privacy loss is bounded by $\epsilon$. The PRI indicates the effect of $\epsilon$ on each individual, thus becoming an actionable indicator that aids in selecting a suitable value. We describe two approaches to help controllers choose $\epsilon$ that leads to exactly equal or statistically similar PRIs. We show that both are impractical, which motivates our new algorithm (Section 3.2).
Figure 1: Example of using DP to equalize the PRIs. The goal is to find the largest $\epsilon$ that can make the PRIs equal.
Can we make the PRIs equal? Figure 1 illustrates this using an example dataset with four records. If the output was released directly without adding noise (thus no DP), some individuals would be more at risk than others: those with higher PRI, as shown in the leftmost table. By decreasing $\epsilon$, more noise is added to the output, so the PRIs become more similar. Then, to protect all individuals equally, the controller should choose the $\epsilon$ that makes the PRIs equal. Continuing with the example of Figure 1, setting $\epsilon$ to $0.5$ leads to equal PRIs among the four individuals while $\epsilon=1$ does not. When $\epsilon=0.2$, since it produces larger noise, the PRIs will remain equal, but it also leads to less accurate output. Therefore, $\epsilon=0.5$ is the best choice among the three values as it achieves the best privacy-accuracy trade-off.
Although it helps as an illustration, making PRIs exactly equal across individuals is generally infeasible unless an excessive amount of noise is added to the query output, which is undesirable because this means the output will not be useful for the analyst. It is also possible that a non-zero $\epsilon$ does not exist for which all PRIs are equal, and the only solution is to add an infinite amount of noise and make the PRIs all infinity.
Making PRIs statistically similar? One strategy is to consider the PRIs similar enough if they are statistically similar. Concretely, we can take two biased samples of PRIs: the set of the highest PRIs and the lowest PRIs. Then, we test the null hypothesis $H0$: the two samples come from the same distribution, using the statistical hypothesis framework. We consider the PRIs to be similar enough if we fail to reject $H0$.
Since we choose to compare two biased samples, any hypothesis tests that are sensitive to differences in locations (e.g., Mann-Whitney[37], Kolmogorov-Smirnov[29] or Mood’s Median Test) do not work well in practice since the ranks or the medians of the two samples will likely be very different, thus the tests will always tend to reject the null hypothesis. On the other hand, if we change the strategy to choosing random samples of PRIs and compare them, it becomes too easy to conclude the samples are similar. Therefore, an alternative strategy is needed to determine whether the PRIs provide adequate protection.
3.2 The Find-$\epsilon$-from-PRI Algorithm
Two key observations we make lead to our strategy for determining whether the PRIs provide adequate protection.
O1: Compute ratio between two PRIs. The first observation is we should compute the ratio between the PRIs of two individuals to determine whether they receive similar protection, since the comparison depends on the magnitude of the PRI. Consider two cases: two individuals with PRI of 2 and 1, and another two individuals with PRI of 10 and 9. In both cases, the difference between the two PRIs is 1, but the latter is preferable for equal protection since the ratio is close to 1.
Figure 2: Left tail with similar larger PRIs (green line) is better than right tail with similar smaller PRIs (red line)
O2: Left tail is better than right tail. The second observation is that if we cannot make the PRIs similar to each other as adding too much noise will make the output useless, it is better only to have left tail while making the larger PRIs similar than have right tail; because right tail means some individuals will have much higher risk than the rest, as shown in Figure 2. Even though the green line shows more individuals have relatively larger PRIs, their PRIs are similar enough so that they still receive similar protection; that is, they become indistinguishable with similar privacy loss.
These two key observations lead to a simple strategy for determining whether the PRIs provide adequate privacy protection. Based on O1, we can compute the ratio of the maximum PRI over the minimum PRI. Intuitively, as the ratio approaches 1, the PRIs become more similar and closer to the blue line shown in Figure 2. However, in many cases, this requires adding too much noise. Therefore, O2 allows the controller to trade privacy for more accuracy: instead of choosing the minimum PRI, the controller can choose the PRI at a specific percentile $p$. We fix the maximum PRI as the numerator to ensure the riskiest individuals will always be protected. The smaller $p$ is, the larger the ratio; therefore, more noise must be added (smaller $\epsilon$) to bring the ratio closer to 1. Although $p$ is an additional parameter the controller needs to set, it is more interpretable than $\epsilon$, since the controller only needs to decide at which point all the larger PRIs should be similar. For example, if the controller chooses the median (the 50th percentile), a ratio close to 1 indicates no individual is at considerably higher risk than the median individual.
Algorithm. Since PRI and $\epsilon$ are connected in a principled way, we can find a suitable $\epsilon$ that is consistent with the controller’s privacy requirement by computing the PRIs for different $\epsilon$, and choose the $\epsilon$ that is associated with PRIs such that the ratio between the maximum PRI and the PRI at the $p$th percentile is close to 1. The main intuition is to test different values of $\epsilon$ in descending order, and find the largest suitable $\epsilon$ that satisfies the controller’s privacy requirement while preserving as much accuracy as possible (note Figure 1).
Input : Dataset $x=(x_{1},\cdots,x_{n})$; query $q$; DP mechanism $\mathcal{M}$; consumed privacy budget $\epsilon_{c}$; percentile $p$; threshold $\tau$
Output : Chosen privacy parameter $\epsilon$, query output $r$
1 /* Initialize set of $\epsilon$ values to test; minimum value is $\epsilon_{c}$/
2 $\mathcal{E}\leftarrow$ initialize-eps-to-test($\epsilon_{c}$)
3 $x^{p}\leftarrow\textsc{project-query-attrs}(x,q)$
4 /* Compute $q(x_{-i})$ for all $i$ using memoization */
5 $C\leftarrow map()$, $L\leftarrow list()$
6 for $x^{p}_{i}\in x^{p}$ do
7       if $x^{p}_{i}\notin C$ then
8             $C[x^{p}_{i}]\leftarrow q(x^{p}_{-i})$
9      $L.add(C[x^{p}_{i}])$
10for $\epsilon\in\mathcal{E}$ do
11       /* Compute query output using DP with current $\epsilon$/
12       $r\leftarrow\mathcal{M}(x,q,\epsilon)$
13       /* Compute PRI for each individual */
14       $PRI\leftarrow list()$
15       for $l\in L$ do
16             $PRI_{i}\leftarrow\lVert r-l)\rVert_{1}$
17             $PRI.add(PRI_{i})$
18      $PRI_{max},PRI_{p}\leftarrow\max(PRI)$, get-percentile$(PRI,p)$
19       /* A suitable $\epsilon$ is found when the ratio is close to 1 */
20       if $PRI_{max}/PRI_{p}<1+\tau$ then
21             $\epsilon_{c}\leftarrow\epsilon_{c}+\epsilon$
22             return $r$
23
Algorithm 1 Find-$\epsilon$-from-PRI
Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output. First, the algorithm initializes the set of $\epsilon$ to choose from (line 1). Although the number of possible $\epsilon$ can be infinite, to ensure the algorithm runs in bounded time, we set a range (e.g., from $\epsilon_{c}$ to 10) and include representative $\epsilon$ value in fixed intervals (e.g., 0.1). The lower bound $\epsilon_{c}$ starts at 0 and updates for each answered query (we explain how to handle multiple queries in Section 3.3). The algorithm then projects relevant query attributes from the dataset (line 1) and computes the query output for each neighboring dataset (line 1 - 1). Next, the algorithm iterates over the list of $\epsilon$ from largest to smallest (line 1) and stops when it finds the first suitable value. For each $\epsilon$ it tests, the algorithm computes the current query output $r$ by applying the given DP mechanism $\mathcal{M}$ with parameter $\epsilon$ (line 1), and use it to compute the corresponding PRI for each individual (line 1 - 1). Finally, the algorithm computes the ratio of the maximum PRI over the PRI at the given $p$th percentile (line 1). If the ratio is close to 1 within the threshold $\tau$ (we use 0.01 in our implementation), then a suitable $\epsilon$ has been found. The algorithm updates the lower bound $\epsilon_{c}$ and returns the DP output $r$ (line 1 - 1).
Algorithm data flow. The algorithm is used by controllers to find $\epsilon$ from the PRIs and release DP output. The controller runs the algorithm on their input dataset and query. The algorithm first computes the PRIs, finds a suitable $\epsilon$ based on the controller’s privacy requirement, and returns the DP output computed using this $\epsilon$. Nothing is released by default to analysts. The PRIs are a by-product of which the sole purpose is to help the controller find $\epsilon$, and they are never released to the data analysts (we will discuss how such promise is enforced in Section 4.3. The $\epsilon$ value may need to be released on certain occasions; since finding $\epsilon$ is a post-processing step to computing PRI, which is DP, its release is also DP.
DP guarantee. The algorithm satisfies $\epsilon$-differential privacy. The algorithm only returns the query output $r$ when a suitable $\epsilon$ can be found. The output $r$ is computed using a standard $\epsilon$-differentially private mechanism (i.e., Laplace Mechanism); thus, the algorithm shares the same DP guarantee. This is no different than a standard DP deployment, only that $\epsilon$ is chosen automatically instead of manually by the controller.
3.3 Bounding Privacy Loss Across Queries
The challenge of running multiple queries. In practice, analysts often want to run multiple queries. Combining these queries’ output will reveal new information, eventually breaking privacy as per the fundamental Law of Information Recovery [22]. In today’s deployment of DP, the data controller needs to set a total privacy loss budget a priori on the dataset and stop accepting queries when the cumulative $\epsilon$ exceeds the budget. While the total budget provides a rigorous bound of privacy loss in the context of DP, it abstracts away from the actual privacy outcomes, and setting it without context is extremely challenging [23]. To further exacerbate the problem, the data analyst needs to decide the queries they want to run a priori since they need to decide how to utilize the total budget for their queries. This poses a major obstacle for analysts who do not have a pre-determined set of queries.
The solution. Instead of quantifying the maximum acceptable privacy loss, which is hard, we keep track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ of all previously answered queries; we call it $\epsilon_{c}$. For an incoming query, when using Algorithm 1 to find $\epsilon$, we truncate the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1). We reject answering this query if we fail to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$, meaning the PRIs cannot offer adequate privacy protection for any possible $\epsilon$.
Our solution creates a dynamic bound of total privacy loss. As the cumulative privacy loss, $\epsilon_{c}$, keeps growing for each answered query, it becomes more difficult to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$ since less noise can be added to make the PRIs meet the controller’s requirement. At each point that we fail to find a suitable $\epsilon$ for a query, the total privacy loss is bounded. Using DP’s sequential composition property, releasing the output of those answered queries satisfies $\epsilon_{c}$-differential privacy. The bound is refreshed if a suitable $\epsilon$ is found for a different query. This resulting dynamic bound helps controllers bound the total privacy loss without the need to set the total budget a priori. Since the algorithm automatically finds $\epsilon$ (or rejects queries) based on this dynamic bound, it also removes the need for the analyst to decide how to distribute their privacy budget across different queries a priori.
3.4 Discussion and Extensions
Algorithm complexity and optimizations. The algorithm has a worst-case time complexity of $O(nm)$, where $n$ is the number of records in the dataset, and $m$ is the number of $\epsilon$ values we test. Computing the PRI of each individual requires computing the query output of each neighboring dataset (line 1 - 1). When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$. In addition, the query computation can be parallelized since each computation is independent of the other. We implement multi-processing and compute the query output for a chunk of records in each process, and we show in Section 5.3.1 that the algorithm is both efficient and scalable.
Expressing privacy preferences. A modification of our algorithm permits the controller to specify protection preferences over certain groups of individuals whose privacy should be protected similarly. For example, they can specify individuals whose ages are above 60 and those younger need to receive the same privacy protection. This feature can easily be incorporated into the algorithm by letting the controller change the ratio to compute. In the example, the controller can choose to compare the median PRI of the two age groups and choose the $\epsilon$ that makes the ratio of the two medians close to 1.
Communicating Privacy-Accuracy Trade-off. In some circumstances, the data controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual. Prior work such as DPComp[28] and ViP[40] have explored different ways to visualize privacy-accuracy trade-off in differential privacy. The PRIs can complement the existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
4 Helping Analysts Interpret Utility
In this section, we first propose a simple utility signaling protocol to help analysts understand the impact of DP on their downstream applications to address Challenge 2 (Section 4.1). Then, we assemble a set of requirements for a DP platform, which we derive from the solutions proposed (Section 4.2). In Section 4.3, we introduce a new implementation of our new algorithm and protocol on top of a data escrow that meets those requirements.
4.1 The utility signaling protocol
Today’s DP deployments. In the central DP model, we differentiate among individuals who own the original data, controller that manages individuals’ data on their behalf, platform the infrastructure where computation takes place, and analyst the actor demanding computation over individuals’ data. The analyst submit queries to a platform operated by a controller. The platform computes query output using DP and releases them to the analyst with the controller’s approval.
We consider the case where the analyst has a downstream task with an associated evaluation function, g, that returns the utility of the downstream task. The function takes the set of query output, S, that the analyst received as input. Following the notation used in Section 2.2), $S=\{\mathcal{M}(x,q,\epsilon)\mid x\in\mathcal{X}^{n}\}$, $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$. For example, an analyst queries census data at the county level (i.e., the US Census Bureau is the controller and citizens are the individuals). The set of output $S$ are aggregate statistics on population count, education level, housing prices, etc. The analyst’s task is to build a machine learning model to predict the median income of each county based on those statistics, and the evaluation function returns the prediction accuracy, which is g(S). Since the queries are computed using DP, releasing $g(S)$ is also DP by post-processing.
The problem. The controller uses DP to release noisy output to protect the privacy of individuals, but the analyst does not know how much such noise affects their downstream task and whether the output is useful. In the example, an analyst who obtains a prediction accuracy of $88\%$ does not know how much the accuracy would have been without noise; thus, they cannot decide whether it is worth investing more effort in improving the accuracy of a model. Clearly, the platform cannot tell the analyst how much noise was added because that would jeopardize the privacy guarantee.
A proposed solution: a utility signaling protocol. We propose a protocol to help analysts interpret the implications of the added noise on the quality of their downstream task by letting them compare to the quality they would have obtained over the ground truth output. Let $S^{\prime}=\{q(x)\mid x\in\mathcal{X}^{n}\}$ be the set of ground truth output (i.e., without applying DP). The protocol allows the analyst to obtain both g(S) and g(S’), thus signaling the utility loss due to DP.
Sending the ground truth output to the analyst so they run their downstream task is not feasible as it breaks the premise of protecting the data in the first place. Instead, the analyst sends their downstream task to the platform. The platform invokes the evaluation function $g$ over both the set of noisy output $S$ and the set of ground truth output $S^{\prime}$. Then, if the controller approves, both $g(S)$ and $g(S^{\prime})$ are released to the analyst to signal the loss in utility due to DP. In the example, the platform trains a model over both the DP census statistics and the ground truth statistics, then applies the evaluation function the analyst shared, obtaining $88\%$ and $90\%$ accuracy, respectively. Upon receipt of both numbers, the analyst understands the utility effect of DP and can decide whether it is worth investing more effort in improving the model.
Assessing Privacy Implications of Releasing $g(S^{\prime})$: Any information release may leak information about the individuals’ data, e.g., even the accuracy of a model (consider the trivial case where $S^{\prime}$ contains only one value, the analyst can simply use the evaluation function to test different values until they obtain an accuracy of 100%). Since DP is not being applied when releasing $g(S^{\prime})$, it is ultimately at the discretion of the controller to decide whether or not to release $g(S^{\prime})$. In the example, the controller may be comfortable releasing the accuracy of the machine learning model trained on the ground truth census statistics (in addition to the model accuracy on the DP statistics), but not releasing the model itself. Our utility signaling protocol does not attempt to address the possibility of various attacks the analyst may perform to infer individual data from $g(S^{\prime})$. Instead, the escrow-based platform we describe next allows the controller to examine the analyst’s downstream task and associated evaluation function and ensures $g(S^{\prime})$ will not be released without the controller’s permission. For many applications, the benefit of allowing the analyst to observe $g(S^{\prime})$ can be tremendous.
4.2 Platform Requirements
We propose a new platform to support the Find-$\epsilon$-from-PRI algorithm that helps controllers (addressing Challenge 1) and the utility signaling protocol that helps analysts (addressing Challenge 2). Such a platform must meet the following requirements:
  • R1. Control Data Flow: Analysts should never see the raw data or PRIs. And they should not submit queries and downstream tasks over the raw data directly without the controller’s approval (e.g., the controller may approve releasing the model accuracy to signal utility loss).
  • R2. Trust: The new protocol requires more trust on the platform. Trust is a broad concept, and its meaning varies depending on context[7]. In the context of DP, trust means the platform adheres to prespecified behavior and the DP mechanism upholds its privacy guarantee[7][43]. Individuals must trust the controller will protect their data in a responsible manner (i.e., using DP). Controllers must trust the platform keeps the data secure and runs computations on the data correctly. With the new utility signaling protocol, analysts must also trust the platform runs their downstream task correctly. We want to reduce the trust those entities have on the platform.
  • R3. High Performance: The platform must limit the overhead, which is at odds with the pressures of the new Find-$\epsilon$-from-PRI algorithm that helps controllers at the cost of higher computational overhead.
We highlight a tension between R2 and R3. Improving trust often requires employing techniques (e.g., cryptographic primitives) that increase the computational overhead, potentially leading to impractical solutions. We now explain the design of a platform that addresses these three requirements.
4.3 Escrow-Based DP Platform
We propose a new platform, based on an escrow architecture similar to the proposed in  [50], that lets controllers maintain fine-grained control of data flow (meeting R1), reduces the trust on the controller and platform by both individuals and analysts (meeting R2) all while maintaining high performance to meet R3.
4.3.1 Controlling Data Flow to Implement the utility signaling protocol
A data escrow supports delegated computation, which means the escrow runs computation on behalf of the computation owner. Implementing a platform with an escrow lets analysts delegate their downstream tasks to the platform, thus obtaining the two utilities $g(S)$ and $g(S^{\prime})$. We present the end-to-end protocol in Figure 3, which consists of seven steps.
Figure 3: End-to-end utility signaling protocol. We illustrate the protocol in the central DP deployment scenario where the controller has access to the raw data.
1. Registering data. The controller registers the dataset containing individual records with the escrow.
2. Registering application. The controller registers an application that uses DP to run queries on the registered dataset, which includes the Find-$\epsilon$-from-PRI algorithm. Any derived data product, such as the PRI, will not be released without the controller’s permission. The PRI can also be computed separately as an independent function inside the application if the controller wants to use it for other purposes.
3. Creating policy. The data controller creates a policy to let the analyst run queries on the registered dataset using the registered application; a policy in the escrow indicates who can run what computation on what data.
4. Issuing Queries. The analyst issues queries to the application. The application answers each query and stops when a suitable $\epsilon$ can no longer be found. After checking that the query computation only ran on data accessible to the analyst, the escrow releases the set of DP output $S$ to the analyst because the controller wrote the corresponding policy. The set of ground truth output $S^{\prime}$ is registered inside the escrow since it is a derived data product of the computation.
Up to this point, the protocol provides analysts with the same information they can access in a typical central DP deployment. Now, we extend the protocol to implement the utility signaling component to help analysts understand the impact of DP on their downstream tasks.
5. Registering downstream task. The analyst registers their downstream task and an associated evaluation function $g$ as an application in the escrow. The evaluation function is the only function within the application that is available for the escrow to invoke and anyone to create policy on; if exists, the rest of the functions should be invoked from the evaluation function to perform other necessary computations. The escrow ensures that only the evaluation function’s output (and nothing else) will be released to the analyst.
6. Creating policy for downstream task. After inspecting the application submitted by the analyst and determining it is acceptable to release $g(S^{\prime})$, the controller creates a policy to run the evaluation function $g$ on the ground truth output $S^{\prime}$.
7. Evaluating downstream task. With the controller’s permission, the analyst obtains both $g(S)$ and $g(S^{\prime})$. By inspecting the difference between the two utilities, the analyst learns the impact of DP on their task, thus helping them decide whether the DP output is sufficiently good for their task.
By exploiting the delegated computation features of the data escrow, the platform controls data flow, which is necessary to implement the utility signaling protocol.
4.3.2 From Trusted to Trustworthy, while Fast
A trusted platform adheres to prespecified behaviors[7], in the context of DP this means an adversary only has access to the output of a DP mechanism, but not any of its internal state and any raw or derived data[22]. This is infeasible to achieve in practice. A central platform, as in the central DP model, can claim to be trusted, but such claim does not preclude it from being compromised and intentionally leaking data in the clear, thus breaking the DP guarantee[43]. A trustworthy platform, on the other hand, implements different techniques to enforce the promise that not only will it never behave unexpectedly, it will not pose harm (to the users) in any way[27]. An escrow-based platform reduces the impractical assumptions of a trusted platform and supports trustworthy computation by ensuring:
  • No data or derived data product (e.g., query output) ever leaves the platform without explicit permission from the controller, and only explicitly allowed computation runs on existing data; this is the basis of the control of data flow.
  • Data is encrypted end-to-end, even during computation. End-to-end encryption can be achieved in many ways; for instance, DP platforms leveraging secure multiparty computation[42][9][10][49][45], homomorphic encryption[4][46], Oblivious RAM[12] and trusted execution environment[44]. The prototype we use for our evaluation uses AMD Secure Encrypted Virtualization(SEV)[33].
  • Tasks that are sent to the escrow are the tasks that run. The escrow is registered with an escrow application; the platform ensures that the correct version of the application is running using techniques such as remote attestation[15] or formal verification[38][35]. Our prototype uses hardware-rooted remote attestation enabled by AMD SEV.
Figure 4: Left: central platform. Right: escrow-based platform
By computationally enforcing the promises above, the escrow reduces the trust that both individuals and analysts must place in the platform while maintaining the same accuracy guarantee as the central DP model. Furthermore, the escrow application can be written so that controllers do not gain access to the raw data but only the PRIs (i.e., by writing a policy that only reveals the PRIs to the controller), thus illustrating multiple escrow-based deployment scenarios. Consider Figure 4, which shows how individuals, controller, and analyst interact with the central platform in CDP and the escrow-based platform. In CDP, the controller manages individuals’ data using the central platform and sets $\epsilon$ manually; the individuals need to trust both the controller to protect their privacy (i.e., by choosing a good $\epsilon$) and the platform to operate correctly without leaking their data. Whereas using the escrow-based platform, individuals send data directly to the platform, and the controller only has access to PRIs computed by Algorithm 1. The escrow invokes Algorithm 1 to find $\epsilon$ on the controller’s behalf and releases the output to the analyst, who delegates their downstream task to the escrow to obtain utility.
5 Evaluation
Our evaluation shows our solutions address both Challenge 1 to help controllers, and Challenge 2 to help analysts. We answer the following research questions for our evaluation:
  • RQ1: Does presenting the privacy risk indicator help data controllers choose $\epsilon$?
[…]
  • Microbenchmarks: We conduct experiments to show the
    Find-$\epsilon$-from-PRI algorithm is efficient and scalable. In addition, we study how sensitive is our algorithm to different values of parameter $p$.
5.1 RQ1: Using PRI To Choose Epsilon
To determine whether the PRI help the data controllers choose $\epsilon$, we conducted an IRB-approved user study to analyze how PRI affects the controller’s decision-making regarding choosing $\epsilon$. Our goal is to show that by presenting the PRI to the participants, they will be more consistent on the $\epsilon$ they choose compared to if they don’t have the PRI. To this end, we conducted a between-subject user study where the participants are divided into control and treatment groups. To prevent learning effects, only the treatment group sees the PRI. We chose to conduct the study in the form of an online survey, because our tasks do not require in-person guidance and it is more convenient for participants to complete the study. In addition, it allows us to collect data from a more representative sample.
5.1.1 Recruitment and data collection
We recruited 100 participants using Prolific[1]. To complete the study, each participant must sign a consent form, which describes the purpose and content of the study, compensation, and data confidentiality. We do not collect identifiable information, and the survey responses are anonymized, meaning the participant’s IP Address, location data, and contact information are not being recorded.
Prescreening criteria. When recruiting on Prolific, we prescreen participants based on three criteria: i) they are fluent in English, to avoid confusion due to misunderstanding language; ii) they have completed an undergraduate degree or above; iii) they study Computer Science, Computing (IT) or Mathematics; this is to ensure that the participants have a sufficient quantitative background to complete the study.
Tutorial and Screening Quiz. Since the main goal of the study is to ask the participant to choose $\epsilon$ when using DP to release query output, the participant needs to have basic knowledge about DP and understand the implications of the $\epsilon$ they chose. To ensure each participant knows enough about DP to complete the study, we ask each participant to read a short tutorial about DP at the beginning. This tutorial explains the intuition of DP and the effect of the parameter $\epsilon$ at a high level. Mainly, it explains DP is generally achieved by adding random noise to the query output, and $\epsilon$ is inversely proportional to the amount of noise added.
After reading the tutorial, the participant needs to complete a screening quiz that tests their knowledge about the tutorial they just read. Only participants who pass the quiz proceed to conduct the study. Out of the 100 participants we recruited, 56 of them passed the quiz and their responses are used for the study analysis. This quiz ensures that participants have understood the principles of DP and thus ensures the validity of the user study.
5.1.2 Study procedure
First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group and 27 participants are in the treatment group. The survey presents a sample from the UCI Adult dataset[2] and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$. Then, it presents two queries in random order and asks the participant to choose an $\epsilon$ value from a slider, with values from 0.5 to 5.0 and each interval being 0.5. The participant’s task is to choose an $\epsilon$ value that offers equal privacy protection to all people in the dataset while preserving as much accuracy as possible.
For the control group, the participant is presented with a table that shows the range of possible output for each $\epsilon$, indicating output accuracy, along with the ground truth (non-DP) output. The table includes information that any controller would have access to in a deployment today when choosing $\epsilon$.
For the treatment group, the participant is presented with the same table, but with an additional column that includes a plot showing the range of PRIs for each $\epsilon$ (e.g., a box drawn from the minimum to maximum PRI). Note that this is the only change between the control and treatment groups to ensure we isolate the intervention. We explain beforehand what the PRI is at a high level, and that a smaller range represents more similar PRIs among people in the dataset. A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include these to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
5.1.3 Results
Figure 5: $\epsilon$ chosen by each participant
We compare the distribution of $\epsilon$ chosen by participants in the control and treatment groups. Figure 5 shows a swarm plot where each point represents the $\epsilon$ value chosen by each participant in the control and treatment group for each query. For both queries, 19 out of 27 participants in the treatment group chose either 2 or 1.5, showing a more consistent answer than the control group. This is because the PRI range when $\epsilon$ is 2 or 1.5 is considerably smaller than that of larger $\epsilon$, and smaller $\epsilon$ has lower accuracy. This shows that most participants could connect the PRIs and their task goal, which is to offer equal privacy protection while ensuring high accuracy, and they were able to use the PRIs to choose the appropriate $\epsilon$ that satisfies the goal. On the contrary, the $\epsilon$ values chosen by participants in the control group are more spread out and have greater variance, suggesting participants shared no general consensus as to what is a good $\epsilon$ to pick.
In addition, we ran the Mann–Whitney test [37] on the $\epsilon$ values in the control and treatment group, with the alternative hypothesis being the two samples come from different distributions. For both queries, we obtained $p<0.005$. Thus we reject the null hypothesis and conclude that showing the PRIs helped participants choose $\epsilon$.
In the follow-up question for the treatment group, all 27 participants answered yes when asked whether they used the PRIs when choosing $\epsilon$, confirming our hypothesis that presenting the PRIs helped participants choose the $\epsilon$ that satisfies the task goal.
Figure 6: $\epsilon$ chosen by participants who do not know DP (left) and those who know DP (right)
Does knowing DP matter? Another interesting question is whether the participant’s a priori familiarity with DP affects how they choose $\epsilon$. At the start of the survey before showing the tutorial, we asked each participant whether they know about DP. 27 participants (out of a total of 56) answered they don’t know anything about DP and 29 participants answered they know about DP.
Figure 6 shows the $\epsilon$ chosen by participants based on their familiarity. We find similar patterns. Participants in the treatment group tend to choose either 2 or 1.5, regardless of their familiarity with DP, whereas participants in the control group chose $\epsilon$ with greater variance. This suggests participant’s a priori knowledge of DP does not transfer to making better choices of $\epsilon$; on the other hand, showing PRIs helped participants make more consistent choices whether or not they know about DP previously.
How do different ways of presenting PRIs affect participants’ decisions when choosing $\epsilon$? Since this user study relies on the participants inspecting and interpreting the PRIs with their own eyes to choose $\epsilon$, instead of using the algorithm to automatically find one, how we present the PRIs to the participants can greatly affect their choices. We explored an alternative way of presenting PRIs - by using histograms. To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and asked them to complete the same survey questions in the treatment group, with the only change being presenting histograms of PRIs instead of range plots.
Figure 7: $\epsilon$ chosen by participants who were presented with histograms or range plots, or in the control group
Figure 7 shows $\epsilon$ chosen by participants who were presented with histograms or range plots, and we include the $\epsilon$ chosen by participants in the previous control group as reference. Participants who saw histograms did not perform better than participants who saw range plots. Especially for Q2, 8 participants chose larger $\epsilon$ (3.5, 4, and 4.5). Although we cannot be certain since the responses are anonymous, one possible explanation might be that for the large $\epsilon$, the majority of PRI are already small and similar to each other. Only a small portion of PRI have large values. The participants might think that as long as the privacy of the majority of people is being protected, the task goal is achieved; hence, they chose larger $\epsilon$, so the results are more accurate. However, if the participants only saw the range of PRI, they would be more inclined to choose smaller $\epsilon$ with a smaller range of PRI. Another possible explanation could be that some participants had difficulties interpreting histograms, whereas the range plots are generally much easier to understand.
The discrepancy of results by tweaking how we present the PRIs demonstrates how we convey information to the users affects their ability and reasoning for choosing $\epsilon$. Exploring other alternative solutions for presenting and helping users interpret PRI remains an interesting problem.
5.1.4 Threats to validity
External validity. We recruited participants from diverse backgrounds across the world via Prolific, with few prescreen criteria such that the participants can complete the study without difficulty.
Internal validity. To ensure all participants have the minimum knowledge of DP needed to complete the study, each participant must read a tutorial and pass a quiz. We only allowed the participant to take the quiz once to minimize the chance of participants abusing it if they got multiple chances. To avoid learning effects, we used a between-subject study to isolate the condition we intended to test. We also randomized the two questions for choosing $\epsilon$.
5.2 RQ2: DP On Escrow-based Platform
[…]
5.2.1 Baselines for comparison
We evaluate and compare the performance of the four following baselines:
  • Data Escrow: We implemented a private query escrow application that takes a SQL query as input and applies the Laplace Mechanism to release the query output. Optionally, if the parameter $\epsilon$ is not given, it will use Algorithm 1 to find $\epsilon$ automatically. We simulated the protocol where the data controller registers their dataset to the escrow, creates a policy that allows the data analyst to issue queries to their data using the private query application, and the analyst issues their query and receives the query output.
[…]
  • Standalone without DP: This baseline simply runs the given query without DP. The dataset is in plaintext.
  • Standalone with DP: This baseline represents the central DP model. It runs the given query and applies the Laplace Mechanism to the query output. The dataset is in plaintext.
[…]
5.2.2 Setup
We ran the experiments on n2d-standard-8 instances with 8 virtual CPUs (AMD Rome) and 64GB memory from Google Compute Engine. We used AMD’s Secure Encrypted Virtualization (SEV) when running the data escrow. We used Ubuntu 20.04 and Python 3.8 for all baselines.
5.2.3 Dataset
[…]
5.2.4 Queries
We run the 3 queries independently for all baselines:
  • Q1: Number of people between age 20 and 30 (simple predicate with range)
  • Q2: Number of people by each race (aggregate & group-by)
  • Q3: Number of people who are female and have income greater than 50K (two simple predicates)
5.2.5 Results
[…]
Figure 8: Average query time over 10 runs
[…]
5.3 Microbenchmarks
We evaluate the runtime and scalability of the Find-$\epsilon$-from-PRI algorithm in Section 5.3.1 and study how different values of the parameter $p$ affect the $\epsilon$ chosen in Section 5.3.2.
5.3.1 Is Find-$\epsilon$-from-PRI practical?
Since the algorithm can run multiple times if there are multiple queries , the algorithm needs to be efficient so running it multiple times will still be practical. We ran the algorithm as a registered application in the escrow.
Datasets. We controlled the experiment by generating datasets of size 1000, 10000, 100000, and 1 million based on the Adult dataset. We sampled records for size datasets of size 1000 and 10000. For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they are not the same copy of each other. We included all attributes and all records in the dataset.
Queries. We ran the algorithm to find $\epsilon$ for the five SQL queries below. We chose a diverse set of queries with attributes of different domain sizes and types. The queries are run independently.
  • Q1: SELECT COUNT(*) FROM adult WHERE income==’>50K’ AND education_num=13 AND age==25
  • Q2: SELECT marital_status, COUNT(*) FROM adult WHERE race==’Asian-Pac-Islander’ AND age>=30 AND age<=40 GROUP BY marital_status
  • Q3: SELECT COUNT(*) FROM adult WHERE native_country!=’United-States’ AND sex==’Female’
  • Q4: SELECT AVG(hours_per_week) FROM adult WHERE workclass==’Federal-gov’ OR workclass==’Local-gov’ or workclass==’State-gov’
  • Q5: SELECT SUM(fnlwgt) FROM adult WHERE capital_gain>0 AND income==’<=50K’ AND occupation==’Sales’
Parameters. We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,0.099,\cdots,0.001]$ (37 values in total). We set $\epsilon_{c}=0$, $p=50$ and $\tau=0.01$. We used six parallel processes for computing the PRIs.
Figure 9: Average runtime to find $\epsilon$ over 10 runs
Results. We first note that all queries on all 10 runs were able to find a suitable $\epsilon$ on every dataset we tested. Figure 9 shows a line plot where the x-axis is the data size and the y-axis is the runtime, both in log scale. Each line represents a query, with the error band showing the standard deviation.
As shown in the figure, the runtime increases linearly with the data size. Most queries took less than a few minutes to finish, even for 1 million records, except Q5, which took more than an hour. The runtime for Q5 is slower since it involves the attribute fnlwgt, which is a numerical attribute with a large domain size and many unique values, thus the query involves significantly more unique records than the other queries. For example, for 1 million records, Q5 has approximately 100000 unique records of which the PRIs need to be computed; whereas for other queries, the algorithm only needs to compute PRIs for no more than a few thousand records and the rest can be directly retrieved using memoization.
5.3.2 Effect of parameter $p$
To find a suitable $\epsilon$, the algorithm computes the ratio of the maximum PRI over the PRI at the $p$th percentile; if the ratio is close to 1, the PRIs are considered to reflect adequate privacy protection by the controller’s privacy requirement. Specifically, we are interested in choosing the 00th (minimum), $25$th, and $50$th (medium) percentile. The smaller $p$ is, the stricter the privacy requirement and the harder it is to find $\epsilon$, meaning more noise needs to be added to achieve a ratio closer to 1.
Setup. We used the Adult dataset with 100000 rows that we generated in Section 5.3.1, and we used the same settings (except $p$). We ran the five queries in Section 5.3.1 10 times and we report the $\epsilon$ selected by our algorithm in each run.
Figure 10: $\epsilon$ found when using the PRI at the 0, 25, and 50th percentile to compare against the maximum PRI
Results. Figure 10 shows the $\epsilon$ found when $p=0,25,50$. As expected, using higher percentiles tends to find larger $\epsilon$ than lower percentiles, because less noise needs to be added to bring the maximum PRI closer to the PRI at the larger percentile. When using the minimum PRI, Q1 – Q4 find very small $\epsilon$ that are between 0 and 1. Q1 and Q3, both counting queries, are able to find larger $\epsilon$ since the queries have a global sensitivity of 1, i.e., the maximum difference between PRIs will also be fixed at 1, it becomes easier to make the ratio closer to 1. Whereas Q2 and Q4, a group-by query and an average query with higher sensitivities, require more noise to be added to bring the PRIs closer. For Q5, using the minimum PRI only finds $\epsilon$ once, since the sensitivity is too large, primarily due to the domain size of fnlwgt), the DP mechanism will not be able to add enough noise of such scale beyond certain $\epsilon$ (when $\epsilon=0.1$, the noise scale will be larger than 10,000,000, which is the threshold we use to determine whether the DP mechanism can be applied). However, when the requirement is relaxed, it is able to find larger $\epsilon$ such that the noise scale does not exceed the threshold.
6 Related Work
Choosing privacy loss budget. Different approaches [30] [26] [23] [36] [39] [40] [28] [48] [32] [28] aim to help users choose $\epsilon$ when using DP. One line of work [39] [40] [28] [48] [32] implements an interface to visualize the effect of different variables (e.g., utility metrics) when adjusting $\epsilon$. Nanayakkara et al. visualizes the disclosure risk in relation to $\epsilon$; they use the definition of disclosure risk from Lee and Clifton, which is the probability of disclosing the presence or absence of any individual in the database. John et al. defines and visualizes the data sharing risk; Both metrics quantify potential data leakage, but neither quantifies leakage with respect to each individual in the data, unlike our PRI. Ge et al. translates a user-specified accuracy bound to a suitable DP mechanism, along with an $\epsilon$ that incurs the least privacy loss while satisfying the accuracy bound. Hsu et al. proposes an economic model for choosing $\epsilon$ by balancing the interests of two different parties: the data analyst and the prospective participant who will choose to participate if the benefit outweighs the risk. However, this model assumes that the participant can quantitatively express the "cost" of whether to participate in the study, which is impractical for non-experts. In addition, Cummings et al. and Xiong et al. investigate user’s understanding of DP guarantees through multiple user studies, which offer insights on facilitating communicating the privacy implications of DP to users and help them choose $\epsilon$. Jayaraman and Evans evaluates the impact of $\epsilon$ on ML tasks and specifically their privacy-utility tradeoffs. To the best of our knowledge, no known approach helps users understand the implications of $\epsilon$ on individuals’ privacy, and use that to design an algorithm that finds $\epsilon$ automatically.
Interpreting utility. Tao et al. proposes DPXPlain to explain aggregate query answers under DP by comparing answers with and without DP, and assessing whether the difference is due to the noise introduced by DP or properties of the data, thus helping analysts interpret whether the answers are useful for their analysis. This idea, complementary to ours, relies on analysts to come up with comparison questions that require explanation in order to understand any anomalies or trends in the query answers. Moreover, feasibility studies [8][14] have been conducted to evaluate real-world usage of DP.
CDP LDP Shuffler
Crypto
Primitive
Data
Escrow
T?
No Yes Yes Yes
Yes
Add.
Trust
Asm.
Trust
indv’ls
add noise
correctly
Trust
shuffler
Trust
crypto
impl.
Trust
SEV &
escrow
impl.
Encr.
Data?
No
Does not
collect
raw data
Does not
collect
raw data
Yes Yes
Error $O(1)$ $O(\sqrt{n}$) Near $O(1)$ $O(1)$ $O(1)$
Table 1: Comparisons of different DP deployment scenarios. T stands for trustworthy.
Deploying differential privacy. Alternatives beyond the central model aim to deploy DP with fewer trust assumptions. We summarize those different deployment scenarios in Table 1. The local model (LDP) has been proposed to eliminate the need for a trusted data curator and it has been adopted in various real-world applications for collecting client data from their software and devices [24] [18] [17]. In LDP, each individual perturbs their own data using LDP mechanisms before sending it to the data curator; the data curator does not need to be trusted since it only sees the differentially private version of data. However, the primary disadvantage of the local DP model is that it introduces a much larger error than the central DP model. The shuffle model creates an anonymous communication channel that amplifies the privacy guarantees [34] [6], thus it requires less noise to be added when users perturb their data and offers better accuracy than the local DP model. However, it requires additional trust assumptions as the shuffler must be trusted. Lastly, techniques that leverage cryptographic primitives, specifically secure multiparty computation, homomorphic encryption, oblivious RAM, and trusted hardware enclaves, have been proposed to offer security guarantees when deploying DP without a trusted data curator [20] [42] [9] [4] [45] [46] [49] [10] [44] [11] [12]. We showed that the escrow-based solution we introduce presents an interesting alternative with a lower computational cost.
7 Conclusions
To ease the deployment of DP, we address challenges faced by controllers and analysts by 1) defining the privacy risk indicator (PRI) that indicates the relative privacy loss of each individual for a given $\epsilon$ and output and an algorithm to choose $\epsilon$ using the PRIs. 2) introducing an utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks. We propose a platform implementation on top of a data escrow. We demonstrated our contributions through an IRB-approved user study and empirical evaluation with other DP platforms.

### v2

@@
 Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform
-Abstract
+Abstract.
 Differential privacy (DP) enables private data analysis but is hard to use in practice.
@@
 For data analysts who submit queries, it is hard to understand the impact of the noise introduced by DP on their tasks.
-To address these two challenges: 1) we define a privacy risk indicator that indicates the impact of choosing $\epsilon$ on individuals’ privacy and use that to design an algorithm that chooses $\epsilon$ automatically; 2) we introduce a utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks.
-We implement the algorithm and the protocol inside a new platform built on top of a data escrow, which allows the controller to control the data flow and achieve trustworthiness while maintaining high performance.
+To address these two challenges: 1) we define a privacy risk indicator that indicates the impact of choosing $\epsilon$ on individuals’ privacy and use that to design an algorithm to choose $\epsilon$ and release output based on controllers’ privacy preferences; 2) we introduce a utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks.
+We implement the algorithm and the protocol inside a new platform built on top of a data escrow, which allows controllers to control dataflows while maintaining high performance.
 We demonstrate our contributions through an IRB-approved user study, extensive experimental evaluations, and comparison with other DP platforms.
 All in all, our work contributes to making DP easier to use by lowering adoption barriers.
-1 Introduction
-Analyzing private data is essential for various tasks, and differential privacy (DP) is a promising tool to achieve privacy.
-In a typical DP deployment, data analyst runs queries on a dataset hosted on a platform that is accessible to a data controller.
+PVLDB Reference Format:
+PVLDB, 14(1): XXX-XXX, 2020.
+doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
+Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
+For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
+Copyright is held by the owner/author(s).
+Publication rights licensed to the VLDB Endowment.
+Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
+doi:XX.XX/XXX.XX
+1. Introduction
+Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006b) is a promising tool to analyze data about individuals, but it remains challenging to use in practice.
+In a typical DP deployment, a data analyst runs queries on a dataset hosted on a platform that is accessible to a data controller.
 The dataset contains information about individuals, and releasing this information risks revealing the identity of those individuals.
-The data controller wants to protect their identity while releasing useful output to the data analyst.
+The data controller wants to protect individuals’ privacy while releasing useful output to the data analyst.
 DP guarantees that the output of the query is likely to be the same if we change or remove any individual’s record in the dataset.
 This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: smaller $\epsilon$ leads to more accurate output but weaker privacy guarantees.
-Despite bespoke deployments by large organizations such as the US Census Bureau[3], Apple[5], and Google[24], it remains difficult to use DP in practice [23][43].
-In this paper, we consider two important challenges controllers and analysts face.
-Challenge 1.
-The controller’s perspective.
+Despite some successful deployments of DP (Abowd, 2018; Apple, 2017; Erlingsson et al., 2014; Ding et al., 2017), a few roadblocks prevent broader DP deployment.
+In this paper, we consider three such roadblocks:
+Challenge 1: Choosing $\epsilon$.
 Controllers need to decide the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$.
-Dwork et al.[23] showed that controllers have no clear consensus on how to choose $\epsilon$ or how to approach this problem in a principled manner.
+Dwork et al.
+showed that controllers have no clear consensus on how to choose $\epsilon$ or how to approach this problem in a principled manner.
 Our contribution.
@@
 In response, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual.
-Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a query output affects individuals’ privacy.
-And because the PRI and $\epsilon$ are connected in a principled way, there is an opportunity to automatically find an $\epsilon$ that produces a profile of privacy risk indicators consistent with the controller’s privacy requirement.
-We exploit this opportunity by designing an algorithm that automatically finds an $\epsilon$ and releases query output such that all individuals receive adequate protection based on the controller’s privacy requirement.
-Challenge 2.
-The analyst’s perspective.
-Analysts submit queries and receive noisy output that deviates from the ground truth output, i.e., the output if DP had not been applied.
-Without access to such ground truth, analysts do not know the impact of DP on the quality of their downstream tasks; this hinders the analysts’ ability to use the DP output in practice.
+Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a specific query output affects individuals’ privacy.
+And because PRI is connected to $\epsilon$, there is an opportunity to automatically find an $\epsilon$ that produces a profile of privacy risk indicators consistent with the controller’s privacy preferences; we exploit this opportunity by designing an algorithm, called Find-$\epsilon$-from-PRI, to choose $\epsilon$ and release DP output given a simple description of controllers’ privacy preferences.
+Challenge 2: Interpreting utility of DP output.
+Analysts submit queries and receive noisy output that deviates from the raw output, i.e., the ground truth output if DP had not been applied.
+Without access to such ground truth, analysts do not know the impact of DP on the quality of their downstream tasks, which is the reason why they sought the DP output in the first place.
+This hinders the analysts’ ability to use the DP output in practice.
 Our contribution.
-In this paper, we propose a utility signaling protocol to help analysts interpret the impact of DP on their downstream tasks while still allowing controllers to fully control the data flow.
+In this paper, we propose a utility signaling protocol to help analysts interpret the impact of DP on their downstream tasks while still allowing controllers to fully control data release.
 In the protocol we propose, the analyst delegates their downstream task with an associated evaluation function to the platform.
-With the controller’s permission, the platform invokes the evaluation function on both the DP output and the ground truth output and releases both results, thereby signaling the utility loss on the analyst’s downstream task due to noise introduced by DP.
-Putting everything together: Escrow-Based Platform.
-The utility signaling protocol requires analysts to place more trust in the platform than in a traditional DP deployment.
-We want to make the platform trustworthy while maintaining high performance, which is at odds with the increased overhead introduced by the algorithm we propose to help controllers.
-Our solution is to build a DP platform on top of a data escrow, which permits controlling data flow and improving trust with significantly lower overhead compared to other DP platforms.
-Summary of related work.
-Many approaches have been proposed to help controllers choose $\epsilon$ [30] [26] [23] [36] [39] [40] [28] [48] [32] [28], [26] [30] [36] [39] [40], however, none concentrates on communicating the effects of $\epsilon$ on specific individuals represented in the dataset, as we do.
-In addition, approaches for communicating the privacy guarantees offered by DP are mainly targeted toward end-users[41][25], i.e., the privacy risk when sharing their data, whereas our approach focuses on helping controllers, who can only choose $\epsilon$ if they understand its privacy implications on those end-users.
-Exposing error statistics, e.g., error bound, can help analysts interpret the output; such is the approach followed by the US Census [13].
-However, such statistics are not sufficient in many cases.
-Our utility signaling protocol also complements existing approaches [47] to help analysts gain a deeper understanding of the effect of DP on their output.
-Lastly, different deployment of DP has received a lot of attention [20] [42] [9] [4] [45] [46] [49] [10] [44] [11].
-Our escrow-based implementation hits a new point in the design space between the degree of trust required by participants and the resulting performance.
+With the controller’s permission, the platform invokes the evaluation function on both the DP output and the raw output and releases both results, thereby signaling the utility loss on the analyst’s downstream task due to noise introduced by DP.
+Challenge 3: Deploying our solutions in a DP platform.
+Deploying our solutions to the previous two challenges brings a new challenge on the platform design.
+Specifically, implementing the new Find-$\epsilon$-from-PRI algorithm (addressing Challenge 1) and utility signaling protocol (addressing Challenge 2) calls for a platform that meets the four requirements: (1) it allows controllers to control dataflows, (2) it implements necessary security measures to protect against unintended dataflows, (3) it facilitates controllers to inspect dataflows, and (4) it adds little overhead.
+Various DP platforms have been proposed, but none meet all four requirements.
+For instance, the central DP model assumes adversaries are incapable of corrupting the platform and forcing unintended dataflows, which is unrealistic in practice (Near et al., 2021); platforms leveraging cryptographic primitives (Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018) make fewer assumptions about the adversaries, but often have significant computation overhead.
+Our contribution.
+We propose a new platform built on top of a data escrow (Xia et al., 2022) which enables the implementation of the Find-$\epsilon$-from-PRI algorithm and utility signaling protocol, meeting the four platform requirements we curated.
+We show that the escrow-based platform hits a new point in the design space, offering a better trade-off between security, usability, and performance.
 […]
+Summary of Contributions and Related Work
+Helping controller choose $\epsilon$ using PRI.
+Many approaches have been proposed to help controllers choose $\epsilon$ (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Ligett et al., 2017), however, none concentrate on communicating the effects of $\epsilon$ on specific individuals represented in the dataset, as we do.
+In addition, approaches for communicating the privacy guarantees offered by DP are mainly targeted toward end-users who share their data (Nanayakkara et al., 2023; Franzen et al., 2022), whereas our approach focuses on helping controllers, who can only make an informed choice of $\epsilon$ if they understand its privacy implications on those end-users.
+Helping analyst via utility signaling protocol.
+Adding DP-style noise to the output affects the quality and usability of the data for the analyst’s downstream tasks (Hotz et al., 2022).
+For instance, the adoption of DP by the US Census has raised major concerns regarding the reliability of the census statistics after noise infusion and its effect on research and policy-making (Ruggles et al., 2019; Santos-Lozada et al., 2020; Kenny et al., 2021; ABOWD, [n.d.]).
+Our utility signaling protocol takes a step forward in helping analysts interpret the effect of DP on the released output.
+Escrow-based DP platform.
+Various DP platforms have been proposed (Dwork et al., 2006a; Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018; Nguyen et al., 2020; Bittau et al., 2017), but none meets the requirements for deploying our solutions to help controllers and analysts.
+Our escrow-based platform meets those requirements, enabling practical deployment of DP.
+Outline.
 The rest of the paper is organized as follows.
 We present the new privacy risk indicator in Section 2 and the algorithm we designed to automatically find $\epsilon$ and release DP output in Section 3.
-Section 4 presents the utility signaling protocol and the escrow-based platform.
+Section 4 presents the utility signaling protocol and Section 5 presents the escrow-based platform.
 We finish with evaluation (Section 6), related work (Section 7) and conclusions in Section 8.
-2 Privacy Risk Indicator (PRI)
-In this section, we present the new privacy risk indicator. We first provide background on differential privacy.
-Differential privacy (DP) [19][21] is a mathematical definition of privacy that bounds the effect a single individual can have on the output of a query.
+2. Privacy Risk Indicator (PRI)
+Background.
+Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006b) is a mathematical definition of privacy that bounds the effect a single individual has on the output of a query over a dataset of records, where each record represents an individual.
 Formally, a randomized mechanism $\mathcal{M}$ satisfies $\epsilon$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
 $\Pr(\mathcal{M}(x)\in O)\leq e^{\epsilon}\Pr(\mathcal{M}(x^{\prime})\in O).$
-The parameter $\epsilon$ is a tunable parameter called privacy loss budget.
-A typical way of achieving differential privacy is by adding noise to the output.
-The amount of noise is controlled by $\epsilon$; smaller $\epsilon$ leads to larger noise and less accurate output.
-Therefore, $\epsilon$ is the key parameter for navigating the privacy-accuracy trade-off in differential privacy.
-We present two important properties of differential privacy:
-Post-processing.
-[22] DP is immune to post-processing, which means any computation on the output of a DP mechanism will not undermine the DP guarantees.
-Formally, let $\mathcal{M}(x)$ be an $\epsilon$-DP mechanism.
-For any arbitrary function $f$, $f(\mathcal{M}(x))$ is $\epsilon$-DP.
-Sequential composition.
-[22] Sequentially combining multiple DP mechanisms on the same input dataset results in quantifiable degradation of privacy.
-Formally, given $k$ independent mechanisms $\mathcal{M}_{1},\ldots,\mathcal{M}_{k}$ that satisfy $\epsilon_{1},\ldots,\epsilon_{k}$-DP, then any arbitrary function $g(\mathcal{M}_{1},\ldots,\mathcal{M}_{k})$ is $\sum_{i=1}^{i=k}\epsilon_{i}$-DP.
-2.1 Overview
-A good $\epsilon$ provides adequate privacy protection to all individuals while making the output still useful to analysts.
-However, it is inherently difficult to reason about the privacy implications of choosing $\epsilon$ for an specific individual in the dataset.
-Because $\epsilon$ bounds the privacy loss for any individual in expectation, but its concrete effect on a specific individual’s privacy is unknown.
-This is the source of difficulty for controllers whose ultimate goal is to protect individuals in the dataset.
-We observe that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice.
-To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to others.
-For instance, a controller can set an $\epsilon$, compute the corresponding query output, and observe the resulting PRI for each individual.
-If the PRIs are equal, then all individuals will be equally protected.
-If they look different, that is an indication that certain individuals are at a higher risk of being identified and a suggestion for the controller to try a different $\epsilon$.
-The direct connection between the new PRI and differential privacy that we explain next is key to both providing a principled understanding of the new indicator and designing strategies to find $\epsilon$, including an algorithm that finds $\epsilon$ automatically, which we will present in Section 3.
-2.2 Privacy Risk Indicator: Formal Definition
+One way of achieving differential privacy is by constructing a mechanism that adds noise to the query output.
+The amount of noise is controlled by a tunable privacy loss budget parameter, $\epsilon$; smaller $\epsilon$ leads to larger noise and less accurate output: the choice of $\epsilon$ determines the privacy-accuracy tradeoff (Near et al., 2021).
+2.1. Why Choosing $\epsilon$ is Hard and PRI Overview
+Controllers decide what query output to release through $\epsilon$, which impacts individuals’ privacy.
+The parameter $\epsilon$ bounds the maximum privacy loss for any individual, but it is challenging for controllers to reason about the actual privacy implications of releasing an output for specific individuals in the dataset after choosing $\epsilon$ (Dwork et al., 2019).
+We hypothesize that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice.
+To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to other individuals in the dataset, thus providing a direct indication of the privacy impact of releasing a query output.
+We construct PRI so that it is directly connected with the concept of privacy loss, which we explain next.
+This connection is key to both providing a principled understanding of the new indicator and designing strategies to find a suitable $\epsilon$.
+One strategy includes an algorithm to find $\epsilon$ based on the controller’s privacy preferences expressed in terms of PRIs, which we will explain in Section 3.
+2.2. Privacy Risk Indicator: Formal Definition
 We represent a dataset as a set of $n$ records $x=(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, drawn from a data universe $\mathcal{X}$.
@@
 We denote an $\epsilon$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$.
-We consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers.
+Specifically, we consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers.
 Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and the privacy loss budget $\epsilon$ as input, we define the privacy risk indicator of individual $i\in[n]$ as follows:
 Definition 0 (Privacy Risk Indicator).
+The privacy risk indicator of an individual $i$ is defined as:
 $PRI_{i}=\lVert r-q(x_{-i})\rVert_{1},$
-where $x_{-i}$ is the neighboring dataset of $x$ with individual $i$’s record $x_{i}$ removed. The difference is in the L1 norm.
-Computing PRI is $\epsilon$-DP.
-Since $r$ is the output of an $\epsilon$-differentially private mechanism, computing PRI also satisfies $\epsilon$-differential privacy by DP’s post-processing property.
+where $x_{-i}$ is the neighboring dataset of $x$ without $i$’s record $x_{i}$.
 Connection between PRI and privacy loss.
-The PRI has a direct connection to the concept of privacy loss as defined in differential privacy.
-We formally show that when the Laplace Mechanism is used to compute the query output $r$, the connection is simple: larger PRI corresponds to larger privacy loss and vice versa.
-First, we introduce the definition of privacy loss in DP[22].
+Next, we show that when the Laplace Mechanism (Dwork et al., 2014) is used to compute the query output $r$, PRI becomes a relative indicator of individual privacy loss.
+First, we show the definition of privacy loss in DP (Dwork et al., 2014).
 Definition 0 (Privacy Loss).
-Given two neighboring dataset $x,x^{\prime}\in\mathcal{X}^{n}$ such that $\lVert x-x^{\prime}\rVert_{1}\leq 1$, for an output $r\sim\mathcal{Y}$, the privacy loss incurred upon observing $r$ is defined as:
+Given two neighboring datasets $x,x^{\prime}\in\mathcal{X}^{n}$ such that $\lVert x-x^{\prime}\rVert_{1}\leq 1$, for an output $r\sim\mathcal{Y}$, the privacy loss incurred upon observing $r$ is defined as:
 $\mathcal{L}_{\mathcal{M}(x,q,\epsilon)\|\mathcal{M}(x^{\prime},q,\epsilon)}^{(r)}=\ln\left(\frac{\Pr\left[\mathcal{M}(x,q,\epsilon)=r\right]}{\Pr\left[\mathcal{M}(x^{\prime},q,\epsilon)=r)\right]}\right).$
-Next, we introduce the definition of Laplace Mechanism[22], a classic $\epsilon$-differentially private mechanism that we use to compute the query output $r$.
+Next, we introduce the definition of Laplace Mechanism (Dwork et al., 2014), an $\epsilon$-differentially private mechanism that we use to compute the query output $r$.
 Definition 0 (Laplace Mechanism).
@@
 $\mathcal{M}_{L}(x,q,\epsilon)=q(x)+(Z_{1},\ldots,Z_{k}),$
-where $Z_{i}\sim Lap(\frac{GS_{q}}{\epsilon})$ are i.i.d random variables drawn from the Laplace distribution  [^1] The global sensitivity of q in the $\ell_{1}$ norm is defined as $GS_{q}=\max|q(x)-q(x^{\prime})|$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
-Using the Laplace Mechanism, the relative privacy loss of a specific individual $i$ can be expressed as:
+where $Z_{i}$ are i.i.d random variables drawn from the Laplace distribution with probability density function $p_{b}(z)=\frac{1}{2b}\exp(-|z|/b)$ and the scale $b=\frac{GS_{q}}{\epsilon}$  [^1] The global sensitivity of q in the $\ell_{1}$ norm is defined as $GS_{q}=\max|q(x)-q(x^{\prime})|$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
+Using the Laplace Mechanism, we express the relative privacy loss of individual $i$ as:
 $\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}$ $=\ln\left(\prod_{j=1}^{k}\frac{\exp(-\epsilon|r_{j}-q(x)_{j}|/GS_{q})}{\exp(-\epsilon|r_{j}-q(x_{-i})_{j}|/GS_{q})}\right)$
 $=\frac{\epsilon(\lVert r-q(x_{-i})\rVert_{1}-\lVert r-q(x)\rVert_{1})}{GS_{q}}.$
-Since $\lVert r-q(x_{-i})\rVert_{1}$ is the definition of $PRI_{i}$, there exists a clear connection between $PRI_{i}$ and individual $i$’s privacy loss: larger PRI corresponds to larger privacy loss.
-Similarly to PRI, the privacy loss for a specific individual is data and query-dependent; essentially, it is the privacy loss individual $i$ will suffer if they participate in the analysis and the corresponding query output $r$ is released, compared to if they do not participate.
-Note that since we are using the Laplace Mechanism, which is $\epsilon$-DP, the privacy loss will always be bounded by $\epsilon$.
-We have shown that the PRI indicates the relative privacy loss of each individual if a given output is released using DP.
-A data controller can leverage the PRI for many purposes; one natural goal is to exploit the relationship between PRI and DP to choose an $\epsilon$ and the corresponding output to release.
-In the next section, we explain different approaches the controller can use to achieve this goal and propose an algorithm.
-3 Deriving Epsilon based on PRI
-In this section, we describe how to leverage the PRI we introduced to address Challenge 1.
-We explain the two general approaches of using the PRI to find $\epsilon$ in Section 3.1.
-We describe the key observations we made to determine whether the PRIs offer adequate protection and the detailed algorithm for choosing $\epsilon$ and releasing query output in Section 3.2.
-In Section 3.3, we consider multiple queries and discuss how to bound the total privacy loss.
-Finally, we discuss the algorithm complexities and extensions in Section 3.4.
-3.1 Equal PRI means Equal Protection
-The goal of DP is to protect each individual equally in expectation, such that no individual suffers additional privacy loss than others by participating in the analysis; the privacy loss is bounded by $\epsilon$.
-The PRI indicates the effect of $\epsilon$ on each individual, thus becoming an actionable indicator that aids in selecting a suitable value.
-We describe two approaches to help controllers choose $\epsilon$ that leads to exactly equal or statistically similar PRIs.
-We show that both are impractical, which motivates our new algorithm (Section 3.2).
-Figure 1: Example of using DP to equalize the PRIs. The goal is to find the largest $\epsilon$ that can make the PRIs equal.
-Can we make the PRIs equal?
-Figure 1 illustrates this using an example dataset with four records.
-If the output was released directly without adding noise (thus no DP), some individuals would be more at risk than others: those with higher PRI, as shown in the leftmost table.
-By decreasing $\epsilon$, more noise is added to the output, so the PRIs become more similar.
-Then, to protect all individuals equally, the controller should choose the $\epsilon$ that makes the PRIs equal.
-Continuing with the example of Figure 1, setting $\epsilon$ to $0.5$ leads to equal PRIs among the four individuals while $\epsilon=1$ does not.
-When $\epsilon=0.2$, since it produces larger noise, the PRIs will remain equal, but it also leads to less accurate output.
-Therefore, $\epsilon=0.5$ is the best choice among the three values as it achieves the best privacy-accuracy trade-off.
-Although it helps as an illustration, making PRIs exactly equal across individuals is generally infeasible unless an excessive amount of noise is added to the query output, which is undesirable because this means the output will not be useful for the analyst.
-It is also possible that a non-zero $\epsilon$ does not exist for which all PRIs are equal, and the only solution is to add an infinite amount of noise and make the PRIs all infinity.
-Making PRIs statistically similar?
-One strategy is to consider the PRIs similar enough if they are statistically similar.
-Concretely, we can take two biased samples of PRIs: the set of the highest PRIs and the lowest PRIs.
-Then, we test the null hypothesis $H0$: the two samples come from the same distribution, using the statistical hypothesis framework.
-We consider the PRIs to be similar enough if we fail to reject $H0$.
-Since we choose to compare two biased samples, any hypothesis tests that are sensitive to differences in locations (e.g., Mann-Whitney[37], Kolmogorov-Smirnov[29] or Mood’s Median Test) do not work well in practice since the ranks or the medians of the two samples will likely be very different, thus the tests will always tend to reject the null hypothesis.
-On the other hand, if we change the strategy to choosing random samples of PRIs and compare them, it becomes too easy to conclude the samples are similar.
-Therefore, an alternative strategy is needed to determine whether the PRIs provide adequate protection.
-3.2 The Find-$\epsilon$-from-PRI Algorithm
-Two key observations we make lead to our strategy for determining whether the PRIs provide adequate protection.
-O1: Compute ratio between two PRIs.
-The first observation is we should compute the ratio between the PRIs of two individuals to determine whether they receive similar protection, since the comparison depends on the magnitude of the PRI.
-Consider two cases: two individuals with PRI of 2 and 1, and another two individuals with PRI of 10 and 9.
-In both cases, the difference between the two PRIs is 1, but the latter is preferable for equal protection since the ratio is close to 1.
-Figure 2: Left tail with similar larger PRIs (green line) is better than right tail with similar smaller PRIs (red line)
-O2: Left tail is better than right tail.
-The second observation is that if we cannot make the PRIs similar to each other as adding too much noise will make the output useless, it is better only to have left tail while making the larger PRIs similar than have right tail; because right tail means some individuals will have much higher risk than the rest, as shown in Figure 2.
-Even though the green line shows more individuals have relatively larger PRIs, their PRIs are similar enough so that they still receive similar protection; that is, they become indistinguishable with similar privacy loss.
-These two key observations lead to a simple strategy for determining whether the PRIs provide adequate privacy protection.
-Based on O1, we can compute the ratio of the maximum PRI over the minimum PRI.
-Intuitively, as the ratio approaches 1, the PRIs become more similar and closer to the blue line shown in Figure 2.
-However, in many cases, this requires adding too much noise.
-Therefore, O2 allows the controller to trade privacy for more accuracy: instead of choosing the minimum PRI, the controller can choose the PRI at a specific percentile $p$.
-We fix the maximum PRI as the numerator to ensure the riskiest individuals will always be protected.
-The smaller $p$ is, the larger the ratio; therefore, more noise must be added (smaller $\epsilon$) to bring the ratio closer to 1.
-Although $p$ is an additional parameter the controller needs to set, it is more interpretable than $\epsilon$, since the controller only needs to decide at which point all the larger PRIs should be similar.
-For example, if the controller chooses the median (the 50th percentile), a ratio close to 1 indicates no individual is at considerably higher risk than the median individual.
-Algorithm.
-Since PRI and $\epsilon$ are connected in a principled way, we can find a suitable $\epsilon$ that is consistent with the controller’s privacy requirement by computing the PRIs for different $\epsilon$, and choose the $\epsilon$ that is associated with PRIs such that the ratio between the maximum PRI and the PRI at the $p$th percentile is close to 1.
-The main intuition is to test different values of $\epsilon$ in descending order, and find the largest suitable $\epsilon$ that satisfies the controller’s privacy requirement while preserving as much accuracy as possible (note Figure 1).
-Input : Dataset $x=(x_{1},\cdots,x_{n})$; query $q$; DP mechanism $\mathcal{M}$; consumed privacy budget $\epsilon_{c}$; percentile $p$; threshold $\tau$
+Claim: For all individuals $i,j$ where $i\neq j$, if $PRI_{i}>PRI_{j}$, then
+$\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}>\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-j},q,\epsilon)}^{(r)}.$
+The proof is straightforward: if $PRI_{i}>PRI_{j}$, then $\lVert r-q(x_{-i})\rVert_{1}$ $\lVert r-q(x_{-j})\rVert_{1}$, thus the inequality holds.
+Note that since we are using the Laplace Mechanism, which is $\epsilon$-DP, the privacy loss of any individual will always be bounded by $\epsilon$.
+We now make several remarks:
+PRI gives controllers additional information to decide what to release.
+The PRI becomes an additional indicator associated with each individual in the dataset that permits controllers to directly observe the effect of releasing a query output.
+In this paper, we show how PRI helps controllers decide which output to release.
+PRI is data-dependent and thus leaks information about individuals.
+Controllers already have access to individuals’ raw data, so when they gain access to the PRI, they do not learn anything extra about the individuals.
+But this means the PRIs must remain private to the controllers and never be shared with anyone else.
+Controllers may give PRI different uses.
+Controllers may express their privacy preferences for individuals in the dataset in terms of their associated PRIs, and derive an $\epsilon$ that produces that set of PRIs.
+This means that eliciting controllers’ privacy preferences based on PRIs paves the way for choosing $\epsilon$.
+We exploit this insight in the next section to design an algorithm to find $\epsilon$.
+We also discuss other ways in which PRIs may be used in Section 3.5.
+3. Deriving Epsilon based on the PRI
+In this section, we describe an algorithm to choose $\epsilon$ given controllers’ privacy preferences in terms of PRIs, thus addressing Challenge 1 (Sections 3.1, 3.2).
+In Section 3.3, we discuss how to bound the total privacy loss in the presence of multiple queries.
+In Section 3.4, we provide the end-to-end DP guarantee for the algorithm.
+We discuss the algorithm’s complexity and extensions in Section 3.5.
+3.1. Using PRI to Express Privacy Preferences
+Traditionally, controllers need to choose $\epsilon$ to convey their privacy preferences.
+However, choosing $\epsilon$ is challenging because it is hard to understand the specific effect of $\epsilon$ on the privacy of any one individual in the dataset.
+In contrast, the per-individual PRI indicates the relative privacy loss of an individual compared to other individuals in the dataset.
+The consequence is the PRI becomes an actionable indicator that helps controllers decide whether a concrete query output matches their privacy preferences, in which case they may agree to release the output.
+Next, we convey the intuition on how the PRI is useful to controllers through a vignette.
+Figure 1. Example of using PRI to find $\epsilon$. For illustrative purposes, we show the PRI of each student (with the corresponding query output) and leave out students’ actual GPA.
+Vignette illustrating one use of the PRI.
+Figure 1 shows the resulting PRIs of four students after answering a query about their average GPA.
+If the query output were released without using DP (a), the analyst observing the output would learn a lot about those students’ grades, as indicated by the PRI (remember that the PRI is a proxy to privacy loss).
+If the controller used DP, they would have to choose $\epsilon$ and obtain the query output by applying a DP mechanism.
+This would lead to a noisy output and a PRI associated with each student (note that running the mechanism two times with the same $\epsilon$ does not guarantee the same PRIs are obtained).
+We observe that decreasing $\epsilon$ (b, c, d) results in noisier output, which causes the PRIs to become more similar across students.
+In the extreme, when the PRIs are equal (c), no student in the dataset is at a higher risk than others.
+Thus, between (b) and (c), the controller may be more comfortable releasing the specific output (average GPA of 3.2) produced by choosing $\epsilon=1$ because it protects each student equally, which may be a particular privacy preference the controller wants to achieve.
+Lastly, the PRIs in both (c) and (d) indicate no student will be at a higher risk than others, but (d) produces a query output with worse accuracy.
+The controller may choose the larger $\epsilon$ (c) as it yields higher utility for the analyst receiving the output.
+The above is an illustrative example: finding an $\epsilon$ that makes the PRI exactly equal across individuals is generally infeasible: either such an $\epsilon$ does not exist, or it does, but it requires adding too much noise that the output offers little utility to the analyst.
+3.2. The Find-$\epsilon$-from-PRI Algorithm
+When a controller runs a query on a dataset with a DP mechanism, observing the PRIs associated with individuals in the dataset communicates to the controller the relative privacy loss of each individual compared to others.
+Controllers could then change $\epsilon$, rerun the DP mechanism, and observe the resulting PRIs: they could repeat this process until they find a set of PRIs that represents their preferences.
+We introduce an algorithm that implements this intuition, named Find-$\epsilon$-from-PRI.
+In Section 3.1, we have shown one criterion for choosing $\epsilon$ that leads to exactly equal PRIs, which is typically infeasible.
+Instead, we propose an alternative criterion to let controllers decide whether a query output respects their privacy preferences.
+We explain the key idea and then the algorithm.
+Key Idea.
+We let controllers express their privacy preferences by indicating a cutoff for the highest privacy loss of any individual in the dataset.
+This is done by requesting exactly one parameter from the controllers, a percentage p ($0\leq p\leq 100$).
+We compute the minimum and maximum PRI, $PRI_{min}$ and $PRI_{max}$.
+Then we find an $\epsilon$ that produces a list of PRIs (and associated DP output) such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$.
+In other words, we let controllers specify that the maximum PRI should deviate no more than $p$ percent from the minimum PRI.
+For instance, setting $p=0$ means the PRIs should be exactly equal.
+Unlike $\epsilon$, the parameter $p$ directly conveys the level of risk conferred to individuals after releasing a specific query output.
+Smaller $p$ means more privacy at the cost of worse accuracy, hence controllers can trade privacy for accuracy by setting a larger $p$.
+The algorithm aims to find the largest $\epsilon$ that produces a list of PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$.
+Input : Dataset $x=(x_{1},\cdots,x_{n})$; query $q$; percentage $p$
 Output : Chosen privacy parameter $\epsilon$, query output $r$
-1 /* Initialize set of $\epsilon$ values to test; minimum value is $\epsilon_{c}$/
-2 $\mathcal{E}\leftarrow$ initialize-eps-to-test($\epsilon_{c}$)
-3 $x^{p}\leftarrow\textsc{project-query-attrs}(x,q)$
-4 /* Compute $q(x_{-i})$ for all $i$ using memoization */
-5 $C\leftarrow map()$, $L\leftarrow list()$
-6 for $x^{p}_{i}\in x^{p}$ do
-7       if $x^{p}_{i}\notin C$ then
-8             $C[x^{p}_{i}]\leftarrow q(x^{p}_{-i})$
-9      $L.add(C[x^{p}_{i}])$
-10for $\epsilon\in\mathcal{E}$ do
+1 /* Initialize set of $\epsilon$ values to test */
+2 $\mathcal{E}\leftarrow$ initialize-eps-to-test()
+3 /* Compute $q(x_{-i})$ for all individual $i$ using memoization */
+4 $C\leftarrow map()$
+5 for $x_{i}\in x$ do
+6       if $x_{i}\notin C$ then
+7             $C[x_{i}]\leftarrow q(x_{-i})$
+8
+9 for $\epsilon\in\mathcal{E}$ do
 10       /* Compute query output using DP with current $\epsilon$/
-12       $r\leftarrow\mathcal{M}(x,q,\epsilon)$
-13       /* Compute PRI for each individual */
-14       $PRI\leftarrow list()$
-15       for $l\in L$ do
-16             $PRI_{i}\leftarrow\lVert r-l)\rVert_{1}$
-17             $PRI.add(PRI_{i})$
-18      $PRI_{max},PRI_{p}\leftarrow\max(PRI)$, get-percentile$(PRI,p)$
-19       /* A suitable $\epsilon$ is found when the ratio is close to 1 */
-20       if $PRI_{max}/PRI_{p}<1+\tau$ then
-21             $\epsilon_{c}\leftarrow\epsilon_{c}+\epsilon$
+11       $r\leftarrow\mathcal{M}_{L}(x,q,\epsilon)$
+12       /* Compute corresponding PRI of each individual */
+13       $PRIs\leftarrow list()$
+14       for $x_{i}\in C$ do
+15             $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{1}$
+16             $PRIs.add(PRI_{i})$
+17      $PRI_{min},PRI_{max}\leftarrow\min(PRIs),\max(PRIs)$
+18       /* Check whether $\epsilon$ satisfies controller’s privacy preference */
+19       if $PRI_{min}/PRI_{max}\geq 1-(p/100)$ then
 20             return $r$
-23
+21
 Algorithm 1 Find-$\epsilon$-from-PRI
+Detailed Algorithm.
 Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output.
 First, the algorithm initializes the set of $\epsilon$ to choose from (line 1).
-Although the number of possible $\epsilon$ can be infinite, to ensure the algorithm runs in bounded time, we set a range (e.g., from $\epsilon_{c}$ to 10) and include representative $\epsilon$ value in fixed intervals (e.g., 0.1).
-The lower bound $\epsilon_{c}$ starts at 0 and updates for each answered query (we explain how to handle multiple queries in Section 3.3).
-The algorithm then projects relevant query attributes from the dataset (line 1) and computes the query output for each neighboring dataset (line 1 - 1).
-Next, the algorithm iterates over the list of $\epsilon$ from largest to smallest (line 1) and stops when it finds the first suitable value.
-For each $\epsilon$ it tests, the algorithm computes the current query output $r$ by applying the given DP mechanism $\mathcal{M}$ with parameter $\epsilon$ (line 1), and use it to compute the corresponding PRI for each individual (line 1 - 1).
-Finally, the algorithm computes the ratio of the maximum PRI over the PRI at the given $p$th percentile (line 1).
-If the ratio is close to 1 within the threshold $\tau$ (we use 0.01 in our implementation), then a suitable $\epsilon$ has been found.
-The algorithm updates the lower bound $\epsilon_{c}$ and returns the DP output $r$ (line 1 - 1).
-Algorithm data flow.
-The algorithm is used by controllers to find $\epsilon$ from the PRIs and release DP output.
-The controller runs the algorithm on their input dataset and query.
-The algorithm first computes the PRIs, finds a suitable $\epsilon$ based on the controller’s privacy requirement, and returns the DP output computed using this $\epsilon$.
-Nothing is released by default to analysts.
-The PRIs are a by-product of which the sole purpose is to help the controller find $\epsilon$, and they are never released to the data analysts (we will discuss how such promise is enforced in Section 4.3.
-The $\epsilon$ value may need to be released on certain occasions; since finding $\epsilon$ is a post-processing step to computing PRI, which is DP, its release is also DP.
-DP guarantee.
-The algorithm satisfies $\epsilon$-differential privacy.
-The algorithm only returns the query output $r$ when a suitable $\epsilon$ can be found.
-The output $r$ is computed using a standard $\epsilon$-differentially private mechanism (i.e., Laplace Mechanism); thus, the algorithm shares the same DP guarantee.
-This is no different than a standard DP deployment, only that $\epsilon$ is chosen automatically instead of manually by the controller.
-3.3 Bounding Privacy Loss Across Queries
-The challenge of running multiple queries.
+Although the number of possible $\epsilon$ values is infinite, we set an upper bound and a range, e.g., 10 and 0.1, to ensure a finite set of options, e.g., 100 in this example.
+Then, for each $x$’s neighboring dataset $x_{-i}$, the algorithm computes $q(x_{-i})$ (line 1 - 1), which will be used to compute the PRI for each individual $i$.
+Next, the algorithm iterates over the set of $\epsilon$ in descending order (line 1).
+For each $\epsilon$, the algorithm computes the query output $r$ by applying the Laplace Mechanism $\mathcal{M}_{L}$ with parameter $\epsilon$ (line 1) and computes the corresponding PRI of each individual (line 1 - 1).
+Then, given the list of PRIs, the algorithm computes $PRI_{min}$ and $PRI_{max}$ (line 1).
+If $PRI_{min}/PRI_{max}\geq 1-(p/100)$, then the algorithm has found a suitable $\epsilon$ and a query output, so it stops and returns the output.
+Figure 2. Example of $\epsilon$ chosen by the Find-$\epsilon$-from-PRI algorithm with different percentage $p$ over 50 runs
+Running the algorithm with the same dataset $x$, query $q$ and parameter $p$ may identify different $\epsilon$ because the same $\epsilon$ may produce different PRI profiles across executions of the mechanism; simultaneously, similar PRI profiles may be produced by different $\epsilon$ values.
+What is important is that the final query output the analyst releases has been subject to DP and the privacy loss of the individuals is consistent with the controller’s privacy preferences as expressed by $p$.
+In addition, we show empirically in Section 6.3.2 that stricter privacy preferences (i.e., smaller $p$) generally lead to smaller $\epsilon$.
+Figure 2 shows a boxplot representing the distributions of $\epsilon$ chosen by the algorithm with different $p$ (5, 25, 50, 75, 95) over 50 runs; for demonstration, we choose one of the five queries (Q5) we have run.
+As shown in the figure, as $p$ increases, the algorithm tends to choose larger $\epsilon$.
+Setting $p=5$ means the minimum and maximum PRI needs to be very close (within 5%), thus the algorithm mainly finds $\epsilon$ that is less than 1; in contrast, setting $p=95$ means the maximum PRI can be as much as 95% larger than the minimum PRI, thus the $\epsilon$ found can be very large.
+Figure 3. Dataflow of Find-$\epsilon$-from-PRI Algorithm
+Dataflows and leakage.
+It is critically important to understand what information is released to whom.
+Figure 3 shows the involved dataflows in a scenario where a controller uses the Find-$\epsilon$-from-PRI algorithm.
+Individuals send their data to the controller; this dataflow remains unchanged from traditional central DP models (Near et al., 2021).
+The analyst submits a query to the controller.
+The controller expresses their privacy preferences in terms of $p$ (and not in terms of $\epsilon$) and runs the algorithm to find a suitable query output, which is an output produced by a standard DP mechanism (Laplace Mechanism in our work) with a value for $\epsilon$ selected by the algorithm.
+At this point, no information has been returned to the analyst yet.
+The PRI associated with each individual is never released to the analyst.
+The query output that satisfies the controller’s privacy preferences is only released with the controller’s approval.
+The value for $\epsilon$ that led to the query output is also not released by default.
+Releasing $\epsilon$.
+Under certain scenarios (Dwork et al., 2019), $\epsilon$ may also need to be released.
+Since $\epsilon$ is chosen based on PRIs, which are data-dependent, the controller should release $\epsilon$ differential privately by incorporating techniques such as Exponential Mechanism (e.g., Report Noisy Max) (McSherry and Talwar, 2007) or Sparse Vector Technique (Dwork et al., 2009; Lyu et al., 2016) in the algorithm.
+Instead of choosing the largest $\epsilon$ that passes the PRI criterion, the algorithm will choose and return the first $\epsilon$ (which may or may not be the largest) that passes the PRI criterion while preserving DP.
+3.3. Bounding Privacy Loss Across Queries
 In practice, analysts often want to run multiple queries.
-Combining these queries’ output will reveal new information, eventually breaking privacy as per the fundamental Law of Information Recovery [22].
-In today’s deployment of DP, the data controller needs to set a total privacy loss budget a priori on the dataset and stop accepting queries when the cumulative $\epsilon$ exceeds the budget.
-While the total budget provides a rigorous bound of privacy loss in the context of DP, it abstracts away from the actual privacy outcomes, and setting it without context is extremely challenging [23].
-To further exacerbate the problem, the data analyst needs to decide the queries they want to run a priori since they need to decide how to utilize the total budget for their queries.
-This poses a major obstacle for analysts who do not have a pre-determined set of queries.
+Combining these queries’ output will reveal new information, eventually breaking privacy as per the fundamental Law of Information Recovery (Dwork et al., 2014).
+In today’s deployment of DP, controllers need to set a total privacy budget on the dataset a priori and stop accepting queries when the cumulative $\epsilon$ exceeds the budget (Dwork et al., 2019).
+The need for setting a total budget poses a critical challenge for controllers:
+Controllers do not know the privacy implications of total privacy budget.
+Similar to choosing $\epsilon$ for a single query, when deciding which total budget to set, controllers face the same challenge of not knowing the actual privacy implications on individuals after releasing multiple query outputs.
+Oftentimes, they resort to selecting the total budget arbitrarily or through trial-and-error (Dwork et al., 2019).
 The solution.
-Instead of quantifying the maximum acceptable privacy loss, which is hard, we keep track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ of all previously answered queries; we call it $\epsilon_{c}$.
-For an incoming query, when using Algorithm 1 to find $\epsilon$, we truncate the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1).
-We reject answering this query if we fail to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$, meaning the PRIs cannot offer adequate privacy protection for any possible $\epsilon$.
+Instead of quantifying the maximum acceptable privacy loss, which is hard, the algorithm keeps track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ of all previously answered queries; we call it $\epsilon_{c}$.
+For an incoming query, the algorithm truncates the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1).
+The algorithm rejects the query if it fails to find a suitable $\epsilon$ larger than $\epsilon_{c}$, meaning the controller’s privacy preferences cannot be met for any possible $\epsilon$.
 Our solution creates a dynamic bound of total privacy loss.
-As the cumulative privacy loss, $\epsilon_{c}$, keeps growing for each answered query, it becomes more difficult to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$ since less noise can be added to make the PRIs meet the controller’s requirement.
-At each point that we fail to find a suitable $\epsilon$ for a query, the total privacy loss is bounded.
-Using DP’s sequential composition property, releasing the output of those answered queries satisfies $\epsilon_{c}$-differential privacy.
+As $\epsilon_{c}$ keeps growing for each answered query, it becomes more difficult to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$ since larger $\epsilon$ may not produce PRIs that satisfy controllers’ privacy preferences.
+At each point the algorithm fails to find a suitable $\epsilon$ for a query, the total privacy loss is bounded.
 The bound is refreshed if a suitable $\epsilon$ is found for a different query.
-This resulting dynamic bound helps controllers bound the total privacy loss without the need to set the total budget a priori.
-Since the algorithm automatically finds $\epsilon$ (or rejects queries) based on this dynamic bound, it also removes the need for the analyst to decide how to distribute their privacy budget across different queries a priori.
-3.4 Discussion and Extensions
+A potential concern is whether there is additional privacy leakage using this dynamic bound.
+In Section 3.4, we show the bound can be formalized as a privacy odometer in fully-adaptive composition (Rogers et al., 2016) and there is no leakage.
+Considering analysts’ preferences over queries.
+A natural tension exists between analysts and controllers: analysts want more accurate output, whereas controllers want to protect individuals’ privacy by releasing noisy output with degrading accuracy.
+Because it gets increasingly harder to find a suitable $\epsilon$ after answering multiple queries as $\epsilon_{c}$ keeps increasing, analysts may need to prioritize queries of which they need more accurate output.
+One approach to incorporate analysts’ preferences when finding $\epsilon$ is to allow them to assign weights (i.e., priority scores) to their queries, then controllers can adjust the parameter $p$ in the algorithm based on analysts’ preferences, e.g., they can set a higher $p$ for high-priority queries so the algorithm can find larger $\epsilon$ with more accurate output.
+Nevertheless, solving related problems, such as how to communicate analysts’ preferences to controllers, and how to help controllers decide an appropriate $p$ that takes both analysts’ preferences and their own privacy preferences into account, are outside the scope of the paper but remain interesting directions to explore.
+3.4. End-to-End DP Guarantee
+Implications of choosing $\epsilon$ adaptively.
+First of all, if the algorithm has only answered one query, releasing the query output to the analyst satisfies DP and the total privacy loss is bounded by the $\epsilon$ that is chosen by the algorithm.
+However, if the algorithm answered multiple queries (for the same dataset), for each query $\epsilon$ is chosen adaptively because it is chosen based on previously answered queries — a suitable $\epsilon$ needs to be larger than $\epsilon_{c}$.
+Then the total privacy loss is not bounded by any fixed constant since the algorithm does not require setting a total privacy budget.
+This brings the question of whether releasing multiple query outputs computed using adaptively chosen $\epsilon$ causes additional privacy leakage.
+Privacy odometer in fully-adaptive composition.
+To answer this question, we construct a privacy odometer (Rogers et al., 2016) for the algorithm, which provides a probabilistic running upper bound on the realized privacy loss after answering each query.
+The privacy odometer is proposed to provide privacy guarantees in fully-adaptive composition (Rogers et al., 2016), in which the privacy parameters are chosen adaptively as a function of previous interactions.
+Theorem 3.4 in  (Rogers et al., 2016) has proved that there is no additional leakage to choose privacy parameters adaptively if the privacy bound is computed using sequential composition.
+Therefore, the privacy odometer for the Find-$\epsilon$-from-PRI algorithm under sequential composition (Dwork et al., 2014) can simply be defined as:
+Definition 0 (Privacy Odometer).
+[…]
+The privacy odometer gives the same privacy bound as sequential composition; since the algorithm uses $\epsilon$-DP mechanism, $\delta_{g}$ (the global failure parameter in $(\epsilon,\delta)$-differential privacy) can be set to 0.
+The privacy odometer updates its running upper bound on privacy loss to $\epsilon_{c}$ after the algorithm answers each query, and releasing the composed output of all previously answered queries (on the same dataset) satisfies $\epsilon_{c}$-differential privacy.
+It is worth noting that although  (Rogers et al., 2016) showed the privacy odometer gives a worse bound on privacy loss than the standard advanced composition theorem (Dwork et al., 2010), more recently (Whitehouse et al., 2023) proved that in the fully-adaptive setting the privacy odometer can achieve the same asymptotic bound as advanced composition under certain conditions.
+For simplicity, we define the privacy odometer of our algorithm for sequential composition instead of advanced composition.
+3.5. Discussion and Extensions
 Algorithm complexity and optimizations.
@@
 When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$.
-In addition, the query computation can be parallelized since each computation is independent of the other.
-We implement multi-processing and compute the query output for a chunk of records in each process, and we show in Section 5.3.1 that the algorithm is both efficient and scalable.
+In addition, the query computation can be parallelized since each computation is independent of the other, i.e., the problem is embarrassingly parallel.
+Thus, we implement multi-processing and compute the query output for a chunk of records in each process, and we show in Section 6.3.1 that the algorithm is both efficient and scalable.
 Expressing privacy preferences.
-A modification of our algorithm permits the controller to specify protection preferences over certain groups of individuals whose privacy should be protected similarly.
-For example, they can specify individuals whose ages are above 60 and those younger need to receive the same privacy protection.
-This feature can easily be incorporated into the algorithm by letting the controller change the ratio to compute.
-In the example, the controller can choose to compare the median PRI of the two age groups and choose the $\epsilon$ that makes the ratio of the two medians close to 1.
+By default, the algorithm considers the list of PRIs for all individuals in the dataset when finding $\epsilon$.
+A modification of our algorithm permits the controller to specify their privacy preferences in different granularity by selecting subsets of records.
+For example, a controller can divide the dataset $x$ consisting of individuals’ ages to the subset of individuals over 60 ($x^{\prime}$) and those younger ($x^{\prime\prime}$), and specify that the two groups need similar privacy protection.
+This preference can easily be incorporated into the algorithm by letting the controller change the two PRI values to compare instead of the global $PRI_{min}$ and $PRI_{min}$ for all individuals; e.g., they can compute $PRI_{med}^{\prime}$ and $PRI_{med}^{\prime\prime}$, the median PRI for individuals in $x^{\prime}$ and $x^{\prime}$ respectively, and choose an $\epsilon$ that makes $PRI_{med}^{\prime}/PRI_{med}^{\prime\prime}\geq 1-(p/100)$.
 Communicating Privacy-Accuracy Trade-off.
-In some circumstances, the data controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual.
-Prior work such as DPComp[28] and ViP[40] have explored different ways to visualize privacy-accuracy trade-off in differential privacy.
-The PRIs can complement the existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
-4 Helping Analysts Interpret Utility
-In this section, we first propose a simple utility signaling protocol to help analysts understand the impact of DP on their downstream applications to address Challenge 2 (Section 4.1).
-Then, we assemble a set of requirements for a DP platform, which we derive from the solutions proposed (Section 4.2).
-In Section 4.3, we introduce a new implementation of our new algorithm and protocol on top of a data escrow that meets those requirements.
-4.1 The utility signaling protocol
-Today’s DP deployments.
-In the central DP model, we differentiate among individuals who own the original data, controller that manages individuals’ data on their behalf, platform the infrastructure where computation takes place, and analyst the actor demanding computation over individuals’ data.
-The analyst submit queries to a platform operated by a controller.
-The platform computes query output using DP and releases them to the analyst with the controller’s approval.
-We consider the case where the analyst has a downstream task with an associated evaluation function, g, that returns the utility of the downstream task.
-The function takes the set of query output, S, that the analyst received as input.
-Following the notation used in Section 2.2), $S=\{\mathcal{M}(x,q,\epsilon)\mid x\in\mathcal{X}^{n}\}$, $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$.
-For example, an analyst queries census data at the county level (i.e., the US Census Bureau is the controller and citizens are the individuals).
-The set of output $S$ are aggregate statistics on population count, education level, housing prices, etc.
-The analyst’s task is to build a machine learning model to predict the median income of each county based on those statistics, and the evaluation function returns the prediction accuracy, which is g(S).
-Since the queries are computed using DP, releasing $g(S)$ is also DP by post-processing.
-The problem.
-The controller uses DP to release noisy output to protect the privacy of individuals, but the analyst does not know how much such noise affects their downstream task and whether the output is useful.
-In the example, an analyst who obtains a prediction accuracy of $88\%$ does not know how much the accuracy would have been without noise; thus, they cannot decide whether it is worth investing more effort in improving the accuracy of a model.
-Clearly, the platform cannot tell the analyst how much noise was added because that would jeopardize the privacy guarantee.
-A proposed solution: a utility signaling protocol.
-We propose a protocol to help analysts interpret the implications of the added noise on the quality of their downstream task by letting them compare to the quality they would have obtained over the ground truth output.
-Let $S^{\prime}=\{q(x)\mid x\in\mathcal{X}^{n}\}$ be the set of ground truth output (i.e., without applying DP).
-The protocol allows the analyst to obtain both g(S) and g(S’), thus signaling the utility loss due to DP.
-Sending the ground truth output to the analyst so they run their downstream task is not feasible as it breaks the premise of protecting the data in the first place.
-Instead, the analyst sends their downstream task to the platform.
-The platform invokes the evaluation function $g$ over both the set of noisy output $S$ and the set of ground truth output $S^{\prime}$.
-Then, if the controller approves, both $g(S)$ and $g(S^{\prime})$ are released to the analyst to signal the loss in utility due to DP.
-In the example, the platform trains a model over both the DP census statistics and the ground truth statistics, then applies the evaluation function the analyst shared, obtaining $88\%$ and $90\%$ accuracy, respectively.
-Upon receipt of both numbers, the analyst understands the utility effect of DP and can decide whether it is worth investing more effort in improving the model.
-Assessing Privacy Implications of Releasing $g(S^{\prime})$: Any information release may leak information about the individuals’ data, e.g., even the accuracy of a model (consider the trivial case where $S^{\prime}$ contains only one value, the analyst can simply use the evaluation function to test different values until they obtain an accuracy of 100%).
-Since DP is not being applied when releasing $g(S^{\prime})$, it is ultimately at the discretion of the controller to decide whether or not to release $g(S^{\prime})$.
-In the example, the controller may be comfortable releasing the accuracy of the machine learning model trained on the ground truth census statistics (in addition to the model accuracy on the DP statistics), but not releasing the model itself.
-Our utility signaling protocol does not attempt to address the possibility of various attacks the analyst may perform to infer individual data from $g(S^{\prime})$.
-Instead, the escrow-based platform we describe next allows the controller to examine the analyst’s downstream task and associated evaluation function and ensures $g(S^{\prime})$ will not be released without the controller’s permission.
-For many applications, the benefit of allowing the analyst to observe $g(S^{\prime})$ can be tremendous.
-4.2 Platform Requirements
-We propose a new platform to support the Find-$\epsilon$-from-PRI algorithm that helps controllers (addressing Challenge 1) and the utility signaling protocol that helps analysts (addressing Challenge 2).
-Such a platform must meet the following requirements:
+In some circumstances, the controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual.
+Prior work such as DPComp (Hay et al., 2016) and ViP (Nanayakkara et al., 2022) have explored different ways to visualize the privacy-accuracy trade-off in DP.
+The PRI complements the existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
+4. Helping Analysts Interpret Utility
+In this section, we concentrate on solutions to help analysts interpret the utility of DP output.
+We revisit Challenge 2 (Section 4.1) and present a solution we call utility signaling protocol (Section 4.2).
+4.1. Challenge Overview and Key Insight
+Controllers work on a platform that has access to the individuals’ data and that is responsible for answering analysts’ queries using DP.
+An analyst submits a set of queries $Q=\{q\mid q\in\mathcal{Q}\}$ ($\mathcal{Q}$ is the query space) to the platform, and (if a suitable $\epsilon$ is found for each query) receives a set of noisy query output $S=\{\mathcal{M}(x,q,\epsilon)\mid x\in\mathcal{X}^{n},q\in Q\}$ ($\mathcal{M}$ is a DP mechanism and $\mathcal{X}$ is the data universe).
+A well-known challenge of using DP is that analysts do not know the utility of $S$ (Hotz et al., 2022).
+Concretely, analysts do not know the effect of using $S$ on their downstream task, which is the purpose for analysts to query the data in the first place.
+Example 1/3.
+An analyst queries census data at the county level.
+The set of query output $S$ is aggregate data on population count, education level, and housing prices that are released using DP.
+The analyst uses $S$ to train a binary classifier to predict the income level of each county and obtains an accuracy of 68%.
+The analyst does not know if the (low) accuracy is due to the effect of DP or inappropriate training data.
+The difficulty in interpreting the utility of $S$ makes it challenging for analysts to understand how the DP output will impact their downstream tasks; this is one of the core issues surrounding the discussion of the use of DP for the US Census (Ruggles et al., 2019; Santos-Lozada et al., 2020; Kenny et al., 2021; ABOWD, [n.d.]).
+Observation.
+Analysts use the DP output $S$ for some downstream task, such as a machine learning or inference task.
+Downstream tasks often have an associated evaluation function, $g(\cdot)\rightarrow\mathbb{R}$, which returns a number indicating the quality of the downstream task when evaluated on a certain input.
+In the example above, the downstream task is to train a binary classifier, the evaluation function measures the quality of the classifier trained on DP output and returns 68% as the quality indication (accuracy in this case).
+Insight: signaling DP effect.
+Analysts would learn the impact of DP if they had access to both the result of the evaluation function on the DP output, $g(S)$ [^2] We slightly abuse notation for simplification.
+and the raw output, $g(S^{\prime})$, where $S^{\prime}=\{q(x)\mid x\in\mathcal{X}^{n},q\in Q\}$ (the set of query output without using DP).
+Example 2/3.
+Besides having access to the quality of the classifier on the DP census aggregates (68%), the analyst also gains access to the quality of the classifier if it had been trained on the raw census aggregates $S^{\prime}$, say 80%.
+By comparing 68% ($g(S)$) and 80% ($g(S^{\prime})$), the analyst learns the effect of added DP noise on their downstream task.
+Gaining access to $g(S)$ and $g(S^{\prime})$ helps analysts interpret the impact of DP on their downstream tasks.
+The remaining challenge is how to provide $g(S^{\prime})$ to the analysts without giving them access to $S^{\prime}$, which is the raw query output and cannot be released to the analyst without violating the privacy of individuals in the dataset.
+4.2. The utility signaling protocol
+Rather than the platform directly sending $S^{\prime}$ to the analyst, which would break individuals’ privacy, the analyst submits their downstream task and associated evaluation function $g(\cdot)$ to the platform.
+Upon successful inspection by the controller, the platform executes the downstream task and evaluation function on behalf of the analyst, producing $g(S)$ and $g(S^{\prime})$.
+Last, if the controller approves, the platform releases both $g(S)$ and $g(S^{\prime})$ to the analyst, thus signaling the impact of DP on the analysts’ task without revealing $S^{\prime}$.
+Example 3/3.
+Instead of only submitting the queries to the platform, with the utility signaling protocol, the analyst submits their queries, the downstream task, and the evaluation function.
+The controller inspects each and allows the platform to signal the impact of using DP to the analyst by releasing both utility results 68% and 80%.
+4.2.1. Protocol Step by Step
+Figure 4. End-to-end example of utility signaling protocol. We illustrate the protocol in the central DP deployment scenario where the controller has access to the raw data.
+Figure 4 illustrates the utility signal protocol, which we describe next:
+0.
+Platform.
+The platform has access to the dataset $x$ of individuals.
+It also has the capacity to run queries using DP.
+It may optionally implement the Find-$\epsilon$-from-PRI algorithm presented in Section 3.2333While we envision the algorithm being deployed to help controllers, both contributions are orthogonal.
+1. Analyst submits queries to the platform. The platform notifies the controller.
+2.
+Controller inspects queries.
+The controller inspects the analyst’s queries and allows the platform to execute the queries as long as a suitable $\epsilon$ is selected.
+The platform withholds both the set of DP output $S$ and raw output $S^{\prime}$ without releasing anything to the analyst by default.
+3.
+Platform releases $S$ upon controller’s approval.
+The platform notifies the controller after computing $S$ and $S^{\prime}$.
+The controller gives permission for the platform to release $S$ to the analyst.
+Up to this point, the protocol provides the analyst with the same information they can access in a typical central DP deployment.
+Now, we extend the protocol to implement the utility signaling component to help the analyst understand the impact of DP on their downstream task.
+4. Analyst submits downstream task and evaluation function $g(\cdot)$ to the platform. The platform notifies the controller.
+5.
+Controller inspects downstream task and evaluation function.
+The controller inspects the implementation of the analyst’s downstream task and evaluation function $g(\cdot)$, and decides whether it is appropriate to release $g(S)$ and $g(S^{\prime})$.
+With the controller’s approval, the platform executes the downstream task and evaluation function $g(\cdot)$ on both $S$ and $S^{\prime}$, withholding the results by default.
+6.
+Platform releases $g(S)$ and $g(S^{\prime})$ upon controller’s approval.
+The controller can inspect $g(S^{\prime})$ one last time to ensure it is appropriate to release; $g(S)$ preserves DP since it is a post-processing (Dwork et al., 2014) step of DP output $S$.
+If approved, the platform releases $g(S)$ and $g(S^{\prime})$ to the analyst, who learns the impact of DP on their downstream task without ever gaining access to $S^{\prime}$ directly.
+4.2.2. Privacy implications of releasing $g(S^{\prime})$
+Any data release leaks information.
+When releasing an output that is subject to DP, such as $g(S)$, controllers understand the consequences on the individual’s privacy.
+Conversely, when releasing output not subject to DP, such as $g(S^{\prime})$, controllers must assess whether doing so maintains individuals’ privacy.
+That ultimately depends on the dataset $x$, the output $S^{\prime}$, the implementation of the downstream task and evaluation function, $g(\cdot)$, and the adversary model.
+The scope of the utility signaling protocol is to enable analysts to understand the impact of using DP, but not on understanding, controlling, and determining when it is safe to release $g(S^{\prime})$, which is ultimately at the discretion of the controllers and an interesting avenue of future work.
+5. Escrow-Based DP Platform
+We curate a set of requirements for a DP platform that implements the Find-$\epsilon$-from-PRI algorithm and the utility signaling protocol in Section 5.1.
+In Section 5.2, we describe the implementation of a platform built on top of a data escrow that meets these requirements.
+5.1. Platform Requirements
+The solutions to Challenge 1 (the Find-$\epsilon$-from-PRI algorithm), to Challenge 2 (the utility signaling protocol), and the nature of processing over private data call for a platform that meets the following requirements:
   • R1.
-Control Data Flow: Analysts should never see the raw data or PRIs.
-And they should not submit queries and downstream tasks over the raw data directly without the controller’s approval (e.g., the controller may approve releasing the model accuracy to signal utility loss).
+Control dataflows: Given the three relevant entities (analyst, platform, and controller), a dataflow indicates data movement between any two of those entities.
+A requirement for the platform is to ensure that only permitted dataflows take place.
+Concretely, controllers indicate what dataflows take place, and the platform enforces the controller’s decision.
+In the context of our work, analysts can access for example the DP output $S$ after the controller approves, and they cannot access PRIs or any other intermediates not explicitly approved by the controller.
   • R2.
-Trust: The new protocol requires more trust on the platform.
-Trust is a broad concept, and its meaning varies depending on context[7].
-In the context of DP, trust means the platform adheres to prespecified behavior and the DP mechanism upholds its privacy guarantee[7][43].
-Individuals must trust the controller will protect their data in a responsible manner (i.e., using DP).
-Controllers must trust the platform keeps the data secure and runs computations on the data correctly.
-With the new utility signaling protocol, analysts must also trust the platform runs their downstream task correctly.
-We want to reduce the trust those entities have on the platform.
+Secure platform: Adversarial agents may break the requirement of controlling dataflows by breaking into the platform and forcing unintended dataflows.
+This is an assumed risk in the central DP model, which can be reduced by using a local DP deployment (Near et al., 2021).
+Unfortunately, local DP reduces the accuracy of the output compared to central DP.
+A requirement for the platform is to ameliorate the risk of unintended dataflows.
   • R3.
-High Performance: The platform must limit the overhead, which is at odds with the pressures of the new Find-$\epsilon$-from-PRI algorithm that helps controllers at the cost of higher computational overhead.
-We highlight a tension between R2 and R3.
-Improving trust often requires employing techniques (e.g., cryptographic primitives) that increase the computational overhead, potentially leading to impractical solutions.
-We now explain the design of a platform that addresses these three requirements.
-4.3 Escrow-Based DP Platform
-We propose a new platform, based on an escrow architecture similar to the proposed in  [50], that lets controllers maintain fine-grained control of data flow (meeting R1), reduces the trust on the controller and platform by both individuals and analysts (meeting R2) all while maintaining high performance to meet R3.
-4.3.1 Controlling Data Flow to Implement the utility signaling protocol
-A data escrow supports delegated computation, which means the escrow runs computation on behalf of the computation owner.
-Implementing a platform with an escrow lets analysts delegate their downstream tasks to the platform, thus obtaining the two utilities $g(S)$ and $g(S^{\prime})$.
-We present the end-to-end protocol in Figure 3, which consists of seven steps.
-Figure 3: End-to-end utility signaling protocol. We illustrate the protocol in the central DP deployment scenario where the controller has access to the raw data.
-1. Registering data. The controller registers the dataset containing individual records with the escrow.
-2.
-Registering application.
-The controller registers an application that uses DP to run queries on the registered dataset, which includes the Find-$\epsilon$-from-PRI algorithm.
-Any derived data product, such as the PRI, will not be released without the controller’s permission.
-The PRI can also be computed separately as an independent function inside the application if the controller wants to use it for other purposes.
-3.
-Creating policy.
-The data controller creates a policy to let the analyst run queries on the registered dataset using the registered application; a policy in the escrow indicates who can run what computation on what data.
-4.
-Issuing Queries.
-The analyst issues queries to the application.
-The application answers each query and stops when a suitable $\epsilon$ can no longer be found.
-After checking that the query computation only ran on data accessible to the analyst, the escrow releases the set of DP output $S$ to the analyst because the controller wrote the corresponding policy.
-The set of ground truth output $S^{\prime}$ is registered inside the escrow since it is a derived data product of the computation.
-Up to this point, the protocol provides analysts with the same information they can access in a typical central DP deployment.
-Now, we extend the protocol to implement the utility signaling component to help analysts understand the impact of DP on their downstream tasks.
-5.
-Registering downstream task.
-The analyst registers their downstream task and an associated evaluation function $g$ as an application in the escrow.
-The evaluation function is the only function within the application that is available for the escrow to invoke and anyone to create policy on; if exists, the rest of the functions should be invoked from the evaluation function to perform other necessary computations.
-The escrow ensures that only the evaluation function’s output (and nothing else) will be released to the analyst.
-6.
-Creating policy for downstream task.
-After inspecting the application submitted by the analyst and determining it is acceptable to release $g(S^{\prime})$, the controller creates a policy to run the evaluation function $g$ on the ground truth output $S^{\prime}$.
-7.
-Evaluating downstream task.
-With the controller’s permission, the analyst obtains both $g(S)$ and $g(S^{\prime})$.
-By inspecting the difference between the two utilities, the analyst learns the impact of DP on their task, thus helping them decide whether the DP output is sufficiently good for their task.
-By exploiting the delegated computation features of the data escrow, the platform controls data flow, which is necessary to implement the utility signaling protocol.
-4.3.2 From Trusted to Trustworthy, while Fast
-A trusted platform adheres to prespecified behaviors[7], in the context of DP this means an adversary only has access to the output of a DP mechanism, but not any of its internal state and any raw or derived data[22].
-This is infeasible to achieve in practice.
-A central platform, as in the central DP model, can claim to be trusted, but such claim does not preclude it from being compromised and intentionally leaking data in the clear, thus breaking the DP guarantee[43].
-A trustworthy platform, on the other hand, implements different techniques to enforce the promise that not only will it never behave unexpectedly, it will not pose harm (to the users) in any way[27].
-An escrow-based platform reduces the impractical assumptions of a trusted platform and supports trustworthy computation by ensuring:
-  • No data or derived data product (e.g., query output) ever leaves the platform without explicit permission from the controller, and only explicitly allowed computation runs on existing data; this is the basis of the control of data flow.
-  • Data is encrypted end-to-end, even during computation.
-End-to-end encryption can be achieved in many ways; for instance, DP platforms leveraging secure multiparty computation[42][9][10][49][45], homomorphic encryption[4][46], Oblivious RAM[12] and trusted execution environment[44].
-The prototype we use for our evaluation uses AMD Secure Encrypted Virtualization(SEV)[33].
-  • Tasks that are sent to the escrow are the tasks that run.
-The escrow is registered with an escrow application; the platform ensures that the correct version of the application is running using techniques such as remote attestation[15] or formal verification[38][35].
-Our prototype uses hardware-rooted remote attestation enabled by AMD SEV.
-Figure 4: Left: central platform. Right: escrow-based platform
-By computationally enforcing the promises above, the escrow reduces the trust that both individuals and analysts must place in the platform while maintaining the same accuracy guarantee as the central DP model.
-Furthermore, the escrow application can be written so that controllers do not gain access to the raw data but only the PRIs (i.e., by writing a policy that only reveals the PRIs to the controller), thus illustrating multiple escrow-based deployment scenarios.
-Consider Figure 4, which shows how individuals, controller, and analyst interact with the central platform in CDP and the escrow-based platform.
-In CDP, the controller manages individuals’ data using the central platform and sets $\epsilon$ manually; the individuals need to trust both the controller to protect their privacy (i.e., by choosing a good $\epsilon$) and the platform to operate correctly without leaking their data.
-Whereas using the escrow-based platform, individuals send data directly to the platform, and the controller only has access to PRIs computed by Algorithm 1.
-The escrow invokes Algorithm 1 to find $\epsilon$ on the controller’s behalf and releases the output to the analyst, who delegates their downstream task to the escrow to obtain utility.
-5 Evaluation
-Our evaluation shows our solutions address both Challenge 1 to help controllers, and Challenge 2 to help analysts. We answer the following research questions for our evaluation:
-  • RQ1: Does presenting the privacy risk indicator help data controllers choose $\epsilon$?
+Protect against adversarial downstream and evaluation functions: Adversarial analysts could submit downstream tasks and evaluation functions crafted to release more information than needed.
+While controllers must inspect these functions before agreeing to release the output and they are the ultimate arbiter, a requirement of the platform is to facilitate the controllers’ job.
+  • R4.
+High Performance: The Find-$\epsilon$-from-PRI algorithm increases processing overhead.
+Simultaneously, security safeguards to address R2 are often at odds with high-performance computation.
+Together, these two facts risk making the platform impractical.
+A requirement is to implement the solutions to Challenge 1, Challenge 2, and meet the requirements R1, R2, and R3 while maintaining low overhead.
+We now explain the implementation of a platform that addresses these three requirements.
+5.2. Platform Design and Implementation
+To demonstrate the feasibility of the solutions to Challenge 1 and Challenge 2, we describe how to implement the solutions above on a platform that leverages a data escrow (Xia et al., 2022) architecture.
+Data escrow overview.
+A data escrow is a system that acts as an intermediary for storing, processing, and releasing data.
+Different entities interact within the escrow.
+Controllers register data and control the dataflows.
+The data is never stored or computed in the clear using end-to-end encryption.
+Analysts run computations on registered data in the escrow.
+For any data or its derived data product to leave the escrow, controllers have to create a policy on the escrow.
+The policies are tamper-proof since the escrow validates the identity of the controller who creates the policy and the policy remains encrypted, thus ensuring the controller fully controls the dataflow.
+Meeting R1.
+No raw data (e.g., $x$) or derived data product (e.g., PRIs, $g(S)$, $g(S^{\prime})$) leave the escrow without an explicit policy that permits that.
+Policies then determine what dataflows take place.
+Controllers are in charge of writing policies.
+Meeting R2. Data is encrypted end-to-end, even during computation. Adversaries that break into the platform cannot read plain-text data. In other words, no unintended dataflows take place.
+Meeting R3.
+The platform can let analysts submit their downstream task and associated evaluation function, and put the burden on the controller to verify the implementation of those functions.
+Or it can offer standard implementation of downstream tasks and evaluation functions and let analysts call these, thus requiring scrutinizing the functions once and using them many times.
+The data escrow permits the deployment of programs expressing a desired functionality.
+In specific DP scenarios, e.g., the US Census, it is conceivable that one could provide programs implementing (at least) the more standard functionalities and query interfaces, and rely on controllers to scrutinize only one-off tasks.
+Meeting R4.
+End-to-end data encryption (necessary to meet R2) can be achieved in many ways; for instance, DP platforms leveraging Secure Multiparty Computation (Narayan and Haeberlen, 2012; Bater et al., 2016, 2018; Wang et al., 2022; Roth et al., 2019), Homomorphic Encryption (Agarwal et al., 2018; Roy Chowdhury et al., 2020), Oblivious RAM (Bogatov et al., 2021) and Trusted Execution Environment (Nguyen et al., 2020).
+Our evaluation prototype uses AMD Secure Encrypted Virtualization (Kaplan et al., 2016) because it permits running unmodified software while maintaining the integrity and confidentiality of memory contents.
+The escrow design protects access to external resources (e.g., network and disk), ensuring data is encrypted transparently to any application.
+As we show in the evaluation (Section 6.2), this solution virtually removes the processing overhead.
+In addition, we implement a series of optimizations to the Find-$\epsilon$-from-PRI algorithm to demonstrate its efficiency and scalability when run in the escrow (Section 6.3.1).
+6. Evaluation
+Our evaluation shows our solutions address Challenge 1 to help controllers, and Challenge 2 and 3 using an escrow-based platform. We answer the following research questions:
+  • RQ1: Does presenting the PRI help data controllers choose $\epsilon$?
 […]
-  • Microbenchmarks: We conduct experiments to show the
-    Find-$\epsilon$-from-PRI algorithm is efficient and scalable. In addition, we study how sensitive is our algorithm to different values of parameter $p$.
-5.1 RQ1: Using PRI To Choose Epsilon
-To determine whether the PRI help the data controllers choose $\epsilon$, we conducted an IRB-approved user study to analyze how PRI affects the controller’s decision-making regarding choosing $\epsilon$.
-Our goal is to show that by presenting the PRI to the participants, they will be more consistent on the $\epsilon$ they choose compared to if they don’t have the PRI.
-To this end, we conducted a between-subject user study where the participants are divided into control and treatment groups.
+  • Microbenchmarks: We conduct experiments to show that
+    Find-$\epsilon$-from-PRI algorithm is efficient and scalable. In addition, we study the effect of varying parameter $p$ in the algorithm.
+6.1. RQ1: Using PRI To Choose Epsilon
+To determine whether PRI helps the data controllers choose $\epsilon$, we conducted an IRB-approved user study to analyze how PRI affects the controller’s decision-making regarding choosing $\epsilon$.
+To this end, we conducted a between-subjects user study where the participants are divided into control and treatment groups.
 To prevent learning effects, only the treatment group sees the PRI.
-We chose to conduct the study in the form of an online survey, because our tasks do not require in-person guidance and it is more convenient for participants to complete the study.
+We chose to conduct the study in the form of an online survey because our tasks do not require in-person guidance and it is more convenient for participants to complete the study.
 In addition, it allows us to collect data from a more representative sample.
-5.1.1 Recruitment and data collection
-We recruited 100 participants using Prolific[1].
+6.1.1. Recruitment and data collection
+We recruited 100 participants using Prolific (pro, [n.d.]).
 To complete the study, each participant must sign a consent form, which describes the purpose and content of the study, compensation, and data confidentiality.
@@
 This tutorial explains the intuition of DP and the effect of the parameter $\epsilon$ at a high level.
-Mainly, it explains DP is generally achieved by adding random noise to the query output, and $\epsilon$ is inversely proportional to the amount of noise added.
+Mainly, it explains that DP is generally achieved by adding random noise to the query output, and larger $\epsilon$ means less noise will be added and vice versa.
 After reading the tutorial, the participant needs to complete a screening quiz that tests their knowledge about the tutorial they just read.
 Only participants who pass the quiz proceed to conduct the study.
-Out of the 100 participants we recruited, 56 of them passed the quiz and their responses are used for the study analysis.
-This quiz ensures that participants have understood the principles of DP and thus ensures the validity of the user study.
-5.1.2 Study procedure
-First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group and 27 participants are in the treatment group.
-The survey presents a sample from the UCI Adult dataset[2] and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$.
-Then, it presents two queries in random order and asks the participant to choose an $\epsilon$ value from a slider, with values from 0.5 to 5.0 and each interval being 0.5.
-The participant’s task is to choose an $\epsilon$ value that offers equal privacy protection to all people in the dataset while preserving as much accuracy as possible.
-For the control group, the participant is presented with a table that shows the range of possible output for each $\epsilon$, indicating output accuracy, along with the ground truth (non-DP) output.
+Of the 100 participants we recruited, 56 passed the quiz and their responses were used for the study analysis.
+This quiz ensures that participants have understood the principles of DP, thus strengthening the validity of the user study.
+6.1.2. Study procedure
+First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group, and 27 are in the treatment group.
+The survey presents a sample from the UCI Adult dataset (uci, [n.d.]) and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$.
+Then, it presents two queries in random order:
+  • The number of people with income greater than 50K and are less than 40 years old.
+  • The number of people who have a Master’s degree or above and have income $\leq$ 50k, grouped by each race.
+For each query, the survey asks the participant to choose an $\epsilon$ value from a slider, with values from 0.5 to 5.0 in an interval of 0.5.
+Since we want all participants to have a clear task goal in mind, the survey includes a description of the task: choose $\epsilon$ to protect individuals’ privacy while yielding as accurate answers as possible.
+For the control group, the participant is presented with a table showing the range of possible output for each $\epsilon$ to indicate output accuracy and the raw (non-DP) output.
 The table includes information that any controller would have access to in a deployment today when choosing $\epsilon$.
-For the treatment group, the participant is presented with the same table, but with an additional column that includes a plot showing the range of PRIs for each $\epsilon$ (e.g., a box drawn from the minimum to maximum PRI).
+For the treatment group, the participant is presented with the same table, but with an additional column showing the range of PRIs for each $\epsilon$ (e.g., a box drawn from the minimum to maximum PRI); we also provide a brief explanation of the PRI.
 Note that this is the only change between the control and treatment groups to ensure we isolate the intervention.
-We explain beforehand what the PRI is at a high level, and that a smaller range represents more similar PRIs among people in the dataset.
 A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include these to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
-5.1.3 Results
-Figure 5: $\epsilon$ chosen by each participant
-We compare the distribution of $\epsilon$ chosen by participants in the control and treatment groups.
-Figure 5 shows a swarm plot where each point represents the $\epsilon$ value chosen by each participant in the control and treatment group for each query.
+6.1.3. Results
+Figure 5. $\epsilon$ chosen by each participant
+If the intervention (showing the PRIs) is effective, we expect this would manifest as more consistent answers among participants, since all participants were given the same training and the same task.
+Therefore, we compare the distribution of $\epsilon$ chosen by participants in the control and treatment groups.
+Figure 5 shows a swarm plot where each point represents the $\epsilon$ value chosen by each participant in the control or treatment group.
 For both queries, 19 out of 27 participants in the treatment group chose either 2 or 1.5, showing a more consistent answer than the control group.
 This is because the PRI range when $\epsilon$ is 2 or 1.5 is considerably smaller than that of larger $\epsilon$, and smaller $\epsilon$ has lower accuracy.
-This shows that most participants could connect the PRIs and their task goal, which is to offer equal privacy protection while ensuring high accuracy, and they were able to use the PRIs to choose the appropriate $\epsilon$ that satisfies the goal.
-On the contrary, the $\epsilon$ values chosen by participants in the control group are more spread out and have greater variance, suggesting participants shared no general consensus as to what is a good $\epsilon$ to pick.
-In addition, we ran the Mann–Whitney test [37] on the $\epsilon$ values in the control and treatment group, with the alternative hypothesis being the two samples come from different distributions.
+This shows that most participants could connect the PRIs and their task goal, which is to protect individuals and release useful output, and they were able to use the PRIs to choose the appropriate $\epsilon$ that satisfies the goal.
+On the contrary, the $\epsilon$ values chosen by participants in the control group are more spread out and have greater variance, suggesting participants shared no general consensus as to what is a good $\epsilon$ to choose.
+We ran Mann–Whitney test  (McKnight and Najab, 2010) on the $\epsilon$ values chosen from the control and treatment group, with the alternative hypothesis being the two samples come from different distributions.
 For both queries, we obtained $p<0.005$.
 Thus we reject the null hypothesis and conclude that showing the PRIs helped participants choose $\epsilon$.
-In the follow-up question for the treatment group, all 27 participants answered yes when asked whether they used the PRIs when choosing $\epsilon$, confirming our hypothesis that presenting the PRIs helped participants choose the $\epsilon$ that satisfies the task goal.
-Figure 6: $\epsilon$ chosen by participants who do not know DP (left) and those who know DP (right)
+Last, we asked participants in the treatment group if they used the PRIs to complete the task: all 27 participants responded they did use that information, confirming they were indeed “treated”.
+Figure 6. $\epsilon$ chosen by participants who do not know DP (left) and those who know DP (right)
 Does knowing DP matter?
-Another interesting question is whether the participant’s a priori familiarity with DP affects how they choose $\epsilon$.
+Another interesting question is whether the participant’s prior familiarity with DP affects how they choose $\epsilon$.
 At the start of the survey before showing the tutorial, we asked each participant whether they know about DP.
-27 participants (out of a total of 56) answered they don’t know anything about DP and 29 participants answered they know about DP.
+27 participants (out of a total of 56) answered they do not know anything about DP and 29 participants answered they know about DP.
 Figure 6 shows the $\epsilon$ chosen by participants based on their familiarity.
@@
 How do different ways of presenting PRIs affect participants’ decisions when choosing $\epsilon$?
-Since this user study relies on the participants inspecting and interpreting the PRIs with their own eyes to choose $\epsilon$, instead of using the algorithm to automatically find one, how we present the PRIs to the participants can greatly affect their choices.
-We explored an alternative way of presenting PRIs - by using histograms.
+Since this user study relies on the participants inspecting and interpreting the PRIs with their own eyes to choose $\epsilon$ instead of using the algorithm to automatically find one, how we present the PRIs to the participants can greatly affect their choices.
+We explored an alternative way of presenting PRIs by using histograms.
 To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and asked them to complete the same survey questions in the treatment group, with the only change being presenting histograms of PRIs instead of range plots.
-Figure 7: $\epsilon$ chosen by participants who were presented with histograms or range plots, or in the control group
+Figure 7. $\epsilon$ chosen by participants who were presented with histograms or range plots, or in the control group
 Figure 7 shows $\epsilon$ chosen by participants who were presented with histograms or range plots, and we include the $\epsilon$ chosen by participants in the previous control group as reference.
 Participants who saw histograms did not perform better than participants who saw range plots.
-Especially for Q2, 8 participants chose larger $\epsilon$ (3.5, 4, and 4.5).
-Although we cannot be certain since the responses are anonymous, one possible explanation might be that for the large $\epsilon$, the majority of PRI are already small and similar to each other.
-Only a small portion of PRI have large values.
+Especially for Q2, eight participants chose larger $\epsilon$ (3.5, 4, and 4.5).
+Since the responses are anonymous, we cannot directly ask participants about their decision-making after the survey ends.
+We deduce one possible explanation might be that for the large $\epsilon$, the majority of PRIs are already small and similar, and only a small portion of PRIs have large values.
 The participants might think that as long as the privacy of the majority of people is being protected, the task goal is achieved; hence, they chose larger $\epsilon$, so the results are more accurate.
@@
 Exploring other alternative solutions for presenting and helping users interpret PRI remains an interesting problem.
-5.1.4 Threats to validity
-External validity. We recruited participants from diverse backgrounds across the world via Prolific, with few prescreen criteria such that the participants can complete the study without difficulty.
-Internal validity.
-To ensure all participants have the minimum knowledge of DP needed to complete the study, each participant must read a tutorial and pass a quiz.
-We only allowed the participant to take the quiz once to minimize the chance of participants abusing it if they got multiple chances.
+6.1.4. Threats to validity
+External validity: We recruited participants from diverse backgrounds worldwide via Prolific, with few prescreen criteria so that the participants could complete the study without difficulty.
+Internal validity: To ensure participants have the minimum knowledge of DP needed to complete the study, each participant must read a tutorial and pass a quiz.
+Participants can only take the quiz once to minimize the chance of them abusing it if they have multiple chances.
 To avoid learning effects, we used a between-subject study to isolate the condition we intended to test.
 We also randomized the two questions for choosing $\epsilon$.
-5.2 RQ2: DP On Escrow-based Platform
+6.2. RQ2: DP On Escrow-based Platform
 […]
-5.2.1 Baselines for comparison
+6.2.1. Baselines for comparison
 We evaluate and compare the performance of the four following baselines:
-  • Data Escrow: We implemented a private query escrow application that takes a SQL query as input and applies the Laplace Mechanism to release the query output.
-Optionally, if the parameter $\epsilon$ is not given, it will use Algorithm 1 to find $\epsilon$ automatically.
-We simulated the protocol where the data controller registers their dataset to the escrow, creates a policy that allows the data analyst to issue queries to their data using the private query application, and the analyst issues their query and receives the query output.
+  • Data Escrow: We implemented an escrow application that takes a SQL query as input and applies the Laplace Mechanism to release the query output.
+Optionally, if the parameter $\epsilon$ is not given, it will use the Find-$\epsilon$-from-PRI algorithm to select a suitable $\epsilon$.
+We simulate the two entities (controller and analyst) in the escrow.
+A policy is first created by the controller to allow the query output to be released to the analyst.
 […]
-  • Standalone without DP: This baseline simply runs the given query without DP. The dataset is in plaintext.
-  • Standalone with DP: This baseline represents the central DP model. It runs the given query and applies the Laplace Mechanism to the query output. The dataset is in plaintext.
+  • Standalone with DP: This baseline runs the given query and applies the Laplace Mechanism to the query output in a central platform.
+The dataset is in plaintext.
+This represents the central DP deployment scenario without the overhead of additional security measures.
+  • Standalone without DP: This baseline simply runs the given query without DP in a central platform. The dataset is in plaintext. We include this baseline to show the overhead of using DP.
 […]
-5.2.2 Setup
+6.2.2. Setup
 We ran the experiments on n2d-standard-8 instances with 8 virtual CPUs (AMD Rome) and 64GB memory from Google Compute Engine.
-We used AMD’s Secure Encrypted Virtualization (SEV) when running the data escrow.
+We used AMD Secure Encrypted Virtualization (SEV) when running the data escrow.
 We used Ubuntu 20.04 and Python 3.8 for all baselines.
-5.2.3 Dataset
+6.2.3. Dataset
 […]
-5.2.4 Queries
+6.2.4. Queries
 We run the 3 queries independently for all baselines:
-  • Q1: Number of people between age 20 and 30 (simple predicate with range)
-  • Q2: Number of people by each race (aggregate & group-by)
-  • Q3: Number of people who are female and have income greater than 50K (two simple predicates)
-5.2.5 Results
+  • Q1: Number of people between age 20 and 30
+  • Q2: Number of people by each race
+  • Q3: Number of females who have income greater than 50K
+6.2.5. Results
 […]
-Figure 8: Average query time over 10 runs
+Figure 8. Average query time (in log scale) over 10 runs
 […]
 […]
-5.3 Microbenchmarks
+6.3. Microbenchmarks
 We evaluate the runtime and scalability of the Find-$\epsilon$-from-PRI algorithm in Section 6.3.1 and study how different values of the parameter $p$ affect the $\epsilon$ chosen in Section 6.3.2.
-5.3.1 Is Find-$\epsilon$-from-PRI practical?
+6.3.1. Is Find-$\epsilon$-from-PRI practical?
 Since the algorithm can run multiple times if there are multiple queries , the algorithm needs to be efficient so running it multiple times will still be practical.
-We ran the algorithm as a registered application in the escrow.
+We ran the algorithm in the data escrow to further demonstrate its performance for computation-heavy tasks.
 Datasets.
 We controlled the experiment by generating datasets of size 1000, 10000, 100000, and 1 million based on the Adult dataset.
-We sampled records for size datasets of size 1000 and 10000.
-For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they are not the same copy of each other.
-We included all attributes and all records in the dataset.
+We sampled records for datasets of sizes 1000 and 10000.
+For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they were not the same copy of each other.
+We included all attributes in the dataset.
 Queries.
@@
 The queries are run independently.
-  • Q1: SELECT COUNT(*) FROM adult WHERE income==’>50K’ AND education_num=13 AND age==25
-  • Q2: SELECT marital_status, COUNT(*) FROM adult WHERE race==’Asian-Pac-Islander’ AND age>=30 AND age<=40 GROUP BY marital_status
-  • Q3: SELECT COUNT(*) FROM adult WHERE native_country!=’United-States’ AND sex==’Female’
-  • Q4: SELECT AVG(hours_per_week) FROM adult WHERE workclass==’Federal-gov’ OR workclass==’Local-gov’ or workclass==’State-gov’
+  • Q1: SELECT COUNT(*) FROM adult WHERE income==’>50K’ AND education_num==13 AND age==25
+  • Q2: SELECT marital_status, COUNT(*) FROM adult WHERE race==
+    ’Asian-Pac-Islander’ AND 30<=age<=40 GROUP BY marital_status
+  • Q3: SELECT COUNT(*) FROM adult WHERE native_country!=
+    ’United-States’ AND sex==’Female’
+  • Q4: SELECT AVG(hours_per_week) FROM adult WHERE workclass in (’Federal-gov’, ’Local-gov’, ’State-gov’)
   • Q5: SELECT SUM(fnlwgt) FROM adult WHERE capital_gain>0 AND income==’<=50K’ AND occupation==’Sales’
 Parameters.
-We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,0.099,\cdots,0.001]$ (37 values in total).
-We set $\epsilon_{c}=0$, $p=50$ and $\tau=0.01$.
+We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,\\ 0.099,\cdots,0.001]$ (37 values in total).
+We set $p=50$.
 We used six parallel processes for computing the PRIs.
-Figure 9: Average runtime to find $\epsilon$ over 10 runs
+Figure 9. Average runtime (in log scale) to find $\epsilon$ over 10 runs
 Results.
@@
 For example, for 1 million records, Q5 has approximately 100000 unique records of which the PRIs need to be computed; whereas for other queries, the algorithm only needs to compute PRIs for no more than a few thousand records and the rest can be directly retrieved using memoization.
-5.3.2 Effect of parameter $p$
-To find a suitable $\epsilon$, the algorithm computes the ratio of the maximum PRI over the PRI at the $p$th percentile; if the ratio is close to 1, the PRIs are considered to reflect adequate privacy protection by the controller’s privacy requirement.
-Specifically, we are interested in choosing the 00th (minimum), $25$th, and $50$th (medium) percentile.
-The smaller $p$ is, the stricter the privacy requirement and the harder it is to find $\epsilon$, meaning more noise needs to be added to achieve a ratio closer to 1.
+6.3.2. Effect of parameter $p$
+The Find-$\epsilon$-from-PRI algorithm finds a suitable $\epsilon$ by letting controllers specify their privacy preferences through a parameter $p$, and checks whether $PRI_{min}/PRI_{max}\geq 1-(p/100)$ for each $\epsilon$.
+Smaller $p$ means stricter privacy preferences because the minimum and maximum PRI need to be closer.
+We test the effect of varying $p$ in the algorithm; specifically, when $p=5,25,50,75,95$.
 Setup.
-We used the Adult dataset with 100000 rows that we generated in Section 5.3.1, and we used the same settings (except $p$).
-We ran the five queries in Section 5.3.1 10 times and we report the $\epsilon$ selected by our algorithm in each run.
-Figure 10: $\epsilon$ found when using the PRI at the 0, 25, and 50th percentile to compare against the maximum PRI
+We use the Adult dataset with 100000 rows that we generated in Section 6.3.1.
+We ran each of the five queries from Section 6.3.1 for 50 times, and report the $\epsilon$ chosen by our algorithm in each run.
+Figure 10. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$
 Results.
-Figure 10 shows the $\epsilon$ found when $p=0,25,50$.
-As expected, using higher percentiles tends to find larger $\epsilon$ than lower percentiles, because less noise needs to be added to bring the maximum PRI closer to the PRI at the larger percentile.
-When using the minimum PRI, Q1 – Q4 find very small $\epsilon$ that are between 0 and 1.
-Q1 and Q3, both counting queries, are able to find larger $\epsilon$ since the queries have a global sensitivity of 1, i.e., the maximum difference between PRIs will also be fixed at 1, it becomes easier to make the ratio closer to 1.
-Whereas Q2 and Q4, a group-by query and an average query with higher sensitivities, require more noise to be added to bring the PRIs closer.
-For Q5, using the minimum PRI only finds $\epsilon$ once, since the sensitivity is too large, primarily due to the domain size of fnlwgt), the DP mechanism will not be able to add enough noise of such scale beyond certain $\epsilon$ (when $\epsilon=0.1$, the noise scale will be larger than 10,000,000, which is the threshold we use to determine whether the DP mechanism can be applied).
-However, when the requirement is relaxed, it is able to find larger $\epsilon$ such that the noise scale does not exceed the threshold.
-6 Related Work
+Figure 10 shows boxplots of $\epsilon$ found by the algorithm for different $p$.
+As shown in the figure, setting larger $p$ tends to find larger $\epsilon$.
+Setting $p=5$ results in very small $\epsilon$ chosen for all five queries because the maximum PRI can only be no more than 5% from the minimum PRI; most $\epsilon$ values are less than 0.1 for Q1–4 and less than 1 for Q5.
+Queries with larger sensitivity (Q2, Q4, Q5) can find larger $\epsilon$ when $p>50$, and queries with sensitivity of 1 (Q1, Q3) tend to find smaller $\epsilon$.
+Overall, the results empirically demonstrate that the algorithm is useful for finding a suitable $\epsilon$ (and corresponding query output) that is consistent with the controller’s privacy preferences.
+7. Related Work
+Defining PRI.
+The PRI shares similar intuition as Personalized Differential Privacy (PDP) (Jorgensen et al., 2015) in that different individuals may need different levels of protection since some records are more sensitive than others.
+In contrast to PDP, which requires each individual to specify their own privacy requirement through a privacy budget, the PRI is computed by controllers to indicate the realized privacy loss of each individual if a DP output were to be released.
 Choosing privacy loss budget.
-Different approaches [30] [26] [23] [36] [39] [40] [28] [48] [32] [28] aim to help users choose $\epsilon$ when using DP.
-One line of work [39] [40] [28] [48] [32] implements an interface to visualize the effect of different variables (e.g., utility metrics) when adjusting $\epsilon$.
+Different approaches (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Hay et al., 2016; Ligett et al., 2017) aim to help users choose $\epsilon$ when using DP.
+One line of work (Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021) implements interfaces to visualize the effect of different variables (e.g., utility metrics) when adjusting $\epsilon$.
 Nanayakkara et al.
-visualizes the disclosure risk in relation to $\epsilon$; they use the definition of disclosure risk from Lee and Clifton, which is the probability of disclosing the presence or absence of any individual in the database.
+visualizes the disclosure risk in relation to $\epsilon$ using the definition of disclosure risk from Lee and Clifton, which is the probability of disclosing the presence or absence of any individual in the database.
 John et al.
 defines and visualizes the data sharing risk; Both metrics quantify potential data leakage, but neither quantifies leakage with respect to each individual in the data, unlike our PRI.
+Ligett et al.
+empirically finds the smallest $\epsilon$ consistent with an accuracy requirement in the private empirical risk minimization setting.
 Ge et al.
@@
 Hsu et al.
-proposes an economic model for choosing $\epsilon$ by balancing the interests of two different parties: the data analyst and the prospective participant who will choose to participate if the benefit outweighs the risk.
-However, this model assumes that the participant can quantitatively express the "cost" of whether to participate in the study, which is impractical for non-experts.
-In addition, Cummings et al.
+proposes an economic model for choosing $\epsilon$ by balancing the interests of the data analyst and the prospective participant who will participate if the benefit outweighs the risk.
+However, this model assumes that the participant can quantitatively express the ”cost” of whether to participate in the study, which is impractical for non-experts.
+Cummings et al.
 and Xiong et al.
-investigate user’s understanding of DP guarantees through multiple user studies, which offer insights on facilitating communicating the privacy implications of DP to users and help them choose $\epsilon$.
-Jayaraman and Evans evaluates the impact of $\epsilon$ on ML tasks and specifically their privacy-utility tradeoffs.
+investigate users’ understanding of DP guarantees through multiple user studies, which offer insights on how to communicate the privacy implications of DP to users.
 To the best of our knowledge, no known approach helps users understand the implications of $\epsilon$ on individuals’ privacy, and use that to design an algorithm that finds $\epsilon$ automatically.
@@
 Tao et al.
-proposes DPXPlain to explain aggregate query answers under DP by comparing answers with and without DP, and assessing whether the difference is due to the noise introduced by DP or properties of the data, thus helping analysts interpret whether the answers are useful for their analysis.
+proposed DPXPlain to explain aggregate query answers by comparing answers with and without DP and assess whether the difference is due to the noise introduced by DP or properties of the data, thus helping analysts interpret whether the answers are useful for their analysis.
 This idea, complementary to ours, relies on analysts to come up with comparison questions that require explanation in order to understand any anomalies or trends in the query answers.
-Moreover, feasibility studies [8][14] have been conducted to evaluate real-world usage of DP.
 CDP LDP Shuffler
 Crypto
-Primitive
-Data
 Escrow
-T?
+T
 No Yes Yes Yes
@@
 Trust
-indv’ls
+individuals
 add noise
@@
 Error $O(1)$ $O(\sqrt{n}$) Near $O(1)$ $O(1)$ $O(1)$
-Table 1: Comparisons of different DP deployment scenarios. T stands for trustworthy.
+Table 1. Comparisons of different DP deployment scenarios. T stands for trustworthy.
 Deploying differential privacy.
-Alternatives beyond the central model aim to deploy DP with fewer trust assumptions.
+The threat model of the central DP model assumes adversaries cannot breach the platform and force unintended data flows, which is unrealistic.
+Alternatives beyond the central model aim to deploy DP with different trust assumptions.
 We summarize those different deployment scenarios in Table 1.
-The local model (LDP) has been proposed to eliminate the need for a trusted data curator and it has been adopted in various real-world applications for collecting client data from their software and devices [24] [18] [17].
-In LDP, each individual perturbs their own data using LDP mechanisms before sending it to the data curator; the data curator does not need to be trusted since it only sees the differentially private version of data.
-However, the primary disadvantage of the local DP model is that it introduces a much larger error than the central DP model.
-The shuffle model creates an anonymous communication channel that amplifies the privacy guarantees [34] [6], thus it requires less noise to be added when users perturb their data and offers better accuracy than the local DP model.
+The local model (LDP) has been proposed to eliminate the need for a trusted platform and it has been adopted in various real-world applications for collecting client data from their software and devices  (Erlingsson et al., 2014; Ding et al., 2017; Differential Privacy Team, 2017).
+In LDP, each individual perturbs their own data using LDP mechanisms before sending it to the platform; adversaries can control the platform but they only gain access to DP data.
+However, LDP also introduces a much larger error ($O(\sqrt{n})$, where n is the number of individuals) than the central DP model ($O(1)$).
+The shuffle model creates an anonymous communication channel that amplifies the privacy guarantees  (Kasiviswanathan et al., 2011; Balle et al., 2018), thus it requires adding less noise when users perturb their data and it offers better accuracy than the local DP model.
 However, it requires additional trust assumptions as the shuffler must be trusted.
-Lastly, techniques that leverage cryptographic primitives, specifically secure multiparty computation, homomorphic encryption, oblivious RAM, and trusted hardware enclaves, have been proposed to offer security guarantees when deploying DP without a trusted data curator [20] [42] [9] [4] [45] [46] [49] [10] [44] [11] [12].
+Lastly, techniques that leverage cryptographic primitives, specifically secure multiparty computation, homomorphic encryption, oblivious RAM, and trusted hardware enclaves, have been proposed to offer security guarantees when deploying DP without a trusted data curator  (Dwork et al., 2006a; Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018; Nguyen et al., 2020; Bittau et al., 2017; Bogatov et al., 2021).
 We showed that the escrow-based solution we introduce presents an interesting alternative with a lower computational cost.
-7 Conclusions
-To ease the deployment of DP, we address challenges faced by controllers and analysts by 1) defining the privacy risk indicator (PRI) that indicates the relative privacy loss of each individual for a given $\epsilon$ and output and an algorithm to choose $\epsilon$ using the PRIs.
-2) introducing an utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks.
-We propose a platform implementation on top of a data escrow.
-We demonstrated our contributions through an IRB-approved user study and empirical evaluation with other DP platforms.
+8. Conclusions
+In this paper, we showed that by including individual-level information about privacy loss (the PRI), controllers gain additional understanding of the effects of applying DP to the results of an analyst’s query.
+This translates into less burden choosing $\epsilon$: we implement this idea via an algorithm, Find-$\epsilon$-from-PRI.
+In addition, we show that it is possible to augment analysts’ understanding of the effects of DP on their results by using a utility signaling protocol that reveals the isolated impact of DP on a downstream task.
+We then gather the requirements gleaned from the previous two contributions and propose a platform, based on a data escrow, to deploy both ideas.

### v3

@@
-Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform
+Making Differential Privacy Easier to Use for Data Controllers using a Privacy Risk Indicator
 Abstract.
-Differential privacy (DP) enables private data analysis but is hard to use in practice.
-For data controllers who decide what output to release, choosing the amount of noise to add to the output is a non-trivial task because of the difficulty of interpreting the privacy parameter $\epsilon$.
-For data analysts who submit queries, it is hard to understand the impact of the noise introduced by DP on their tasks.
-To address these two challenges: 1) we define a privacy risk indicator that indicates the impact of choosing $\epsilon$ on individuals’ privacy and use that to design an algorithm to choose $\epsilon$ and release output based on controllers’ privacy preferences; 2) we introduce a utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks.
-We implement the algorithm and the protocol inside a new platform built on top of a data escrow, which allows controllers to control dataflows while maintaining high performance.
-We demonstrate our contributions through an IRB-approved user study, extensive experimental evaluations, and comparison with other DP platforms.
-All in all, our work contributes to making DP easier to use by lowering adoption barriers.
-PVLDB Reference Format:
-PVLDB, 14(1): XXX-XXX, 2020.
-doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
-doi:XX.XX/XXX.XX
+Differential privacy (DP) enables private data analysis but is difficult to use in practice.
+In a typical DP deployment, data controllers manage individuals’ sensitive data and are responsible for answering data analysts’ queries while protecting individuals’ privacy; they do so by choosing $\epsilon$, the privacy loss budget, which controls how much noise to add to the query output.
+However, it is challenging for data controllers to choose $\epsilon$ because of the difficulty of interpreting the privacy implications of such a choice on the individuals they wish to protect.
+To address this challenge, we first derive a privacy risk indicator (PRI) directly from the definition of ex-post per-instance privacy loss in the DP literature.
+The PRI indicates the impact of choosing $\epsilon$ on individuals’ privacy.
+We then leverage the PRI to design an algorithm to choose $\epsilon$ and release query output based on data controllers’ privacy preferences.
+We design a modification of the algorithm that allows releasing both the query output and $\epsilon$ while satisfying differential privacy, and we propose a solution that bounds the total privacy loss when using the algorithm to answer multiple queries without requiring controllers to set the total privacy loss budget.
+We demonstrate our contributions through an IRB-approved user study and experimental evaluations that show the PRI is useful for helping controllers choose $\epsilon$ and our algorithms are efficient.
+Overall, our work contributes to making DP easier to use for controllers by lowering adoption barriers.
+Differential privacy, data controller, privacy loss budget
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data management systemsccs: Security and privacy Privacy protectionsccs: Security and privacy Usability in security and privacy
 1. Introduction
-Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006b) is a promising tool to analyze data about individuals, but it remains challenging to use in practice.
-In a typical DP deployment, a data analyst runs queries on a dataset hosted on a platform that is accessible to a data controller.
-The dataset contains information about individuals, and releasing this information risks revealing the identity of those individuals.
-The data controller wants to protect individuals’ privacy while releasing useful output to the data analyst.
+Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006) is a promising tool to analyze data about individuals, but it remains challenging to use in practice.
+In a typical DP deployment, a data analyst runs queries on a dataset containing sensitive information about individuals that is accessible to a data controller.
+We define a data controller as an agent with the capacity to (1) directly access and run analysts’ queries on individuals’ raw data (2) make decisions on what query output to release to the analysts since the output may reveal the identity of individuals.
+The controller is responsible for protecting individuals’ privacy while releasing useful output to the analysts.
 DP guarantees that the output of the query is likely to be the same if we change or remove any individual’s record in the dataset.
-This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: smaller $\epsilon$ leads to more accurate output but weaker privacy guarantees.
-Despite some successful deployments of DP (Abowd, 2018; Apple, 2017; Erlingsson et al., 2014; Ding et al., 2017), a few roadblocks prevent broader DP deployment.
-In this paper, we consider three such roadblocks:
-Challenge 1: Choosing $\epsilon$.
-Controllers need to decide the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$.
+This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: larger $\epsilon$ leads to more accurate output but weaker privacy guarantees.
+Despite some successful deployments of DP (Abowd, 2018; Apple, 2017; Erlingsson et al., 2014; Ding et al., 2017), a major challenge exists which prevents broader DP deployment.
+Challenge of choosing $\epsilon$.
+Controllers need to choose the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$.
 Dwork et al.
-showed that controllers have no clear consensus on how to choose $\epsilon$ or how to approach this problem in a principled manner.
-Our contribution.
+(Dwork et al., 2019) showed that controllers have no clear consensus in choosing $\epsilon$ or approaching this problem in a principled manner.
+Our Contributions
+Designing PRI and the Find-$\epsilon$-from-PRI algorithm.
 In this paper, we observe the crux of the problem is that $\epsilon$ does not directly translate into the degree of privacy loss incurred to specific individuals in the dataset.
-In response, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual.
-Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a specific query output affects individuals’ privacy.
-And because PRI is connected to $\epsilon$, there is an opportunity to automatically find an $\epsilon$ that produces a profile of privacy risk indicators consistent with the controller’s privacy preferences; we exploit this opportunity by designing an algorithm, called Find-$\epsilon$-from-PRI, to choose $\epsilon$ and release DP output given a simple description of controllers’ privacy preferences.
-Challenge 2: Interpreting utility of DP output.
-Analysts submit queries and receive noisy output that deviates from the raw output, i.e., the ground truth output if DP had not been applied.
-Without access to such ground truth, analysts do not know the impact of DP on the quality of their downstream tasks, which is the reason why they sought the DP output in the first place.
-This hinders the analysts’ ability to use the DP output in practice.
-Our contribution.
-In this paper, we propose a utility signaling protocol to help analysts interpret the impact of DP on their downstream tasks while still allowing controllers to fully control data release.
-In the protocol we propose, the analyst delegates their downstream task with an associated evaluation function to the platform.
-With the controller’s permission, the platform invokes the evaluation function on both the DP output and the raw output and releases both results, thereby signaling the utility loss on the analyst’s downstream task due to noise introduced by DP.
-Challenge 3: Deploying our solutions in a DP platform.
-Deploying our solutions to the previous two challenges brings a new challenge on the platform design.
-Specifically, implementing the new Find-$\epsilon$-from-PRI algorithm (addressing Challenge 1) and utility signaling protocol (addressing Challenge 2) calls for a platform that meets the four requirements: (1) it allows controllers to control dataflows, (2) it implements necessary security measures to protect against unintended dataflows, (3) it facilitates controllers to inspect dataflows, and (4) it adds little overhead.
-Various DP platforms have been proposed, but none meet all four requirements.
-For instance, the central DP model assumes adversaries are incapable of corrupting the platform and forcing unintended dataflows, which is unrealistic in practice (Near et al., 2021); platforms leveraging cryptographic primitives (Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018) make fewer assumptions about the adversaries, but often have significant computation overhead.
-Our contribution.
-We propose a new platform built on top of a data escrow (Xia et al., 2022) which enables the implementation of the Find-$\epsilon$-from-PRI algorithm and utility signaling protocol, meeting the four platform requirements we curated.
-We show that the escrow-based platform hits a new point in the design space, offering a better trade-off between security, usability, and performance.
-[…]
-Summary of Contributions and Related Work
-Helping controller choose $\epsilon$ using PRI.
-Many approaches have been proposed to help controllers choose $\epsilon$ (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Ligett et al., 2017), however, none concentrate on communicating the effects of $\epsilon$ on specific individuals represented in the dataset, as we do.
-In addition, approaches for communicating the privacy guarantees offered by DP are mainly targeted toward end-users who share their data (Nanayakkara et al., 2023; Franzen et al., 2022), whereas our approach focuses on helping controllers, who can only make an informed choice of $\epsilon$ if they understand its privacy implications on those end-users.
-Helping analyst via utility signaling protocol.
-Adding DP-style noise to the output affects the quality and usability of the data for the analyst’s downstream tasks (Hotz et al., 2022).
-For instance, the adoption of DP by the US Census has raised major concerns regarding the reliability of the census statistics after noise infusion and its effect on research and policy-making (Ruggles et al., 2019; Santos-Lozada et al., 2020; Kenny et al., 2021; ABOWD, [n.d.]).
-Our utility signaling protocol takes a step forward in helping analysts interpret the effect of DP on the released output.
-Escrow-based DP platform.
-Various DP platforms have been proposed (Dwork et al., 2006a; Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018; Nguyen et al., 2020; Bittau et al., 2017), but none meets the requirements for deploying our solutions to help controllers and analysts.
-Our escrow-based platform meets those requirements, enabling practical deployment of DP.
+We hypothesize knowing the effect of choosing $\epsilon$ on each individual’s privacy would help controllers make the choice.
+Therefore, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual.
+The PRI is a relative indicator of ex-post per-instance privacy loss (Redberg and Wang, 2021), which is the realized privacy loss incurred to an individual upon releasing a specific output.
+Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a specific query output affects each individual’s privacy.
+And because PRI is connected to $\epsilon$, there is an opportunity to automatically find an $\epsilon$ that produces a profile of PRIs consistent with the controller’s privacy preferences; we exploit this opportunity by designing an algorithm, called Find-$\epsilon$-from-PRI, to choose $\epsilon$ and release query output given a simple description of controllers’ privacy preferences.
+Privately releasing $\epsilon$.
+The Find-$\epsilon$-from-PRI algorithm only releases the DP output and not the chosen $\epsilon$, since $\epsilon$ is chosen based on the PRIs, which are data-dependent, and releasing $\epsilon$ directly may undermine the DP guarantees.
+Thus we introduce a modification of the algorithm, called Find-and-release-$\epsilon$-from-PRI, which permits releasing both the query output and $\epsilon$ while preserving DP.
+Bounding privacy loss across multiple queries.
+In many cases, analysts submit multiple queries to the controller to obtain different information about the dataset.
+Besides the challenge of choosing $\epsilon$ for each query, controllers also face the challenge of setting the global privacy budget which bounds the maximum privacy loss for a dataset.
+We propose a solution that does not require controllers to quantify the maximum acceptable privacy loss; instead, it dynamically bounds the privacy loss after answering (or denying) each query.
+We construct a privacy odometer (Rogers et al., 2016) to provide an end-to-end DP guarantee for our solution.
+Evaluation.
+In our evaluation, we test the hypothesis that the new privacy risk indicator helps controllers choose $\epsilon$ by conducting an IRB-approved user study with 56 participants.
+We show that our algorithms are efficient and scalable, and we conduct ablation studies to analyze the effect of varying controllers’ privacy preferences.
 Outline.
 The rest of the paper is organized as follows.
+We introduce the background of differential privacy and the different agents in standard DP deployment in Section 2.
 We present the new privacy risk indicator in Section 3 and the algorithm we designed to automatically find $\epsilon$ and release DP output in Section 4.
-Section 4 presents the utility signaling protocol and Section 5 presents the escrow-based platform.
+In Section 5, we introduce the modified algorithm to privately release $\epsilon$ and our solution to dynamically bound privacy loss across multiple queries.
 We finish with evaluation (Section 6), related work (Section 7) and conclusions in Section 8.
+2. Background
+In this section, we first introduce the definition of $(\epsilon,\delta)$-differential privacy (Section 2.1), then explain the different agents in a standard DP deployment (Section 2.2).
+2.1. Differential Privacy
+Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006) is a mathematical definition of privacy that bounds the effect a single individual has on the output of a query over a dataset of records, where each record represents an individual.
+Formally, for an $\epsilon,\delta\geq 0$, a randomized mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
+$\Pr(\mathcal{M}(x)\in O)\leq e^{\epsilon}\Pr(\mathcal{M}(x^{\prime})\in O)+\delta.$
+When the failure parameter $\delta$ is 0, $\mathcal{M}$ satisfies $\epsilon$-differential privacy; otherwise $\delta$ is usually set to be a very small probability (e.g., $\frac{1}{n^{2}}$ where $n$ is the number of records).
+One way of achieving DP is by constructing a mechanism that adds random noise to the query output.
+The amount of noise is controlled by a tunable privacy loss budget parameter, $\epsilon$.
+Smaller $\epsilon$ leads to more noise and less accurate output: the choice of $\epsilon$ determines the privacy-accuracy tradeoff in DP (Near et al., 2021).
+There are two important properties of differential privacy, post processing and sequential composition, which are building blocks for designing sophisticated mechanisms that satisfy DP.
+Post processing (Dwork et al., 2014).
+DP is immune to post processing, which means any computation on the output of a DP mechanism will not undermine the DP guarantees.
+Formally, let $\mathcal{M}(x)$ be an $\epsilon$-DP mechanism.
+For any arbitrary function $f$, $f(\mathcal{M}(x))$ is $\epsilon$-DP.
+Sequential composition (Dwork et al., 2014).
+Sequentially combining multiple DP mechanisms on the same input dataset results in quantifiable degradation of privacy.
+Formally, given $k$ independent mechanisms $\mathcal{M}_{1},\ldots,\mathcal{M}_{k}$ that satisfy $\epsilon_{1},\ldots,\epsilon_{k}$-DP, then any arbitrary function $g(\mathcal{M}_{1},\ldots,\mathcal{M}_{k})$ is $\sum_{i=1}^{i=k}\epsilon_{i}$-DP.
+2.2. Agents in DP Deployment
+Figure 1. Agents in standard DP deployment
+Figure 1 shows a diagram of a standard DP deployment, i.e., the central DP model (Near et al., 2021) (Wagh et al., 2021); we define a dataflow as a function that passes data from one entity to another.
+Individuals send their raw data to a data controller.
+The data analyst submits queries to the controller in order to obtain information about the individuals.
+The goal of the controller is to release useful output to the analyst while protecting individuals’ privacy.
+To achieve this, the controller answers the analyst’s queries using DP mechanisms, and they control both how much protection they want to offer to individuals and how accurate the output should be by setting the $\epsilon$ parameter when applying DP to answer each query.
+In addition, the controller needs to set a global privacy budget to control the maximum acceptable privacy loss incurred on the dataset; answering each query consumes some budget, once the budget is reached, the dataset is retired and can no longer be queried.
+One notable example of the central DP model is the application of DP by the US Census in 2020 (Abowd, 2018).
+In this case, the US Census Bureau is the data controller who has access to individuals’ (i.e., the US population) data.
+The data analysts’ queries are pre-determined by the Census Bureau; the census data is the output of various census queries (e.g., population counts in different geographical levels by age, ethnicity, race, and sex).
+The Census Bureau sets the privacy loss budget $\epsilon$ and allocates the budget to each query (Bureau, 2021b)  [^1] Technically, the Census Bureau allocates the budget using the privacy parameter $\rho$ in zero-concentrated differential privacy (Bun and Steinke, 2016), which can be converted to the $\epsilon$ in $(\epsilon,\delta)$-differential privacy.
+We omit the details for simplicity., then uses DP to release the query output.
+The Census Bureau made considerable effort to choose $\epsilon$ and allocate the budget, including testing different $\epsilon$ values on a demonstration census dataset and seeking extensive feedback from experts (Bureau, 2021a).
+However, when deploying DP in practice, choosing $\epsilon$ is an extremely challenging task for controllers without the assistance of DP experts (Dwork et al., 2019).
+While there are other DP deployment models available, i.e., the local DP (Kasiviswanathan et al., 2011) or shuffler model (Erlingsson et al., 2019), where the agents interact differently (e.g., individuals no longer send their raw data to the controller), in this paper we concentrate on addressing the controller’s challenges of choosing $\epsilon$ in the central model.
 3. Privacy Risk Indicator (PRI)
-Background.
-Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006b) is a mathematical definition of privacy that bounds the effect a single individual has on the output of a query over a dataset of records, where each record represents an individual.
-Formally, a randomized mechanism $\mathcal{M}$ satisfies $\epsilon$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
-$\Pr(\mathcal{M}(x)\in O)\leq e^{\epsilon}\Pr(\mathcal{M}(x^{\prime})\in O).$
-One way of achieving differential privacy is by constructing a mechanism that adds noise to the query output.
-The amount of noise is controlled by a tunable privacy loss budget parameter, $\epsilon$; smaller $\epsilon$ leads to larger noise and less accurate output: the choice of $\epsilon$ determines the privacy-accuracy tradeoff (Near et al., 2021).
 3.1. Why Choosing $\epsilon$ is Hard and PRI Overview
 Controllers decide what query output to release through $\epsilon$, which impacts individuals’ privacy.
-The parameter $\epsilon$ bounds the maximum privacy loss for any individual, but it is challenging for controllers to reason about the actual privacy implications of releasing an output for specific individuals in the dataset after choosing $\epsilon$ (Dwork et al., 2019).
+The parameter $\epsilon$ bounds the maximum privacy loss for any individual a priori before running the DP mechanism to produce the query output, but it is challenging for controllers to reason about the concrete privacy implications for individuals represented in the dataset when releasing a specific output ex-post, i.e., after applying the DP mechanism.
 We hypothesize that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice.
 To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to other individuals in the dataset, thus providing a direct indication of the privacy impact of releasing a query output.
-We construct PRI so that it is directly connected with the concept of privacy loss, which we explain next.
+The PRI is an extension to a more fine-grained characterization of the privacy loss in standard DP, thus it does not depart from or replace the standard DP theory.
+We construct PRI so that it is directly connected with the concept of ex-post per-instance privacy loss (Redberg and Wang, 2021), which we explain next in Section 3.2.
 This connection is key to both providing a principled understanding of the new indicator and designing strategies to find a suitable $\epsilon$.
@@
 WLOG, we assume each individual (labeled $1$ to $n$) contributes exactly one record to $x$.
-We denote an $\epsilon$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$.
+We denote an $(\epsilon,\delta)$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$.
 Specifically, we consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers.
-Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and the privacy loss budget $\epsilon$ as input, we define the privacy risk indicator of individual $i\in[n]$ as follows:
+Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon,\delta)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and $(\epsilon,\delta)$ as input, we define the privacy risk indicator of an individual $i\in[n]$ as follows:
 Definition 0 (Privacy Risk Indicator).
 The privacy risk indicator of an individual $i$ is defined as:
-$PRI_{i}=\lVert r-q(x_{-i})\rVert_{1},$
+$PRI_{i}=\lVert r-q(x_{-i})\rVert_{v},$
 where $x_{-i}$ is the neighboring dataset of $x$ without $i$’s record $x_{i}$.
+The value of $v$ depends on the specific mechanism $\mathcal{M}$: if $\mathcal{M}$ is the Laplace Mechanism (Dwork et al., 2014), $v=1$ (i.e., the $\ell_{1}$ norm); if $\mathcal{M}$ is the Gaussian Mechanism (Dwork et al., 2014), $v=2$ (i.e., the $\ell_{2}$ norm).
+We note that one limitation of our formulation of PRI is that it only applies to numeric queries, but not categorical queries that return an element from a set, since both $r$ and $q(x_{-i})$ must be in $\mathbb{R}^{k}$.
+One interesting direction is to explore a more generic formulation of PRI that adapts to both types of queries.
 Connection between PRI and privacy loss.
-Next, we show that when the Laplace Mechanism (Dwork et al., 2014) is used to compute the query output $r$, PRI becomes a relative indicator of individual privacy loss.
-First, we show the definition of privacy loss in DP (Dwork et al., 2014).
-Definition 0 (Privacy Loss).
-Given two neighboring datasets $x,x^{\prime}\in\mathcal{X}^{n}$ such that $\lVert x-x^{\prime}\rVert_{1}\leq 1$, for an output $r\sim\mathcal{Y}$, the privacy loss incurred upon observing $r$ is defined as:
-$\mathcal{L}_{\mathcal{M}(x,q,\epsilon)\|\mathcal{M}(x^{\prime},q,\epsilon)}^{(r)}=\ln\left(\frac{\Pr\left[\mathcal{M}(x,q,\epsilon)=r\right]}{\Pr\left[\mathcal{M}(x^{\prime},q,\epsilon)=r)\right]}\right).$
-Next, we introduce the definition of Laplace Mechanism (Dwork et al., 2014), an $\epsilon$-differentially private mechanism that we use to compute the query output $r$.
-Definition 0 (Laplace Mechanism).
+Next, we show the connection between PRI and ex-post per-instance privacy loss (Redberg and Wang, 2021), which is the privacy loss conditioned on a fixed output, an individual, and a dataset.
+Specifically, we show that when the Laplace Mechanism (Dwork et al., 2014) or the Gaussian Mechanism (Dwork et al., 2014) is used to compute the query output $r$, PRI becomes a relative indicator of ex-post per-instance privacy loss.
+Definition 0 (Ex-post per-instance privacy loss (Redberg and Wang, 2021)).
+For a randomized mechanism $\mathcal{M}$ and a query $q\in\mathcal{Q}$, the ex-post per-instance privacy loss given a realized query output $r\sim\mathcal{Y}$, a fixed dataset $x$, and an individual’s record $x_{i}\in x$ is
+$\epsilon(r,x,x_{-i})=\left|\ln\left(\frac{\Pr[\mathcal{M}(x,q)=r]}{\Pr[\mathcal{M}(x_{-i},q)=r]}\right)\right|$
+Note the difference between $\epsilon$, which is used as input to a DP mechanism, and $\epsilon(\cdot)$, which is a function that denotes the privacy loss conditioned on some given input, i.e., ex-post per-instance privacy loss.
+Intuitively, using ex-post per-instance privacy loss answers the question: what is the realized privacy loss incurred to each individual in the dataset when $r$ is released?
+Note that the ex-post per-instance privacy loss is a more fine-grained characterization of the privacy loss in standard DP; if $\epsilon(\cdot)\leq\epsilon$ for all $r\subseteq Range(\mathcal{M}),x\in\mathcal{X}^{n}$, and $x$’s neighboring dataset $x^{\prime}$ where $\lVert x-x^{\prime}\rVert_{1}\leq 1$, then $\mathcal{M}$ satisfies $\epsilon$-DP (similar argument can be made for $(\epsilon,\delta)$-DP).
+We derive our notion of PRI from the ex-post per-instance privacy loss so that it describes the realized individual privacy loss when an output of a DP mechanism is released.
+The PRI works as a relative indicator of privacy loss for both Laplace Mechanism and Gaussian Mechanism, which we explain next.
+Definition 0 (Laplace Mechanism (Dwork et al., 2014)).
 Given a query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$ and $\epsilon>0$, the Laplace Mechanism is defined as:
 $\mathcal{M}_{L}(x,q,\epsilon)=q(x)+(Z_{1},\ldots,Z_{k}),$
-where $Z_{i}$ are i.i.d random variables drawn from the Laplace distribution with probability density function $p_{b}(z)=\frac{1}{2b}\exp(-|z|/b)$ and the scale $b=\frac{GS_{q}}{\epsilon}$  [^1] The global sensitivity of q in the $\ell_{1}$ norm is defined as $GS_{q}=\max|q(x)-q(x^{\prime})|$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
-Using the Laplace Mechanism, we express the relative privacy loss of individual $i$ as:
-$\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}$ $=\ln\left(\prod_{j=1}^{k}\frac{\exp(-\epsilon|r_{j}-q(x)_{j}|/GS_{q})}{\exp(-\epsilon|r_{j}-q(x_{-i})_{j}|/GS_{q})}\right)$
-$=\frac{\epsilon(\lVert r-q(x_{-i})\rVert_{1}-\lVert r-q(x)\rVert_{1})}{GS_{q}}.$
+where $Z_{i}$ are i.i.d random variables drawn from the Laplace distribution with probability density function $p_{b}(z)=\frac{1}{2b}\exp(-|z|/b)$ and the scale $b=\frac{\Delta_{1}}{\epsilon}$  [^2] The global sensitivity of q in the $\ell_{1}$ norm is defined as $\Delta_{1}=\max\lVert q(x)-q(x^{\prime})\rVert_{1}$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
+The Laplace Mechanism satisfies $\epsilon$-DP.
+When $r$ is released using the Laplace Mechanism, the ex-post per-instance privacy loss is computed as:
+$\epsilon_{L}(r,x,x_{-i})$ $=\left|\ln\left(\prod_{j=1}^{k}\frac{\exp(-\epsilon|r_{j}-q(x)_{j}|/\Delta_{1})}{\exp(-\epsilon|r_{j}-q(x_{-i})_{j}|/\Delta_{1})}\right)\right|$
+$=\left|\epsilon\frac{\lVert r-q(x_{-i})\rVert_{1}-\lVert r-q(x)\rVert_{1}}{\Delta_{1}}\right|.$
 Claim: For all individuals $i,j$ where $i\neq j$, if $PRI_{i}>PRI_{j}$, then
-$\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}>\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-j},q,\epsilon)}^{(r)}.$
-The proof is straightforward: if $PRI_{i}>PRI_{j}$, then $\lVert r-q(x_{-i})\rVert_{1}$ $\lVert r-q(x_{-j})\rVert_{1}$, thus the inequality holds.
-Note that since we are using the Laplace Mechanism, which is $\epsilon$-DP, the privacy loss of any individual will always be bounded by $\epsilon$.
+$\epsilon_{L}(r,x,x_{-i})>\epsilon_{L}(r,x,x_{-j})$
+The proof is straightforward: $PRI_{i}>PRI_{j}$ means $\lVert r-q(x_{-i})\rVert_{1}$ $\lVert r-q(x_{-j})\rVert_{1}$ (the PRI is in $\ell_{1}$ norm since Laplace Mechanism is used), thus the inequality holds.
+Definition 0 (Gaussian Mechanism (Dwork et al., 2014)).
+Given a query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$ and $\epsilon,\delta>0$, the Gaussian Mechanism is defined as:
+$\mathcal{M}_{G}(x,q,\epsilon,\delta)=q(x)+(Z_{1},\ldots,Z_{k}),$
+where $Z_{i}$ are i.i.d random variables drawn from the Gaussian distribution $\mathcal{N}(\sigma^{2})$ with center 00 and variance $\sigma^{2}=\frac{2\Delta_{2}^{2}\ln(1.25/\delta)}{\epsilon^{2}}$  [^3] The global sensitivity of q in the $\ell_{2}$ norm is defined as $\Delta_{2}=\max\lVert q(x)-q(x^{\prime})\rVert_{2}$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
+The Gaussian Mechanism satisfies $(\epsilon,\delta)$-DP.
+When $r$ is released using the Gaussian Mechanism, the ex-post per-instance privacy loss is computed as:
+$\epsilon_{G}(r,x,x_{-i})$ $=\left|\ln\left(\frac{\exp(-\lVert r-q(x)\rVert_{2}^{2}/2\sigma^{2})}{\exp(-\lVert r-q(x_{-i})\rVert_{2}^{2}/2\sigma^{2})}\right)\right|$
+$=\left|\frac{\lVert r-q(x_{-i}\rVert_{2}^{2}-\lVert r-q(x)\rVert_{2}^{2}}{2\sigma^{2}}\right|.$
+Claim: For all individuals $i,j$ where $i\neq j$, if $PRI_{i}>PRI_{j}$, then
+$\epsilon_{G}(r,x,x_{-i})>\epsilon_{G}(r,x,x_{-j})$
+The proof is similar to that of the Laplace Mechanism, except now the PRI is in $\ell_{2}$ norm. Hence $PRI_{i}>PRI_{j}$ means $\lVert r-q(x_{-i})\rVert_{2}$ $\lVert r-q(x_{-j})\rVert_{2}$.
 We now make several remarks:
@@
 The PRI becomes an additional indicator associated with each individual in the dataset that permits controllers to directly observe the effect of releasing a query output.
+The PRI is private information available to controllers only.
 In this paper, we show how PRI helps controllers decide which output to release.
 PRI is data-dependent and thus leaks information about individuals.
-Controllers already have access to individuals’ raw data, so when they gain access to the PRI, they do not learn anything extra about the individuals.
-But this means the PRIs must remain private to the controllers and never be shared with anyone else.
+Controllers already have access to individuals’ raw data, so when they gain access to the PRIs, they do not learn anything extra about the individuals.
+Similar to how controllers would not release the individuals’ raw data, they will not release the PRIs.
+PRIs must remain private to the controllers and never be shared with any other agents.
 Controllers may give PRI different uses.
@@
 4. Deriving Epsilon based on the PRI
-In this section, we describe an algorithm to choose $\epsilon$ given controllers’ privacy preferences in terms of PRIs, thus addressing Challenge 1 (Sections 3.1, 3.2).
-In Section 3.3, we discuss how to bound the total privacy loss in the presence of multiple queries.
-In Section 3.4, we provide the end-to-end DP guarantee for the algorithm.
+In this section, we describe the intuition of using PRIs to express controllers’ privacy preferences (Section 4.1), then introduce an algorithm to choose $\epsilon$ by using PRIs as a more intuitive proxy (Section 4.2).
 We discuss the algorithm’s complexity and extensions in Section 4.3.
@@
 However, choosing $\epsilon$ is challenging because it is hard to understand the specific effect of $\epsilon$ on the privacy of any one individual in the dataset.
-In contrast, the per-individual PRI indicates the relative privacy loss of an individual compared to other individuals in the dataset.
-The consequence is the PRI becomes an actionable indicator that helps controllers decide whether a concrete query output matches their privacy preferences, in which case they may agree to release the output.
-Next, we convey the intuition on how the PRI is useful to controllers through a vignette.
-Figure 1. Example of using PRI to find $\epsilon$. For illustrative purposes, we show the PRI of each student (with the corresponding query output) and leave out students’ actual GPA.
-Vignette illustrating one use of the PRI.
-Figure 1 shows the resulting PRIs of four students after answering a query about their average GPA.
-If the query output were released without using DP (a), the analyst observing the output would learn a lot about those students’ grades, as indicated by the PRI (remember that the PRI is a proxy to privacy loss).
+In contrast, the per-individual PRI indicates the realized privacy loss of an individual relative to other individuals in the dataset.
+The consequence is that the PRI becomes an actionable indicator that helps controllers decide whether a concrete query output matches their privacy preferences, in which case they may agree to release the output.
+Next, we convey the intuition on how the PRI is useful to controllers through a example.
+Figure 2.
+Example of using PRI to find $\epsilon$.
+The query is to compute the average GPA of the three students.
+For each $\epsilon$ (including when no DP is used), we show a specific query output and the corresponding PRI of each student.
+Example illustrating one use of the PRI.
+Figure 2 shows the resulting PRIs of three students after answering a query about their average GPA.
+If the query output were released without using DP (a), the analyst observing the output would learn much about those students’ grades, as indicated by the PRI (recall that the PRI is a proxy to privacy loss).
 If the controller used DP, they would have to choose $\epsilon$ and obtain the query output by applying a DP mechanism.
@@
 We observe that decreasing $\epsilon$ (b, c, d) results in noisier output, which causes the PRIs to become more similar across students.
-In the extreme, when the PRIs are equal (c), no student in the dataset is at a higher risk than others.
-Thus, between (b) and (c), the controller may be more comfortable releasing the specific output (average GPA of 3.2) produced by choosing $\epsilon=1$ because it protects each student equally, which may be a particular privacy preference the controller wants to achieve.
-Lastly, the PRIs in both (c) and (d) indicate no student will be at a higher risk than others, but (d) produces a query output with worse accuracy.
-The controller may choose the larger $\epsilon$ (c) as it yields higher utility for the analyst receiving the output.
-The above is an illustrative example: finding an $\epsilon$ that makes the PRI exactly equal across individuals is generally infeasible: either such an $\epsilon$ does not exist, or it does, but it requires adding too much noise that the output offers little utility to the analyst.
+In the extreme, when the PRIs are equal (c), no student in the dataset is at a higher risk than others (they have the same privacy loss).
+Thus, between (b) and (c), the controller may be more comfortable releasing the specific output (average GPA of 3.275) produced by choosing $\epsilon=1$ because it protects each student equally, which may be a particular privacy preference the controller wants to achieve.
+Lastly, using a smaller $\epsilon=0.1$ (d) also makes the PRIs roughly similar, but at the cost of producing a query output with much worse accuracy than (c).
+One sensible choice is to choose the larger $\epsilon$ (c) as it yields higher utility for the analyst receiving the output.
+The above is an illustrative example that does not work in practice, since finding an $\epsilon$ that makes the PRI exactly equal across individuals is generally infeasible: either such an $\epsilon$ does not exist, or it does, but it requires adding too much noise that the output offers little utility to the analyst.
 4.2. The Find-$\epsilon$-from-PRI Algorithm
-When a controller runs a query on a dataset with a DP mechanism, observing the PRIs associated with individuals in the dataset communicates to the controller the relative privacy loss of each individual compared to others.
-Controllers could then change $\epsilon$, rerun the DP mechanism, and observe the resulting PRIs: they could repeat this process until they find a set of PRIs that represents their preferences.
+When a controller runs a query on a dataset with a DP mechanism, observing the PRIs associated with individuals in the dataset communicates to the controller the realized privacy loss of each individual compared to others if the query output is released.
+Controllers could then change $\epsilon$, rerun the DP mechanism, and observe the resulting PRIs: they could repeat this process until they find a profile of PRIs that represents their preferences.
+Rather than choosing $\epsilon$ and releasing the DP output, controllers first check the corresponding PRIs to understand how would releasing the output affect each individual’s privacy, and then they decide whether or not to release the output based on their privacy preferences on the individuals.
 We introduce an algorithm that implements this intuition, named Find-$\epsilon$-from-PRI.
@@
 We let controllers express their privacy preferences by indicating a cutoff for the highest privacy loss of any individual in the dataset.
-This is done by requesting exactly one parameter from the controllers, a percentage p ($0\leq p\leq 100$).
+This is done by requesting exactly one parameter from the controllers, a percentage $p$ ($0\leq p\leq 100$), which is easier to interpret than $\epsilon$.
 We compute the minimum and maximum PRI, $PRI_{min}$ and $PRI_{max}$.
-Then we find an $\epsilon$ that produces a list of PRIs (and associated DP output) such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$.
+Then we find an $\epsilon$ that produces a DP output with its corresponding PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$.
 In other words, we let controllers specify that the maximum PRI should deviate no more than $p$ percent from the minimum PRI.
 For instance, setting $p=0$ means the PRIs should be exactly equal.
-Unlike $\epsilon$, the parameter $p$ directly conveys the level of risk conferred to individuals after releasing a specific query output.
+Unlike $\epsilon$, the parameter $p$ directly conveys how close the individuals’ realized privacy loss should be — the PRI is a relative indicator of ex-post privacy loss for each individual.
 Smaller $p$ means more privacy at the cost of worse accuracy, hence controllers can trade privacy for accuracy by setting a larger $p$.
-The algorithm aims to find the largest $\epsilon$ that produces a list of PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$.
-Input : Dataset $x=(x_{1},\cdots,x_{n})$; query $q$; percentage $p$
-Output : Chosen privacy parameter $\epsilon$, query output $r$
-1 /* Initialize set of $\epsilon$ values to test */
+The algorithm aims to find the largest $\epsilon$ that produces a list of PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$, because if controllers’ privacy preferences can be satisfied, larger $\epsilon$ is preferred as it leads to more accurate output.
+Input : DP mechanism $\mathcal{M}$; dataset $x=(x_{1},\cdots,x_{n})$; query $q$; percentage $p$
+Output : Query output $r$
+1
+2/* Initialize set of $\epsilon$ values to test */
 3 $\mathcal{E}\leftarrow$ initialize-eps-to-test()
-3 /* Compute $q(x_{-i})$ for all individual $i$ using memoization */
+4
+5/* Compute $q(x_{-i})$ for all individual $i$ using memoization */
 6 $C\leftarrow map()$
@@
 9             $C[x_{i}]\leftarrow q(x_{-i})$
-8
-9 for $\epsilon\in\mathcal{E}$ do
+10
+11
+12for $\epsilon\in\mathcal{E}$ do
+13
 14      /* Compute query output using DP with current $\epsilon$/
-11       $r\leftarrow\mathcal{M}_{L}(x,q,\epsilon)$
+15       $r\leftarrow\mathcal{M}(x,q,\epsilon,\delta=\frac{1}{n^{2}})$
+16
 17      /* Compute corresponding PRI of each individual */
@@
 19       for $x_{i}\in C$ do
+20             if $\mathcal{M}$ is ’Laplace’ then
 21                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{1}$
+22            if $\mathcal{M}$ is ’Gaussian’ then
+23                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{2}$
 24            $PRIs.add(PRI_{i})$
+25
 26      $PRI_{min},PRI_{max}\leftarrow\min(PRIs),\max(PRIs)$
-18       /* Check whether $\epsilon$ satisfies controller’s privacy preference */
-19       if $PRI_{min}/PRI_{max}\geq 1-(p/100)$ then
+27
+28      /* Check whether PRIs satisfy controller’s privacy preference */
+29       if $PRI_{min}/PRI_{max}\geq 1-\frac{p}{100}$ then
 30             return $r$
-21
+31
+return nil
 Algorithm 1 Find-$\epsilon$-from-PRI
 Detailed Algorithm.
-Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output.
+Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output $r$.
 First, the algorithm initializes the set of $\epsilon$ to choose from (line 1).
@@
 Next, the algorithm iterates over the set of $\epsilon$ in descending order (line 1).
-For each $\epsilon$, the algorithm computes the query output $r$ by applying the Laplace Mechanism $\mathcal{M}_{L}$ with parameter $\epsilon$ (line 1) and computes the corresponding PRI of each individual (line 1 - 1).
+For each $\epsilon$, the algorithm computes the query output $r$ by applying the given DP Mechanism $\mathcal{M}$ with parameter $\epsilon$ and a fixed $\delta$ (line 1), then computes the corresponding PRI of each individual (line 1 - 1).
 Then, given the list of PRIs, the algorithm computes $PRI_{min}$ and $PRI_{max}$ (line 1).
 If $PRI_{min}/PRI_{max}\geq 1-(p/100)$, then the algorithm has found a suitable $\epsilon$ and a query output, so it stops and returns the output.
+Note that at this point $\epsilon$ is not released to the analysts, and only the controller knows which $\epsilon$ is chosen by the algorithm.
 Figure 3. Example of $\epsilon$ chosen by the Find-$\epsilon$-from-PRI algorithm with different percentage $p$ over 50 runs
-Running the algorithm with the same dataset $x$, query $q$ and parameter $p$ may identify different $\epsilon$ because the same $\epsilon$ may produce different PRI profiles across executions of the mechanism; simultaneously, similar PRI profiles may be produced by different $\epsilon$ values.
+Running the algorithm with the same dataset $x$, query $q$ and parameter $p$ may identify different $\epsilon$ because the same $\epsilon$ may produce different PRIs across executions of the mechanism; simultaneously, similar PRIs may be produced by different $\epsilon$ values.
 What is important is that the final query output the analyst releases has been subject to DP and the privacy loss of the individuals is consistent with the controller’s privacy preferences as expressed by $p$.
-In addition, we show empirically in Section 6.3.2 that stricter privacy preferences (i.e., smaller $p$) generally lead to smaller $\epsilon$.
-Figure 2 shows a boxplot representing the distributions of $\epsilon$ chosen by the algorithm with different $p$ (5, 25, 50, 75, 95) over 50 runs; for demonstration, we choose one of the five queries (Q5) we have run.
+In addition, we show empirically in Section 6.3.1 that stricter privacy preferences (i.e., smaller $p$) lead to smaller $\epsilon$, as expected.
+For demonstration, we present one of our evaluation results (Q4 in Figure 10).
+Figure 3 shows a boxplot representing the distributions of $\epsilon$ chosen by the algorithm with different $p$ (5, 25, 50, 75, 95) over 50 runs.
 As shown in the figure, as $p$ increases, the algorithm tends to choose larger $\epsilon$.
-Setting $p=5$ means the minimum and maximum PRI needs to be very close (within 5%), thus the algorithm mainly finds $\epsilon$ that is less than 1; in contrast, setting $p=95$ means the maximum PRI can be as much as 95% larger than the minimum PRI, thus the $\epsilon$ found can be very large.
+Setting $p=5$ means the minimum and maximum PRI needs to be very close (within 5%), thus the algorithm mainly finds $\epsilon$ that is less than 0.1; in contrast, setting $p=95$ means the maximum PRI can be as much as 95% larger than the minimum PRI, thus the $\epsilon$ found can be very large.
 Figure 4. Dataflow of Find-$\epsilon$-from-PRI Algorithm
@@
 It is critically important to understand what information is released to whom.
-Figure 3 shows the involved dataflows in a scenario where a controller uses the Find-$\epsilon$-from-PRI algorithm.
-Individuals send their data to the controller; this dataflow remains unchanged from traditional central DP models (Near et al., 2021).
+Figure 4 shows the involved dataflows in a scenario where a controller uses the Find-$\epsilon$-from-PRI algorithm to find $\epsilon$ and release query output to the analyst.
+Individuals send their data to the controller; this dataflow remains unchanged from the traditional central DP model (Near et al., 2021).
 The analyst submits a query to the controller.
-The controller expresses their privacy preferences in terms of $p$ (and not in terms of $\epsilon$) and runs the algorithm to find a suitable query output, which is an output produced by a standard DP mechanism (Laplace Mechanism in our work) with a value for $\epsilon$ selected by the algorithm.
+The controller expresses their privacy preferences in terms of $p$ (and not in terms of $\epsilon$) and runs the algorithm to release a suitable query output, which is an output produced by a DP mechanism (Laplace Mechanism or Gaussian Mechanism in our work) with an $\epsilon$ selected by the algorithm.
 At this point, no information has been returned to the analyst yet.
@@
 The query output that satisfies the controller’s privacy preferences is only released with the controller’s approval.
-The value for $\epsilon$ that led to the query output is also not released by default.
-Releasing $\epsilon$.
-Under certain scenarios (Dwork et al., 2019), $\epsilon$ may also need to be released.
-Since $\epsilon$ is chosen based on PRIs, which are data-dependent, the controller should release $\epsilon$ differential privately by incorporating techniques such as Exponential Mechanism (e.g., Report Noisy Max) (McSherry and Talwar, 2007) or Sparse Vector Technique (Dwork et al., 2009; Lyu et al., 2016) in the algorithm.
-Instead of choosing the largest $\epsilon$ that passes the PRI criterion, the algorithm will choose and return the first $\epsilon$ (which may or may not be the largest) that passes the PRI criterion while preserving DP.
+The value for $\epsilon$ that led to the query output is also not released since $\epsilon$ is chosen based on PRIs, which are data-dependent, so releasing it directly might jeopardize the DP guarantee.
+In Section 5, we explain a modification of the algorithm that permits releasing $\epsilon$ differential privately.
+Claim: Algorithm 1 is $(\epsilon,\delta)$-DP, where $\epsilon$ is the value found by the algorithm that satisfies the PRI threshold, or 0 if no suitable $\epsilon$ is found.
+Algorithm 1 only releases the query output $r$, which is computed using an $(\epsilon,\delta)$-DP mechanism ($\delta=0$ when $\mathcal{M}$ is Laplace Mechanism and $\delta=\frac{1}{n^{2}}$ when $\mathcal{M}$ is Gaussian Mechanism).
+If no suitable $\epsilon$ is found, the algorithm does not release $r$.
+Therefore, the algorithm satisfies $(\epsilon,\delta)$-DP.
+4.3. Discussion and Extensions
+Algorithm complexity and optimizations.
+The algorithm has a worst-case time complexity of $O(nm)$, where $n$ is the number of records in the dataset, and $m$ is the number of $\epsilon$ values we test.
+Computing the PRI of each individual requires computing the query output of each neighboring dataset (line 1 - 1).
+When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$.
+In addition, the query computation can be parallelized since each computation is independent of the other, i.e., the problem is embarrassingly parallel.
+Thus, we implement parallelism and compute the query output for a chunk of records in each process, and we show in Section 6.2 that the algorithm is both efficient and scalable.
+Expressing privacy preferences.
+By default, the algorithm considers the list of PRIs for all individuals in the dataset when finding $\epsilon$.
+A modification of our algorithm permits the controller to specify their privacy preferences in different granularity by selecting subsets of records.
+For example, a controller can divide the dataset $x$ consisting of individuals’ ages to the subset of individuals over 60 ($x^{\prime}$) and those younger ($x^{\prime\prime}$), and specify that the two groups need similar privacy protection.
+This preference can easily be incorporated into the algorithm by letting the controller change the two PRI values to compare instead of the global $PRI_{min}$ and $PRI_{min}$ for all individuals; e.g., they can compute $PRI_{med}^{\prime}$ and $PRI_{med}^{\prime\prime}$, the median PRI for individuals in $x^{\prime}$ and $x^{\prime}$ respectively, and choose an $\epsilon$ that makes $PRI_{med}^{\prime}/PRI_{med}^{\prime\prime}\geq 1-(p/100)$.
+Communicating Privacy-Accuracy Trade-off.
+In some circumstances, the controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual.
+Prior work such as DPComp (Hay et al., 2016) and ViP (Nanayakkara et al., 2022) have explored different ways to visualize the privacy-accuracy trade-off in DP.
+The PRI complements these existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
+5. Releasing $\epsilon$ and Bounding Privacy Loss
+In Section 5.1, we propose a modification of Find-$\epsilon$-from-PRI that finds $\epsilon$ and releases both the query output and $\epsilon$ simultaneously while preserving DP.
+In Section 5.2, we discuss how to bound the total privacy loss in the presence of multiple queries and we provide the end-to-end DP guarantee in Section 5.3.
+5.1. Privately Releasing $\epsilon$
+Since Algorithm 1 finds $\epsilon$ in a data-dependent manner, releasing the chosen $\epsilon$ directly may reveal additional information about the dataset and violate the DP guarantees.
+Therefore, a modification of the algorithm is needed to ensure DP is preserved if the controller wants to release both the query output $r$ and $\epsilon$ to the analysts.
+Key Idea.
+Our key idea is to leverage the Sparse Vector Technique (SVT) (Dwork et al., 2009) to test multiple $\epsilon$ values (from largest to smallest) and release the first $\epsilon$ that satisfies the PRI threshold specified by the controller.
+Algorithm 1 tests the set of $\epsilon$ values and chooses the largest $\epsilon$ that passes the PRI threshold; by incorporating SVT, the modified algorithm chooses approximately the largest $\epsilon$ that passes the threshold so releasing it preserves DP.
+Input : Dataset $x$; a stream of SVT queries $\mathcal{Q}_{svt}=q_{1},q_{2}\cdots$ with global sensitivity at most $\Delta_{svt}$; threshold $\tau$; privacy budget $\epsilon_{svt}^{\prime}$ and $\epsilon_{svt}^{\prime\prime}$
+Output : A stream of answers $a_{1},a_{2}\cdots$ where each $a_{i}\in\{\top,\bot\}$
+1
+[…]
+3for $q_{i}\in\mathcal{Q}_{svt}$ do
+4
+5      if $q_{i}(x)+Lap\left(\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}\right)\geq\tau+\rho$ then
+6             Output $a_{i}=\top$
+7             return
+8      else
+9             Output $a_{i}=\bot$
+10
+Algorithm 2 SVT algorithm proposed by Lyu et al. (Lyu et al., 2016)
+The SVT algorithm.
+Algorithm 2 shows the standard SVT algorithm proposed by  Lyu et al.
+(Lyu et al., 2016)  [^4] The algorithm from  (Lyu et al., 2016) has an additional parameter $c$ (the maximum number of queries the algorithm can answer $\top$).
+We set $c=1$, since we only need one positive answer for our modified algorithm., which we will adopt and incorporate into our modified algorithm to release $\epsilon$ privately.
+SVT takes a dataset, a stream of queries (each with a global sensitivity of at most $\Delta_{svt}$), a given threshold, and a privacy budget $\epsilon_{svt}$ that is divided into $\epsilon_{svt}^{\prime}$ and $\epsilon_{svt}^{\prime\prime}$ (i.e., $\epsilon_{svt}=\epsilon_{svt}^{\prime}+\epsilon_{svt}^{\prime\prime}$), then outputs a stream of answers indicating whether each query output is above or below the threshold.
+The algorithm works by first initializing a Laplacian noise ($\rho$) scaled to $\frac{\Delta_{svt}}{\epsilon_{svt}^{\prime}}$ (line 2).
+For each SVT query, it generates another term of Laplacian noise scaled to $\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}$ and adds the noise to the query output.
+Then the algorithm compares the noisy query output with the noisy threshold (line 2); if it is above or equal to the threshold, the algorithm outputs $\top$ indicating the query output approximately passes the threshold, otherwise, it outputs $\bot$.
+Note that the algorithm answers the SVT queries in a streaming fashion.
+Overall, the algorithm satisfies $\epsilon_{svt}$-DP.
+Input : DP mechanism $\mathcal{M}$; dataset $x=(x_{1},\cdots,x_{n})$; query $q$; variance threshold $\tau$; SVT budget $\epsilon_{svt}$
+Output : Query output $r$, chosen $\epsilon$
+1
+2/* Assign privacy budget for running SVT */
+3 $\epsilon_{svt}^{\prime},\epsilon_{svt}^{\prime\prime}\leftarrow\frac{\epsilon_{svt}}{1+2^{2/3}},\frac{2^{2/3}\epsilon_{svt}}{1+2^{2/3}}$
+4
+5/* Global sensitivity of variance */
+6 $\Delta\leftarrow\frac{1}{n}$
+7
+8/* Generate noise to add to PRI threshold */
+9 $\rho=Lap\left(\frac{\Delta}{\epsilon_{svt}^{\prime}}\right)$
+10
+11/* Initialize set of $\epsilon$ values to test */
+12 $\mathcal{E}\leftarrow$ initialize-eps-to-test()
+13
+14/* Compute $q(x_{-i})$ for each individual $i$ using memoization */
+15 $C\leftarrow map()$
+16 for $x_{i}\in x$ do
+17       if $x_{i}\notin C$ then
+18             $C[x_{i}]\leftarrow q(x_{-i})$
+19
+20
+21for $\epsilon\in\mathcal{E}$ do
+22
+23      /* Compute query output using DP with current $\epsilon$/
+24       $r\leftarrow\mathcal{M}(x,q,\epsilon,\delta=\frac{1}{n^{2}})$
+25
+26      /* Compute corresponding PRI of each individual */
+27       $PRIs\leftarrow list()$
+28       for $x_{i}\in C$ do
+29             if $\mathcal{M}$ is ’Laplace’ then
+30                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{1}$
+31            if $\mathcal{M}$ is ’Gaussian’ then
+32                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{2}$
+33            $PRIs.add(PRI_{i})$
+34
+35      /* Normalize PRIs and compute variance */
+36       $PRI_{max}\leftarrow\max(PRIs)$
+37       for $i\in range(PRIs.length)$ do
+38             $PRIs[i]\leftarrow\frac{PRIs[i]}{PRI_{max}}$
+39      $Var\leftarrow$ compute-variance($PRIs$)
+40
+41      /* Using SVT to release the output $r$ and $\epsilon$ that approximately passes variance threshold */
+42
+43      if $-Var+Lap\left(\frac{2\Delta}{\epsilon_{svt}^{\prime\prime}}\right)\geq-\tau+\rho$ then
+44             return $r,\epsilon$
+45
+46
+return nil
+Algorithm 3 Find-and-release-$\epsilon$-from-PRI
+What is a good SVT query for choosing $\epsilon$?
+To incorporate SVT in our algorithm, each SVT query corresponds to a specific $\epsilon$ the algorithm is testing, and each query should output a value based on PRIs computed from the $\epsilon$; naturally, this value could be the existing PRI criterion we already have in Algorithm 1, which is $\frac{PRI_{min}}{PRI_{max}}$.
+Then the algorithm applies SVT by comparing the PRI criterion to the threshold (i.e., the percentage $p$ specified by controllers) and releasing $\epsilon$ of which the corresponding PRI criterion approximately passes the threshold.
+However, the current PRI criterion has one caveat that makes it difficult to incorporate with SVT: the SVT query, which is a function that outputs $\frac{PRI_{min}}{PRI_{max}}$, has a sensitivity $\Delta_{svt}=1$.
+This means the algorithm will be adding a Laplacian noise scaled to $\frac{2}{\epsilon_{svt}^{\prime\prime}}$ to $\frac{PRI_{min}}{PRI_{max}}$; clearly, when $\epsilon_{svt}^{\prime\prime}$ is small, the magnitude of added noise will be too large and make the PRI criterion useless since it will be too noisy to choose a reasonable $\epsilon$.
+Using variance as the new PRI criterion.
+To address this caveat, we opt for a different PRI criterion when using SVT — by using the variance of normalized PRIs.
+Each SVT query will output the variance of PRIs for the corresponding $\epsilon$, and the variance is compared to a threshold specified by the controller; if the variance is approximately smaller than the threshold, then a suitable $\epsilon$ is found.
+The normalized PRIs will have values between 0 and 1 — this still preserves the PRI’s meaning as a relative indicator of ex-post per-individual privacy loss.
+The variance of the PRIs will have a sensitivity of $\frac{1}{n}$, where n is the number of records in the dataset; this means when $n$ is relatively large, the sensitivity of the variance (hence the Laplacian noise added to the variance) will be small!
+In addition, the variance of PRIs shares the same intuitiveness as $\frac{PRI_{min}}{PRI_{max}}$: if the controllers want the individuals to have similar realized privacy loss (hence more protected), then they should set a lower variance threshold, and vice versa.
+The Find-and-release-$\epsilon$-from-PRI algorithm.
+We present our modified Find-and-release-$\epsilon$-from-PRI algorithm in Algorithm 3.
+The main modification lies in line 3 -  3 and line 3 -  3.
+In line 3, the algorithm allocate the SVT budget $\epsilon_{svt}$ to have $\epsilon_{svt}^{\prime}:\epsilon_{svt}^{\prime\prime}=1:2^{\frac{2}{3}}$; we choose this allocation strategy from  (Lyu et al., 2016) as it leads to better accuracy of SVT.
+Then the algorithm iterates over each $\epsilon$ and computes its corresponding PRIs the same way as in Algorithm 1 (line 3 -  3).
+It then normalizes each PRI and computes the variance (line 3), and lastly, it compares the noisy variance with the noisy threshold (controllers want $Var\leq\tau$, which is equivalent to $-Var\geq-\tau$, and returns the current output $r$ and $\epsilon$ if it passes the threshold.
+The main advantage of using SVT is that it always consumes a fixed budget $\epsilon_{svt}$ regardless of the number of SVT queries.
+Controllers understand the effect of using a specific $\epsilon_{svt}$ and running SVT to release $\epsilon$, since they know exactly how much noise is added to both the variance and the variance threshold $\tau$ they have specified.
+They can decide whether they are comfortable releasing an output that does not exactly conform to their privacy preferences; if not, they may use a different $\epsilon_{svt}$ or $\tau$ and rerun the algorithm.
+Claim: Algorithm 3 is $(\epsilon+\epsilon_{svt},\delta)$-DP, where $\epsilon$ is the value released using SVT, or 0 if no suitable $\epsilon$ is found.
+Algorithm 3 can be divided into two steps.
+In the first step, it uses a $(\epsilon,\delta)$-DP mechanism ($\delta=0$ when the mechanism is Laplace) to compute the query output $r$.
+In the second step, it uses SVT, which is $\epsilon_{svt}$-DP, to output whether the SVT query for the given $\epsilon,r$ is above the threshold, and releases $\epsilon,r$ if it is above the threshold.
+Therefore, by sequential composition (Dwork et al., 2014), Algorithm 3 satisfies $(\epsilon+\epsilon_{svt},\delta)$-DP.
 5.2. Bounding Privacy Loss Across Queries
+The preceding discussion is valid for one query.
 In practice, analysts often want to run multiple queries.
@@
 The solution.
-Instead of quantifying the maximum acceptable privacy loss, which is hard, the algorithm keeps track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ of all previously answered queries; we call it $\epsilon_{c}$.
+Instead of quantifying the maximum acceptable privacy loss, which is difficult, the algorithm keeps track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ or $\epsilon+\epsilon_{svt}$ of all previously answered queries for Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI respectively; we call it $\epsilon_{c}$.
 For an incoming query, the algorithm truncates the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1).
@@
 Considering analysts’ preferences over queries.
-A natural tension exists between analysts and controllers: analysts want more accurate output, whereas controllers want to protect individuals’ privacy by releasing noisy output with degrading accuracy.
+A natural tension exists between analysts and controllers: analysts want more accurate output, whereas controllers want to protect individuals’ privacy by releasing noisy output.
 Because it gets increasingly harder to find a suitable $\epsilon$ after answering multiple queries as $\epsilon_{c}$ keeps increasing, analysts may need to prioritize queries of which they need more accurate output.
-One approach to incorporate analysts’ preferences when finding $\epsilon$ is to allow them to assign weights (i.e., priority scores) to their queries, then controllers can adjust the parameter $p$ in the algorithm based on analysts’ preferences, e.g., they can set a higher $p$ for high-priority queries so the algorithm can find larger $\epsilon$ with more accurate output.
-Nevertheless, solving related problems, such as how to communicate analysts’ preferences to controllers, and how to help controllers decide an appropriate $p$ that takes both analysts’ preferences and their own privacy preferences into account, are outside the scope of the paper but remain interesting directions to explore.
+One approach to incorporate analysts’ preferences when finding $\epsilon$ is to allow them to assign weights (i.e., priority scores) to their queries, then controllers can adjust the parameter $p$ or $\tau$ in each algorithm based on analysts’ preferences, e.g., they can set a higher $p$ or $\tau$ for high-priority queries so the algorithm can find larger $\epsilon$ with more accurate output.
+Nevertheless, solving related problems, such as how to communicate analysts’ preferences to controllers, and how to help controllers decide an appropriate $p$ or $\tau$ that takes both analysts’ preferences and their own privacy preferences into account, are outside the scope of the paper but remain interesting directions to explore.
 5.3. End-to-End DP Guarantee
 Implications of choosing $\epsilon$ adaptively.
-First of all, if the algorithm has only answered one query, releasing the query output to the analyst satisfies DP and the total privacy loss is bounded by the $\epsilon$ that is chosen by the algorithm.
+If the algorithm has only answered one query, releasing the query output to the analyst satisfies DP and the total privacy loss is bounded by the $\epsilon$ or $\epsilon+\epsilon_{svt}$ for Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI respectively.
 However, if the algorithm answered multiple queries (for the same dataset), for each query $\epsilon$ is chosen adaptively because it is chosen based on previously answered queries — a suitable $\epsilon$ needs to be larger than $\epsilon_{c}$.
@@
 Privacy odometer in fully-adaptive composition.
-To answer this question, we construct a privacy odometer (Rogers et al., 2016) for the algorithm, which provides a probabilistic running upper bound on the realized privacy loss after answering each query.
+To answer this question, we construct a privacy odometer (Rogers et al., 2016), which provides a probabilistic running upper bound on the realized privacy loss after answering each query.
 The privacy odometer is proposed to provide privacy guarantees in fully-adaptive composition (Rogers et al., 2016), in which the privacy parameters are chosen adaptively as a function of previous interactions.
-Theorem 3.4 in  (Rogers et al., 2016) has proved that there is no additional leakage to choose privacy parameters adaptively if the privacy bound is computed using sequential composition.
-Therefore, the privacy odometer for the Find-$\epsilon$-from-PRI algorithm under sequential composition (Dwork et al., 2014) can simply be defined as:
+In short, there is no additional leakage to choose privacy parameters adaptively if the privacy bound is computed using sequential composition.
+Therefore, the privacy odometer for Find-$\epsilon$-from-PRI under sequential composition (Dwork et al., 2014) is defined as:
 Definition 0 (Privacy Odometer).
 […]
-The privacy odometer gives the same privacy bound as sequential composition; since the algorithm uses $\epsilon$-DP mechanism, $\delta_{g}$ (the global failure parameter in $(\epsilon,\delta)$-differential privacy) can be set to 0.
-The privacy odometer updates its running upper bound on privacy loss to $\epsilon_{c}$ after the algorithm answers each query, and releasing the composed output of all previously answered queries (on the same dataset) satisfies $\epsilon_{c}$-differential privacy.
-It is worth noting that although  (Rogers et al., 2016) showed the privacy odometer gives a worse bound on privacy loss than the standard advanced composition theorem (Dwork et al., 2010), more recently (Whitehouse et al., 2023) proved that in the fully-adaptive setting the privacy odometer can achieve the same asymptotic bound as advanced composition under certain conditions.
-For simplicity, we define the privacy odometer of our algorithm for sequential composition instead of advanced composition.
-3.5. Discussion and Extensions
-Algorithm complexity and optimizations.
-The algorithm has a worst-case time complexity of $O(nm)$, where $n$ is the number of records in the dataset, and $m$ is the number of $\epsilon$ values we test.
-Computing the PRI of each individual requires computing the query output of each neighboring dataset (line 1 - 1).
-When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$.
-In addition, the query computation can be parallelized since each computation is independent of the other, i.e., the problem is embarrassingly parallel.
-Thus, we implement multi-processing and compute the query output for a chunk of records in each process, and we show in Section 6.3.1 that the algorithm is both efficient and scalable.
-Expressing privacy preferences.
-By default, the algorithm considers the list of PRIs for all individuals in the dataset when finding $\epsilon$.
-A modification of our algorithm permits the controller to specify their privacy preferences in different granularity by selecting subsets of records.
-For example, a controller can divide the dataset $x$ consisting of individuals’ ages to the subset of individuals over 60 ($x^{\prime}$) and those younger ($x^{\prime\prime}$), and specify that the two groups need similar privacy protection.
-This preference can easily be incorporated into the algorithm by letting the controller change the two PRI values to compare instead of the global $PRI_{min}$ and $PRI_{min}$ for all individuals; e.g., they can compute $PRI_{med}^{\prime}$ and $PRI_{med}^{\prime\prime}$, the median PRI for individuals in $x^{\prime}$ and $x^{\prime}$ respectively, and choose an $\epsilon$ that makes $PRI_{med}^{\prime}/PRI_{med}^{\prime\prime}\geq 1-(p/100)$.
-Communicating Privacy-Accuracy Trade-off.
-In some circumstances, the controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual.
-Prior work such as DPComp (Hay et al., 2016) and ViP (Nanayakkara et al., 2022) have explored different ways to visualize the privacy-accuracy trade-off in DP.
-The PRI complements the existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
-4. Helping Analysts Interpret Utility
-In this section, we concentrate on solutions to help analysts interpret the utility of DP output.
-We revisit Challenge 2 (Section 4.1) and present a solution we call utility signaling protocol (Section 4.2).
-4.1. Challenge Overview and Key Insight
-Controllers work on a platform that has access to the individuals’ data and that is responsible for answering analysts’ queries using DP.
-An analyst submits a set of queries $Q=\{q\mid q\in\mathcal{Q}\}$ ($\mathcal{Q}$ is the query space) to the platform, and (if a suitable $\epsilon$ is found for each query) receives a set of noisy query output $S=\{\mathcal{M}(x,q,\epsilon)\mid x\in\mathcal{X}^{n},q\in Q\}$ ($\mathcal{M}$ is a DP mechanism and $\mathcal{X}$ is the data universe).
-A well-known challenge of using DP is that analysts do not know the utility of $S$ (Hotz et al., 2022).
-Concretely, analysts do not know the effect of using $S$ on their downstream task, which is the purpose for analysts to query the data in the first place.
-Example 1/3.
-An analyst queries census data at the county level.
-The set of query output $S$ is aggregate data on population count, education level, and housing prices that are released using DP.
-The analyst uses $S$ to train a binary classifier to predict the income level of each county and obtains an accuracy of 68%.
-The analyst does not know if the (low) accuracy is due to the effect of DP or inappropriate training data.
-The difficulty in interpreting the utility of $S$ makes it challenging for analysts to understand how the DP output will impact their downstream tasks; this is one of the core issues surrounding the discussion of the use of DP for the US Census (Ruggles et al., 2019; Santos-Lozada et al., 2020; Kenny et al., 2021; ABOWD, [n.d.]).
-Observation.
-Analysts use the DP output $S$ for some downstream task, such as a machine learning or inference task.
-Downstream tasks often have an associated evaluation function, $g(\cdot)\rightarrow\mathbb{R}$, which returns a number indicating the quality of the downstream task when evaluated on a certain input.
-In the example above, the downstream task is to train a binary classifier, the evaluation function measures the quality of the classifier trained on DP output and returns 68% as the quality indication (accuracy in this case).
-Insight: signaling DP effect.
-Analysts would learn the impact of DP if they had access to both the result of the evaluation function on the DP output, $g(S)$ [^2] We slightly abuse notation for simplification.
-and the raw output, $g(S^{\prime})$, where $S^{\prime}=\{q(x)\mid x\in\mathcal{X}^{n},q\in Q\}$ (the set of query output without using DP).
-Example 2/3.
-Besides having access to the quality of the classifier on the DP census aggregates (68%), the analyst also gains access to the quality of the classifier if it had been trained on the raw census aggregates $S^{\prime}$, say 80%.
-By comparing 68% ($g(S)$) and 80% ($g(S^{\prime})$), the analyst learns the effect of added DP noise on their downstream task.
-Gaining access to $g(S)$ and $g(S^{\prime})$ helps analysts interpret the impact of DP on their downstream tasks.
-The remaining challenge is how to provide $g(S^{\prime})$ to the analysts without giving them access to $S^{\prime}$, which is the raw query output and cannot be released to the analyst without violating the privacy of individuals in the dataset.
-4.2. The utility signaling protocol
-Rather than the platform directly sending $S^{\prime}$ to the analyst, which would break individuals’ privacy, the analyst submits their downstream task and associated evaluation function $g(\cdot)$ to the platform.
-Upon successful inspection by the controller, the platform executes the downstream task and evaluation function on behalf of the analyst, producing $g(S)$ and $g(S^{\prime})$.
-Last, if the controller approves, the platform releases both $g(S)$ and $g(S^{\prime})$ to the analyst, thus signaling the impact of DP on the analysts’ task without revealing $S^{\prime}$.
-Example 3/3.
-Instead of only submitting the queries to the platform, with the utility signaling protocol, the analyst submits their queries, the downstream task, and the evaluation function.
-The controller inspects each and allows the platform to signal the impact of using DP to the analyst by releasing both utility results 68% and 80%.
-4.2.1. Protocol Step by Step
-Figure 4. End-to-end example of utility signaling protocol. We illustrate the protocol in the central DP deployment scenario where the controller has access to the raw data.
-Figure 4 illustrates the utility signal protocol, which we describe next:
-0.
-Platform.
-The platform has access to the dataset $x$ of individuals.
-It also has the capacity to run queries using DP.
-It may optionally implement the Find-$\epsilon$-from-PRI algorithm presented in Section 3.2333While we envision the algorithm being deployed to help controllers, both contributions are orthogonal.
-1. Analyst submits queries to the platform. The platform notifies the controller.
-2.
-Controller inspects queries.
-The controller inspects the analyst’s queries and allows the platform to execute the queries as long as a suitable $\epsilon$ is selected.
-The platform withholds both the set of DP output $S$ and raw output $S^{\prime}$ without releasing anything to the analyst by default.
-3.
-Platform releases $S$ upon controller’s approval.
-The platform notifies the controller after computing $S$ and $S^{\prime}$.
-The controller gives permission for the platform to release $S$ to the analyst.
-Up to this point, the protocol provides the analyst with the same information they can access in a typical central DP deployment.
-Now, we extend the protocol to implement the utility signaling component to help the analyst understand the impact of DP on their downstream task.
-4. Analyst submits downstream task and evaluation function $g(\cdot)$ to the platform. The platform notifies the controller.
-5.
-Controller inspects downstream task and evaluation function.
-The controller inspects the implementation of the analyst’s downstream task and evaluation function $g(\cdot)$, and decides whether it is appropriate to release $g(S)$ and $g(S^{\prime})$.
-With the controller’s approval, the platform executes the downstream task and evaluation function $g(\cdot)$ on both $S$ and $S^{\prime}$, withholding the results by default.
-6.
-Platform releases $g(S)$ and $g(S^{\prime})$ upon controller’s approval.
-The controller can inspect $g(S^{\prime})$ one last time to ensure it is appropriate to release; $g(S)$ preserves DP since it is a post-processing (Dwork et al., 2014) step of DP output $S$.
-If approved, the platform releases $g(S)$ and $g(S^{\prime})$ to the analyst, who learns the impact of DP on their downstream task without ever gaining access to $S^{\prime}$ directly.
-4.2.2. Privacy implications of releasing $g(S^{\prime})$
-Any data release leaks information.
-When releasing an output that is subject to DP, such as $g(S)$, controllers understand the consequences on the individual’s privacy.
-Conversely, when releasing output not subject to DP, such as $g(S^{\prime})$, controllers must assess whether doing so maintains individuals’ privacy.
-That ultimately depends on the dataset $x$, the output $S^{\prime}$, the implementation of the downstream task and evaluation function, $g(\cdot)$, and the adversary model.
-The scope of the utility signaling protocol is to enable analysts to understand the impact of using DP, but not on understanding, controlling, and determining when it is safe to release $g(S^{\prime})$, which is ultimately at the discretion of the controllers and an interesting avenue of future work.
-5. Escrow-Based DP Platform
-We curate a set of requirements for a DP platform that implements the Find-$\epsilon$-from-PRI algorithm and the utility signaling protocol in Section 5.1.
-In Section 5.2, we describe the implementation of a platform built on top of a data escrow that meets these requirements.
-5.1. Platform Requirements
-The solutions to Challenge 1 (the Find-$\epsilon$-from-PRI algorithm), to Challenge 2 (the utility signaling protocol), and the nature of processing over private data call for a platform that meets the following requirements:
-  • R1.
-Control dataflows: Given the three relevant entities (analyst, platform, and controller), a dataflow indicates data movement between any two of those entities.
-A requirement for the platform is to ensure that only permitted dataflows take place.
-Concretely, controllers indicate what dataflows take place, and the platform enforces the controller’s decision.
-In the context of our work, analysts can access for example the DP output $S$ after the controller approves, and they cannot access PRIs or any other intermediates not explicitly approved by the controller.
-  • R2.
-Secure platform: Adversarial agents may break the requirement of controlling dataflows by breaking into the platform and forcing unintended dataflows.
-This is an assumed risk in the central DP model, which can be reduced by using a local DP deployment (Near et al., 2021).
-Unfortunately, local DP reduces the accuracy of the output compared to central DP.
-A requirement for the platform is to ameliorate the risk of unintended dataflows.
-  • R3.
-Protect against adversarial downstream and evaluation functions: Adversarial analysts could submit downstream tasks and evaluation functions crafted to release more information than needed.
-While controllers must inspect these functions before agreeing to release the output and they are the ultimate arbiter, a requirement of the platform is to facilitate the controllers’ job.
-  • R4.
-High Performance: The Find-$\epsilon$-from-PRI algorithm increases processing overhead.
-Simultaneously, security safeguards to address R2 are often at odds with high-performance computation.
-Together, these two facts risk making the platform impractical.
-A requirement is to implement the solutions to Challenge 1, Challenge 2, and meet the requirements R1, R2, and R3 while maintaining low overhead.
-We now explain the implementation of a platform that addresses these three requirements.
-5.2. Platform Design and Implementation
-To demonstrate the feasibility of the solutions to Challenge 1 and Challenge 2, we describe how to implement the solutions above on a platform that leverages a data escrow (Xia et al., 2022) architecture.
-Data escrow overview.
-A data escrow is a system that acts as an intermediary for storing, processing, and releasing data.
-Different entities interact within the escrow.
-Controllers register data and control the dataflows.
-The data is never stored or computed in the clear using end-to-end encryption.
-Analysts run computations on registered data in the escrow.
-For any data or its derived data product to leave the escrow, controllers have to create a policy on the escrow.
-The policies are tamper-proof since the escrow validates the identity of the controller who creates the policy and the policy remains encrypted, thus ensuring the controller fully controls the dataflow.
-Meeting R1.
-No raw data (e.g., $x$) or derived data product (e.g., PRIs, $g(S)$, $g(S^{\prime})$) leave the escrow without an explicit policy that permits that.
-Policies then determine what dataflows take place.
-Controllers are in charge of writing policies.
-Meeting R2. Data is encrypted end-to-end, even during computation. Adversaries that break into the platform cannot read plain-text data. In other words, no unintended dataflows take place.
-Meeting R3.
-The platform can let analysts submit their downstream task and associated evaluation function, and put the burden on the controller to verify the implementation of those functions.
-Or it can offer standard implementation of downstream tasks and evaluation functions and let analysts call these, thus requiring scrutinizing the functions once and using them many times.
-The data escrow permits the deployment of programs expressing a desired functionality.
-In specific DP scenarios, e.g., the US Census, it is conceivable that one could provide programs implementing (at least) the more standard functionalities and query interfaces, and rely on controllers to scrutinize only one-off tasks.
-Meeting R4.
-End-to-end data encryption (necessary to meet R2) can be achieved in many ways; for instance, DP platforms leveraging Secure Multiparty Computation (Narayan and Haeberlen, 2012; Bater et al., 2016, 2018; Wang et al., 2022; Roth et al., 2019), Homomorphic Encryption (Agarwal et al., 2018; Roy Chowdhury et al., 2020), Oblivious RAM (Bogatov et al., 2021) and Trusted Execution Environment (Nguyen et al., 2020).
-Our evaluation prototype uses AMD Secure Encrypted Virtualization (Kaplan et al., 2016) because it permits running unmodified software while maintaining the integrity and confidentiality of memory contents.
-The escrow design protects access to external resources (e.g., network and disk), ensuring data is encrypted transparently to any application.
-As we show in the evaluation (Section 6.2), this solution virtually removes the processing overhead.
-In addition, we implement a series of optimizations to the Find-$\epsilon$-from-PRI algorithm to demonstrate its efficiency and scalability when run in the escrow (Section 6.3.1).
+The privacy odometer for Find-and-release-$\epsilon$-from-PRI can be defined using the same principle, which we omit.
+Essentially, the privacy odometer updates its running upper bound on privacy loss to $\epsilon_{c}=\sum_{i=1}^{i=k}\epsilon_{i}$ after the algorithm answers the $k$-th query, and releasing the composed output of all previously answered queries (on the same dataset) satisfies $(\epsilon_{c},\delta_{g})$-DP (we can set $\delta_{g}=\sum_{i=1}^{i=k}\delta$).
+It is also worth noting that under certain conditions (Whitehouse et al., 2023), the privacy odometer achieves the same asymptotic bound as advanced composition (Dwork et al., 2010).
+For simplicity, we define the privacy odometer of our algorithm for sequential composition instead of advanced composition, and we leave further improvements for future work.
 6. Evaluation
-Our evaluation shows our solutions address Challenge 1 to help controllers, and Challenge 2 and 3 using an escrow-based platform. We answer the following research questions:
+We answer the following research questions:
   • RQ1: Does presenting the PRI help data controllers choose $\epsilon$?
-[…]
-  • Microbenchmarks: We conduct experiments to show that
-    Find-$\epsilon$-from-PRI algorithm is efficient and scalable. In addition, we study the effect of varying parameter $p$ in the algorithm.
-6.1. RQ1: Using PRI To Choose Epsilon
+  • RQ2: Are our proposed algorithms, Find-$\epsilon$-from-PRI and
+    Find-and-release-$\epsilon$-from-PRI efficient and scalable?
+  • Microbenchmarks: We study the effect of varying parameter $p$ in Find-$\epsilon$-from-PRI, and parameter $\tau$ in Find-and-release-$\epsilon$-
+    from-PRI
+    , and we compare the chosen $\epsilon$ values by both algorithms.
+6.1. RQ1: Does PRI Help Controllers Choose $\epsilon$?
 To determine whether PRI helps the data controllers choose $\epsilon$, we conducted an IRB-approved user study to analyze how PRI affects the controller’s decision-making regarding choosing $\epsilon$.
@@
 6.1.1. Recruitment and data collection
-We recruited 100 participants using Prolific (pro, [n.d.]).
+We recruited 100 participants using Prolific (pro, [n.
+d.]).
 To complete the study, each participant must sign a consent form, which describes the purpose and content of the study, compensation, and data confidentiality.
@@
 First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group, and 27 are in the treatment group.
-The survey presents a sample from the UCI Adult dataset (uci, [n.d.]) and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$.
+The survey presents a sample from the UCI Adult dataset (uci, [n.
+d.]) and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$.
 Then, it presents two queries in random order:
@@
 Note that this is the only change between the control and treatment groups to ensure we isolate the intervention.
-A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include these to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
+A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include it to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
 6.1.3. Results
@@
 Participants in the treatment group tend to choose either 2 or 1.5, regardless of their familiarity with DP, whereas participants in the control group chose $\epsilon$ with greater variance.
-This suggests participant’s a priori knowledge of DP does not transfer to making better choices of $\epsilon$; on the other hand, showing PRIs helped participants make more consistent choices whether or not they know about DP previously.
+This suggests participant’s a priori knowledge of DP does not transfer to making better choices of $\epsilon$; on the other hand, showing PRIs helped participants make more consistent choices whether or not they knew about DP previously.
 How do different ways of presenting PRIs affect participants’ decisions when choosing $\epsilon$?
@@
 We explored an alternative way of presenting PRIs by using histograms.
-To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and asked them to complete the same survey questions in the treatment group, with the only change being presenting histograms of PRIs instead of range plots.
+To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and we asked them to complete the same survey questions as in the treatment group except that the survey presents histograms of PRIs instead of range plots.
 Figure 7. $\epsilon$ chosen by participants who were presented with histograms or range plots, or in the control group
@@
 We also randomized the two questions for choosing $\epsilon$.
-6.2. RQ2: DP On Escrow-based Platform
-[…]
-6.2.1. Baselines for comparison
-We evaluate and compare the performance of the four following baselines:
-  • Data Escrow: We implemented an escrow application that takes a SQL query as input and applies the Laplace Mechanism to release the query output.
-Optionally, if the parameter $\epsilon$ is not given, it will use the Find-$\epsilon$-from-PRI algorithm to select a suitable $\epsilon$.
-We simulate the two entities (controller and analyst) in the escrow.
-A policy is first created by the controller to allow the query output to be released to the analyst.
-[…]
-  • Standalone with DP: This baseline runs the given query and applies the Laplace Mechanism to the query output in a central platform.
-The dataset is in plaintext.
-This represents the central DP deployment scenario without the overhead of additional security measures.
-  • Standalone without DP: This baseline simply runs the given query without DP in a central platform. The dataset is in plaintext. We include this baseline to show the overhead of using DP.
-[…]
-6.2.2. Setup
-We ran the experiments on n2d-standard-8 instances with 8 virtual CPUs (AMD Rome) and 64GB memory from Google Compute Engine.
-We used AMD Secure Encrypted Virtualization (SEV) when running the data escrow.
-We used Ubuntu 20.04 and Python 3.8 for all baselines.
-6.2.3. Dataset
-[…]
-6.2.4. Queries
-We run the 3 queries independently for all baselines:
-  • Q1: Number of people between age 20 and 30
-  • Q2: Number of people by each race
-  • Q3: Number of females who have income greater than 50K
-6.2.5. Results
-[…]
-Figure 8. Average query time (in log scale) over 10 runs
-[…]
-[…]
-6.3. Microbenchmarks
-We evaluate the runtime and scalability of the Find-$\epsilon$-from-PRI algorithm in Section 6.3.1 and study how different values of the parameter $p$ affect the $\epsilon$ chosen in Section 6.3.2.
-6.3.1. Is Find-$\epsilon$-from-PRI practical?
-Since the algorithm can run multiple times if there are multiple queries , the algorithm needs to be efficient so running it multiple times will still be practical.
-We ran the algorithm in the data escrow to further demonstrate its performance for computation-heavy tasks.
+6.2. RQ2: Are Both Algorithms Practical?
+Since both Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI algorithms can run multiple times if there are multiple queries, the algorithms need to be efficient so running it multiple times will still be practical.
+Thus, we evaluate both algorithms’ runtime for different queries and dataset sizes.
 Datasets.
-We controlled the experiment by generating datasets of size 1000, 10000, 100000, and 1 million based on the Adult dataset.
-We sampled records for datasets of sizes 1000 and 10000.
+We controlled the experiment by generating datasets with 1000, 10000, 100000, and 1 million records based on the Adult dataset (uci, [n.
+d.]).
+We sampled records for datasets of sizes (number of records) 1000 and 10000.
 For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they were not the same copy of each other.
 We included all attributes in the dataset.
-Queries.
-We ran the algorithm to find $\epsilon$ for the five SQL queries below.
-We chose a diverse set of queries with attributes of different domain sizes and types.
-The queries are run independently.
+Queries. We ran both algorithms for the five SQL queries below. We chose a diverse set of queries with attributes of different domain sizes and types. The queries are run independently.
   • Q1: SELECT COUNT(*) FROM adult WHERE income==’>50K’ AND education_num==13 AND age==25
@@
   • Q4: SELECT AVG(hours_per_week) FROM adult WHERE workclass in (’Federal-gov’, ’Local-gov’, ’State-gov’)
-  • Q5: SELECT SUM(fnlwgt) FROM adult WHERE capital_gain>0 AND income==’<=50K’ AND occupation==’Sales’
+  • Q5: SELECT SUM(capital_gain) FROM adult
 Parameters.
 We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,\\ 0.099,\cdots,0.001]$ (37 values in total).
-We set $p=50$.
-We used six parallel processes for computing the PRIs.
-Figure 9. Average runtime (in log scale) to find $\epsilon$ over 10 runs
+We set $p=5$ for Find-$\epsilon$-from-
+PRI
+, and $\epsilon_{svt}=1,\tau=10^{-6}$ for Find-and-release-$\epsilon$-from-PRI. We used eight parallel processes for computing the PRIs.
+Configurations. We ran the experiments on a Chameleon Cloud (cha, [n. d.]) instance with 160 CPUs (Intel Xeon Platinum 8380) and 256GB memory. We used Ubuntu 22.04 and Python 3.10 for all experiments.
+Figure 8. Average runtime (in log scale) to find $\epsilon$ over 10 runs by Find-$\epsilon$-from-PRI (a) and Find-and-release-$\epsilon$-from-PRI (b)
 Results.
-We first note that all queries on all 10 runs were able to find a suitable $\epsilon$ on every dataset we tested.
-Figure 9 shows a line plot where the x-axis is the data size and the y-axis is the runtime, both in log scale.
-Each line represents a query, with the error band showing the standard deviation.
-As shown in the figure, the runtime increases linearly with the data size.
-Most queries took less than a few minutes to finish, even for 1 million records, except Q5, which took more than an hour.
-The runtime for Q5 is slower since it involves the attribute fnlwgt, which is a numerical attribute with a large domain size and many unique values, thus the query involves significantly more unique records than the other queries.
-For example, for 1 million records, Q5 has approximately 100000 unique records of which the PRIs need to be computed; whereas for other queries, the algorithm only needs to compute PRIs for no more than a few thousand records and the rest can be directly retrieved using memoization.
-6.3.2. Effect of parameter $p$
-The Find-$\epsilon$-from-PRI algorithm finds a suitable $\epsilon$ by letting controllers specify their privacy preferences through a parameter $p$, and checks whether $PRI_{min}/PRI_{max}\geq 1-(p/100)$ for each $\epsilon$.
+Figure 8 shows two line plots where the x-axis is the data size and the y-axis is the runtime, both in log scale; the left plot corresponds to the run time of Find-$\epsilon$-from-PRI, and the right plot corresponds to the run time of Find-and-release-$\epsilon$-from-PRI.
+Each line in the plot represents a query, with the error band showing the standard deviation.
+As shown in both plots, the runtime of both algorithms increases linearly with the data size.
+Most queries took less than a few minutes to finish, even for 1 million records.
+The runtime of Find-and-release-$\epsilon$-from-PRI is slightly longer than that of Find-$\epsilon$-from-PRI, especially when the dataset size is small.
+This is because applying SVT in Find-$\epsilon$-from-PRI takes an extra overhead since it needs to normalize the PRIs, compute the variance, and generate two Laplacian noise.
+However, when the dataset gets larger, the overhead becomes negligible and the runtime is dominated by the time to compute PRIs.
+6.3. Microbenchmarks
+Figure 9. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$. The mechanism used is Laplace.
+Figure 10. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$. The mechanism used is Gaussian.
+6.3.1. Effect of parameter $p$ in Find-$\epsilon$-from-PRI
+The Find-$\epsilon$-from
+-PRI
+algorithm finds a suitable $\epsilon$ by letting controllers specify their privacy preferences through a parameter $p$, and checks whether $PRI_{min}/PRI_{max}\geq 1-(p/100)$ for each $\epsilon$.
 Smaller $p$ means stricter privacy preferences because the minimum and maximum PRI need to be closer.
@@
 Setup.
-We use the Adult dataset with 100000 rows that we generated in Section 6.3.1.
-We ran each of the five queries from Section 6.3.1 for 50 times, and report the $\epsilon$ chosen by our algorithm in each run.
-Figure 10. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$
+We use the Adult dataset with 100000 rows that we generated in Section 6.2.
+We ran each of the five queries from Section 6.2 50 times, and recorded the $\epsilon$ chosen by the algorithm in each run.
 Results.
-Figure 10 shows boxplots of $\epsilon$ found by the algorithm for different $p$.
-As shown in the figure, setting larger $p$ tends to find larger $\epsilon$.
-Setting $p=5$ results in very small $\epsilon$ chosen for all five queries because the maximum PRI can only be no more than 5% from the minimum PRI; most $\epsilon$ values are less than 0.1 for Q1–4 and less than 1 for Q5.
-Queries with larger sensitivity (Q2, Q4, Q5) can find larger $\epsilon$ when $p>50$, and queries with sensitivity of 1 (Q1, Q3) tend to find smaller $\epsilon$.
-Overall, the results empirically demonstrate that the algorithm is useful for finding a suitable $\epsilon$ (and corresponding query output) that is consistent with the controller’s privacy preferences.
+Figure 9 and Figure 10 show the boxplots of $\epsilon$ found by the algorithm with different $p$ for each query; Figure 9 shows the results when the DP mechanism $\mathcal{M}$ used in the algorithm (for computing the query output $r$ and the PRIs) is Laplace Mechanism, and Figure 10 shows the results for the Gaussian mechanism.
+As shown in both figures, setting larger $p$ tends to find larger $\epsilon$.
+Setting $p=5$ results in very small $\epsilon$ chosen for all five queries because the maximum PRI can only be no more than 5% from the minimum PRI; most $\epsilon$ values are less than 0.1 when the mechanism is Laplace, and less than 1 when the mechanism is Gaussian.
+In fact, for most of the queries and $p$, using Gaussian caused the algorithm to choose larger $\epsilon$ compared to using Laplace, because the Gaussian Mechanism tends to produce noisier output than the Laplace Mechanism  [^5] This holds when the query’s $\ell_{2}$ sensitivity (used by Gaussian Mechanism) is similar to its $\ell_{1}$ sensitivity (used by Laplace Mechanism); otherwise, if the query’s $\ell_{2}$ sensitivity is much smaller, Gaussian Mechanism adds less noise compared to Laplace Mechanism.
+using the same $\epsilon$.
+In addition, both algorithms find larger $\epsilon$ when $p>50$ for queries with larger sensitivity (Q2, Q4, Q5), and they tend to find smaller $\epsilon$ for queries with sensitivity of 1 (Q1, Q3).
+Overall, the results empirically demonstrate that both algorithms are useful for finding a suitable $\epsilon$ that is consistent with the controller’s privacy preferences.
+Figure 11. $\epsilon$ found by Find-and-release-$\epsilon$-from-PRI algorithm over 50 runs, with $\tau=10^{-3},10^{-6},10^{-9},10^{-12}$.
+6.3.2. Effect of parameter $\tau$ in Find-and-release-$\epsilon$-from-PRI
+The Find-and-release-$\epsilon$-from-PRI algorithm lets controllers specify the PRI variance threshold $\tau$ (and the SVT budget $\epsilon_{svt}$), and it uses SVT to choose the first $\epsilon$ of which the variance of its corresponding PRIs is approximately smaller than the threshold.
+Smaller $\tau$ means stricter privacy preferences.
+We test the effect of varying $\tau$ in the algorithm; specifically, when $\tau=10^{-3},10^{-6},10^{-9},10^{-12}$.
+Setup. We use the same setup as in Section 6.3.1. We set $\epsilon_{svt}=1$ and we use Laplace Mechanism as the DP mechanism $\mathcal{M}$.
+Results.
+Figure 11 shows the boxplots of $\epsilon$ found by the algorithm with different $\tau$ for each query.
+The $\epsilon$ values selected are virtually similar and small when $\tau=10^{-6},10^{-9},10^{-12}$.
+On the contrary, the algorithm mostly selects the largest $\epsilon$ (10) it is considering when $\tau=10^{-3}$ for queries Q1, Q2, Q4, and Q5, and it selects a relatively small $\epsilon$ that is less 1 for Q3.
+This suggests controllers may need to observe the PRIs to decide a reasonable variance threshold to release $\epsilon$ and the query output $r$.
+As we noted before in Section 5.1, controllers may also rerun the algorithm multiple times with different parameters $\tau$ and $\epsilon_{svt}$ to observe how much noise is added and decide what to release that conforms to their privacy preferences.
 7. Related Work
-Defining PRI.
-The PRI shares similar intuition as Personalized Differential Privacy (PDP) (Jorgensen et al., 2015) in that different individuals may need different levels of protection since some records are more sensitive than others.
-In contrast to PDP, which requires each individual to specify their own privacy requirement through a privacy budget, the PRI is computed by controllers to indicate the realized privacy loss of each individual if a DP output were to be released.
+Privacy loss accounting.
+Various DP definitions have been proposed to account the privacy loss in a data-dependent manner or in different granularity.
+Our paper defines PRI as a relative indicator of the ex-post per-instance privacy loss (Redberg and Wang, 2021), which is built upon previous works that separately proposed ex-post privacy loss (Ligett et al., 2017) and per-instance privacy loss (Wang, 2019); ex-post privacy loss is the privacy loss conditioned on the realized output of a DP mechanism, and per-instance privacy loss is the privacy loss conditioned on a fixed dataset and a specific record in it.
+In addition,  Feldman and Zrnic (Feldman and Zrnic, 2021) proposed a privacy filter under Renyi differential privacy, which tracks the personalized privacy loss for each individual and drops an individual’s record when their personalized privacy loss exceeds a fixed budget (for all individuals).
+Seeman et al.
+(Seeman et al., 2024) proposed a formulation of DP called Per Record Differential Privacy; in this formulation, only the policy function on each individual’s privacy loss is published but not the actual privacy loss value, and individuals can compute their own privacy loss using the function.
+Unlike most works which define and use a different DP formation than standard DP, our work does not try to replace standard DP — we simply use the notion of ex-post per-instance privacy loss as a tool to connect our PRI definition to privacy loss in a more fine-grained granularity, and both our algorithms, Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI, satisfy $(\epsilon,\delta)$-DP.
 Choosing privacy loss budget.
@@
 Nanayakkara et al.
-visualizes the disclosure risk in relation to $\epsilon$ using the definition of disclosure risk from Lee and Clifton, which is the probability of disclosing the presence or absence of any individual in the database.
+(Nanayakkara et al., 2022) visualizes the disclosure risk in relation to $\epsilon$ using the definition of disclosure risk from Lee and Clifton (Lee and Clifton, 2011), which is the probability of disclosing the presence or absence of any individual in the database.
 John et al.
@@
 Ligett et al.
-empirically finds the smallest $\epsilon$ consistent with an accuracy requirement in the private empirical risk minimization setting.
+(Ligett et al., 2017) empirically finds the smallest $\epsilon$ consistent with an accuracy requirement in the private empirical risk minimization setting.
 Ge et al.
-translates a user-specified accuracy bound to a suitable DP mechanism, along with an $\epsilon$ that incurs the least privacy loss while satisfying the accuracy bound.
+(Ge et al., 2019) translates a user-specified accuracy bound to a suitable DP mechanism, along with an $\epsilon$ that incurs the least privacy loss while satisfying the accuracy bound.
 Hsu et al.
-proposes an economic model for choosing $\epsilon$ by balancing the interests of the data analyst and the prospective participant who will participate if the benefit outweighs the risk.
+(Hsu et al., 2014) proposes an economic model for choosing $\epsilon$ by balancing the interests of the data analyst and the prospective participant who will participate if the benefit outweighs the risk.
 However, this model assumes that the participant can quantitatively express the ”cost” of whether to participate in the study, which is impractical for non-experts.
 Cummings et al.
-and Xiong et al.
-investigate users’ understanding of DP guarantees through multiple user studies, which offer insights on how to communicate the privacy implications of DP to users.
+(Cummings et al., 2021) and Xiong et al.
+(Xiong et al., 2020) investigate users’ understanding of DP guarantees through multiple user studies, which offer insights on how to communicate the privacy implications of DP to users.
 To the best of our knowledge, no known approach helps users understand the implications of $\epsilon$ on individuals’ privacy, and use that to design an algorithm that finds $\epsilon$ automatically.
-Interpreting utility.
-Tao et al.
-proposed DPXPlain to explain aggregate query answers by comparing answers with and without DP and assess whether the difference is due to the noise introduced by DP or properties of the data, thus helping analysts interpret whether the answers are useful for their analysis.
-This idea, complementary to ours, relies on analysts to come up with comparison questions that require explanation in order to understand any anomalies or trends in the query answers.
-CDP LDP Shuffler
-Crypto
-Escrow
-T
-No Yes Yes Yes
-Yes
-Add.
-Trust
-Asm.
-Trust
-individuals
-add noise
-correctly
-Trust
-shuffler
-Trust
-crypto
-impl.
-Trust
-SEV &
-escrow
-impl.
-Encr.
-Data?
-No
-Does not
-collect
-raw data
-Does not
-collect
-raw data
-Yes Yes
-Error $O(1)$ $O(\sqrt{n}$) Near $O(1)$ $O(1)$ $O(1)$
-Table 1. Comparisons of different DP deployment scenarios. T stands for trustworthy.
-Deploying differential privacy.
-The threat model of the central DP model assumes adversaries cannot breach the platform and force unintended data flows, which is unrealistic.
-Alternatives beyond the central model aim to deploy DP with different trust assumptions.
-We summarize those different deployment scenarios in Table 1.
-The local model (LDP) has been proposed to eliminate the need for a trusted platform and it has been adopted in various real-world applications for collecting client data from their software and devices  (Erlingsson et al., 2014; Ding et al., 2017; Differential Privacy Team, 2017).
-In LDP, each individual perturbs their own data using LDP mechanisms before sending it to the platform; adversaries can control the platform but they only gain access to DP data.
-However, LDP also introduces a much larger error ($O(\sqrt{n})$, where n is the number of individuals) than the central DP model ($O(1)$).
-The shuffle model creates an anonymous communication channel that amplifies the privacy guarantees  (Kasiviswanathan et al., 2011; Balle et al., 2018), thus it requires adding less noise when users perturb their data and it offers better accuracy than the local DP model.
-However, it requires additional trust assumptions as the shuffler must be trusted.
-Lastly, techniques that leverage cryptographic primitives, specifically secure multiparty computation, homomorphic encryption, oblivious RAM, and trusted hardware enclaves, have been proposed to offer security guarantees when deploying DP without a trusted data curator  (Dwork et al., 2006a; Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018; Nguyen et al., 2020; Bittau et al., 2017; Bogatov et al., 2021).
-We showed that the escrow-based solution we introduce presents an interesting alternative with a lower computational cost.
+Customizing Privacy.
+Some works study methods to allow users to specify privacy requirements at the individual level.
+Jorgensen et al.
+(Jorgensen et al., 2015) proposes Personalized Differential Privacy in which each individual specifies their own privacy preferences instead of using a global privacy loss budget.
+Since choosing a numerical privacy preference value is challenging for typical users, the authors recommended associating suitable privacy preference values to different groups (e.g., low/medium/high privacy).
+He et al.
+(He et al., 2014) proposes Blowfish Privacy in which in addition to specifying $\epsilon$, individuals specify their privacy policies consisting of secrets, i.e., information that must be kept secret about the individuals, and constraints, i.e., information that may be known publicly and exploited by adversaries.
+Both works attempt to define the granularity of privacy protection on the individual level but put the burden on the users to specify their privacy preferences.
+Instead, PRI is interpretable and controllers can specify their privacy preferences by observing the PRI of each individual directly.
 8. Conclusions
-In this paper, we showed that by including individual-level information about privacy loss (the PRI), controllers gain additional understanding of the effects of applying DP to the results of an analyst’s query.
-This translates into less burden choosing $\epsilon$: we implement this idea via an algorithm, Find-$\epsilon$-from-PRI.
-In addition, we show that it is possible to augment analysts’ understanding of the effects of DP on their results by using a utility signaling protocol that reveals the isolated impact of DP on a downstream task.
-We then gather the requirements gleaned from the previous two contributions and propose a platform, based on a data escrow, to deploy both ideas.
+In this paper, we showed that by including individual-level information about privacy loss (the PRI), controllers gain more understanding of the effects of applying differential privacy to the output of an analyst’s query.
+This translates into less burden in choosing $\epsilon$: we implement this idea via an algorithm, Find-$\epsilon$-from-PRI.
+In addition, we propose an alternative algorithm, called Find-and-release
+-$\epsilon$-from-PRI
+, which finds and releases $\epsilon$ privately.
+We present solutions to bound privacy loss when answering multiple queries without requiring controllers to set a global privacy budget.
+Overall, our work reduces the burden for controllers to use DP in practice.